A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

1 [904]. 2 × 2 [668, 800]. β [569]. χ^{2} [920]. d′ [569]. n [606]. P [1159]. t [920].

-Series [1159]. -Server [606]. -Test [920].

1 [22, 424, 1261]. 1750 [338]. 1922 [166]. 1924 [456]. 1939 [55]. 1961 [943, 27]. 1965 [1076, 406]. 1967 [112, 86]. 1968 [310, 534, 260, 109]. 1970 [320, 430]. 1971 [688]. 1973 [858]. 1974 [589, 726]. 1976 [1170].

2 [533, 134, 820]. 200th [1321]. 2015 [1169]. 2nd [807].

3 [1062]. 360 [295].

5 [1203].

6002 [869]. 67 [141]. 68 [80]. 69 [195].

71 [655]. 72 [1232]. 74 [1282]. ’75 [1003].

77. [1159]. 7th [1252].

Abraham [957]. Abridged [157]. Absence [643]. Abstract [341, 260]. Accessions [34, 64, 94, 115, 148, 170, 199, 232, 274, 306, 346, 414, 445, 488, 514, 547, 582, 624, 661, 692, 724, 757, 791, 829, 864, 899, 931, 962, 1013, 1043]. Accident [261]. Accidents [261]. Account [1150, 1304]. Accountability [1234]. Accountants [52]. Accounting [1246, 986]. Accounts [54, 196, 1208]. Accumulating [120]. Accuracy [1091, 733, 1241]. Acheson [1109]. Act [310]. Activities [25]. Activity [196]. Actuaries [131, 131]. Actuary [1169]. Acturial [367]. Adamu [1027]. Addendum [90]. Additive [732]. Adelstein [957]. Adi [810]. Adjustment [732]. Administration [521]. Administrative [466]. Adrian [671, 877]. Advanced [977, 1095, 325, 42]. Advances [967, 461, 886]. Advisory [430, 1019]. Afriat [885, 1172]. African [745]. After [40]. Against [155, 613]. Age [151, 233, 669, 836]. Age-Specific [836]. Aged [376]. Ager [643]. Aggregate [6, 243]. Agricole [264]. Agricultural [309]. Agriculture [650, 21]. Agronomiques [22, 134]. Aigner [1107]. Ainsworth [304]. Air [143]. Aird [1109]. Aitchison [1096, 155, 364, 15]. Ajit [336]. Alan [751, 1030, 245]. Albers [779]. Albert [896]. Aldcroft [55]. Alder [1102, 469]. ale [108]. Alexander [785]. Alfréd [525, 92]. Algebra [809, 51]. Algebras [465]. Algol [80]. Algorithm [904, 973]. Algorithmic [80]. Algorithms [1160]. Alison [270]. Allan [1078]. Allard [818]. Allen [822, 263, 926]. Allocation [39, 264, 970, 1254, 721]. Allocations [969, 873]. Alternative [271]. Ambiguity [524]. America [753]. American [481, 335]. Among [410, 1136]. Analyse [505, 881, 992]. Analyses [588]. Analysing [1308, 843]. Analysis [313, 1118, 1302, 401, 977, 466, 1244, 1034, 644, 158, 1047, 335, 976, 1051, 281, 268, 811, 1126, 263, 797, 996, 291, 1119, 704, 921, 1323, 1135, 297, 1297, 883, 178, 90, 14, 332, 1194, 740, 251, 183, 163, 1050, 847, 599, 941, 994, 1023, 271, 701, 367, 707, 1069, 1228, 314, 672, 1284, 334, 940, 746, 850, 995, 1192, 813, 521, 184, 736, 876, 357, 1137, 1279, 500, 611, 1326, 947, 1112, 404, 123, 1153, 836, 1090, 42, 651, 1200, 540, 1121, 890, 1320, 246, 610, 713, 1096, 410, 425, 393, 846, 376, 814, 1097, 1287]. Analysis [912, 502, 319, 1193, 1191, 265, 1320, 298, 509, 1059, 1060, 1295, 812, 258, 1072]. Analytic [1290, 190]. Analytical [1058]. Anant [425]. Anatomy [196]. Anderson [772, 941, 357, 1040, 595]. Andrew [895, 410, 376, 1315]. Andrews [461, 508]. Ann [1029]. Annals [352]. Anne [130]. Anniversary [478, 479, 541, 1321]. Annotated [59, 109]. Annual [102, 209, 355, 497, 632, 768, 909, 1057, 1190, 1313, 1047, 1039, 53, 101, 208, 354, 496, 631, 767, 908, 1054, 1187, 1311]. Answers [468]. Anticipated [1248]. Antituberculeuse [340]. Antonio [671, 877]. Anwendung [947]. APL [295]. APL/360 [295]. Aplicatii [108]. Applicable [657]. Application [466, 644, 528, 141, 1050, 1134, 1033, 334, 188]. Applications [234, 1207, 358, 1164, 22, 134, 288, 1267, 462, 465, 999, 328, 1205, 105, 566, 1094, 989, 1234, 432, 944, 329, 946, 130, 318, 842, 639, 1007, 437, 810, 500, 998, 529, 532, 463, 575, 887, 12, 890, 46, 327, 1170, 370, 194]. Applied [709, 948, 121, 923, 1255, 703, 78, 72, 704, 921, 133, 253, 571]. Appointments [5]. Appraisal [139, 1050, 1068, 267]. Approach [1236, 787, 818, 1077, 71, 290, 452, 1134, 173, 275, 1227, 214, 50, 484, 739, 79, 922, 291, 397, 1132, 295, 1104, 941, 772, 708, 253]. Approaches [1264]. Approximate [606]. Approximating [1278, 1332, 458]. Approximation [800]. Approximations [972, 975, 255]. April [320]. Archaeological [363]. Archaeology [996]. Archives [1273, 648]. Argentine [27]. Arithmetic [294]. Armartya [618]. Armitage [1177, 911, 285]. Armstrong [1322]. Arnold [564, 1125, 887]. Arrow [164]. Arthur [860, 259, 719, 576, 377]. Artis [1295]. Arunachalam [375]. Asbestos [1050]. Ash [186]. Ashton [464]. Aspects [889, 678, 945, 483, 608, 452, 73, 435, 374]. Assay [1316]. Assessing [585].

Assessment [832, 1019]. Assessments [1242]. Assisted [290]. Aster [374]. Asymptotic [975, 597, 779, 1259]. Asymptotically [278]. Atcherley [1136]. Atherton [682]. Athreya [504]. Atkinson [1209]. Atlantic [615]. Atlas [165]. Attainment [1308]. Attempt [122]. Attribute [255]. Auflage [324]. Auspices [320]. Austin [965, 286]. Auswertungsmethoden [324, 17]. Authorities [1243, 410, 173, 275]. Autocorrelation [677]. Autocorrelations [1277]. Avenhaus [1234]. Averill [470]. Avi [372]. Avi-Itzhak [372]. Aycliffe [88]. Aydelotte [646]. Aykac [1205]. Ayres [1267].

B [776, 1100, 544, 442, 504, 399, 920, 1039, 304, 1323, 883, 52, 1289, 258, 777, 1166, 1254, 750, 226, 803, 1003, 1133, 1159, 1162, 1196, 820, 367, 612, 341, 1209, 1102, 778, 221, 1001, 982, 186, 1031, 1229, 643, 929, 884, 83, 469, 1063, 325, 595, 262]. B. [244]. Babbage [1217]. Baccarat [699, 871, 871]. Baccarat-Banque [871]. Bacharach [224]. Back [35, 65, 95, 116, 149, 171, 200, 235, 276, 307, 347, 380, 415, 447, 489, 515, 549, 583, 625, 662, 693, 727, 758, 793, 830, 867, 900, 932, 963, 1014, 1044, 1085, 1115, 1145, 1179, 1212, 1239, 1269, 1305, 1331]. Bagchi [459]. Bailey [889, 777, 59, 292]. Balaam [534, 506]. Balakrishnan [1254]. Balance [1008, 1267, 266]. Ball [1322, 206]. Ballonoff [1065]. Banerjee [1236, 851]. Banker [310]. Banks [688]. Banque [871]. Barlow [528]. Barndorff [1221]. Barndorff-Nielsen [1221]. Barnett [1218, 590]. Baron [927]. Barrie [49]. Barry [268, 580]. Bartholomew [683, 528, 1315, 289, 592]. Bartlett [951, 995, 18, 1123]. Barton [410, 376]. Based [732, 955, 733, 554, 522, 218, 255, 596, 27]. Basic [339, 366, 920, 1290, 1219, 293, 711, 739, 127, 1171, 429, 639, 222, 186, 1030, 16]. Basil [163]. Bassett [289]. Batra [619, 824]. Battacharya [975]. Battersby [371]. Baum [987]. Bayesian [457, 608, 564, 845, 1205, 797, 216, 565, 334, 940]. Beale [194]. Bechhofer [84]. Beck [1125]. Becoming [762, 866]. Bedworth [611]. Beef [334]. Beginners [1027]. Beginning [1028]. Beginnings [77]. Behavior [402, 606, 138, 226]. Behavioral [296, 193]. Behaviour [747, 137, 834, 351, 652]. Behavioural [44, 921, 993, 1023, 1284, 252]. Being [655, 1266]. Belgian [334]. Belinfante [465]. Bell [1204]. Bellagio [1154]. Belz [897]. Ben [706, 810]. Ben-Israel [810]. Benefit [265, 540]. Benjamin [811, 603, 367, 372, 83]. Bennett [993, 456]. Berger [1036]. Berkeley [527, 455, 392, 498, 526]. Berlin [821]. Bernard [465, 685, 538, 42, 1296]. Bernoulli [462]. Bert [402]. Bertier [881]. Beta [903, 904]. Between [390, 686, 268, 687, 1327, 1122]. Beveridge [725, 179]. Bhattacharyya [1128]. Bibby [1223, 745]. Bibliographie [324]. Bibliographies [24]. Bibliography [1, 1144, 180, 554, 59, 69, 109, 608, 534, 24]. Bickel [461]. Bienayme [1248]. Billingsley [1127]. Binary [184]. Binomial [255]. Bioassay [464]. Biological [19, 1316, 946, 1263, 1321]. Biologists [636]. Biology [358, 858, 777, 566, 1231, 1168, 159]. Biomathematics [1294, 18]. Biomedical [1159, 1162]. Biometrical [676]. Biometrika [433]. Biometrischen [160]. Biometry [21, 506, 19]. Bioscience [1164]. Biostatistics [358, 563, 605, 740]. Biproportional [224]. Birnbaum [1133, 1078]. Bishop [876, 91]. Bivariate [1185, 213, 58]. Bjerke [87]. Blackaby [895, 1282]. Blackith [362]. Blackman [296]. Blackstone [112]. Bleddyn [410, 376]. Bliss [159]. Bloomfield [976]. BMDP [1159]. BMDP-77. [1159]. Boas [1327]. Boat [613]. Bock [1201]. Bodmer [373]. Bogue [646]. Bolch [706]. Boldrini [356]. Bonroche [881]. Book [1197, 817, 1172, 459, 776, 859, 1207, 1236, 286, 339, 1036, 1175, 246, 358, 482, 527, 594, 889, 686, 128, 247, 858, 1203, 139, 915, 1231, 401, 784, 561, 155, 215, 320, 359, 457, 633, 703, 939, 977, 1094, 593, 809, 466, 601, 680, 743, 989, 1264, 1283, 852, 978, 365, 508, 573, 600, 646, 808, 1111, 264, 212, 364, 917, 1222, 50, 43, 189, 1022, 161, 407, 709, 710, 711, 782, 1100, 1108, 1109, 1232, 822, 650, 360, 435, 981, 1034, 653, 1113, 76, 997, 1068, 484, 317, 544, 644, 140, 84, 158, 366]. Book [598, 442, 361, 504, 45, 195, 1234, 373, 948, 821, 326, 1004, 616, 56, 1226, 744, 678, 481, 19, 21, 126, 563, 335, 533, 1173, 269, 849, 888, 945, 399, 618, 1061, 185, 323, 602, 712, 927, 976, 1025, 1101, 1139, 1169, 860, 467, 330, 505, 610, 1024, 1095, 1161, 688, 47, 483, 512, 747, 635, 737, 951, 739, 878, 605, 1210, 44, 565, 920, 54, 1318, 436, 928, 127, 268, 677, 713, 811, 316, 683, 684, 1233, 753, 787, 78, 942, 137, 57, 301, 79, 1164, 1103, 1298, 468, 1096, 55, 473, 1039, 1325]. Book [80, 634, 950, 615, 1128, 1035, 395, 537, 613, 210, 299, 652, 72, 1126, 75, 219, 1142, 225, 1171, 1201, 129, 217, 1008, 528, 570, 922, 956, 22, 134, 160, 464, 81, 82, 400, 499, 591, 617, 818, 885, 926, 1073, 263, 949, 1125, 619, 824, 403, 990, 607, 812, 985, 996, 1165, 1288, 291, 895, 259, 704, 374, 567, 227, 749, 60, 925, 1290, 654, 397, 190, 229, 296, 410, 786, 921, 1323, 432, 720, 1135, 1158, 297, 402, 578, 674, 944, 979, 338, 224, 1297, 1077, 569, 883, 590, 30, 143, 597, 771, 910, 954]. Book [955, 975, 1064, 1132, 1198, 1065, 1123, 52, 288, 913, 1289, 1291, 1177, 46, 719, 738, 1168, 880, 294, 300, 655, 642, 856, 1267, 891, 472, 886, 1294, 20, 188, 258, 295, 329, 425, 433, 456, 714, 1218, 1316, 1075, 393, 14, 534, 717, 892, 1160, 1217, 141, 409, 823, 332, 846, 993, 1105, 461, 777, 673, 430, 462, 465, 718, 946, 1141, 1166, 1254, 1286, 368, 51, 1131, 604, 742, 1194, 228, 748, 980, 1000, 1258, 302, 750, 879, 923, 1098, 438, 687, 429, 220, 740, 251, 1127, 1067, 1072, 226, 1041, 162, 803, 376, 814]. Book [1256, 987, 1124, 751, 1028, 1104, 112, 156, 261, 290, 676, 741, 1140, 1224, 24, 182, 183, 187, 458, 525, 1248, 1263, 13, 130, 250, 318, 455, 501, 671, 715, 735, 779, 780, 877, 919, 1058, 1221, 1253, 1285, 1319, 640, 1202, 1252, 193, 218, 562, 752, 59, 854, 609, 1003, 804, 1219, 918, 163, 106, 107, 108, 17, 135, 166, 539, 1037, 1266, 1038, 196, 255, 284, 577, 142, 788, 1206, 216, 283, 392, 423, 424, 608, 1071, 1317, 337, 649, 847, 914, 371, 599, 842, 941, 745, 943, 23, 324, 675, 816, 1281, 754, 460, 1163, 881]. Book [984, 994, 1023, 1097, 1133, 1159, 1162, 1196, 1287, 1321, 953, 820, 1062, 333, 603, 431, 271, 477, 478, 657, 28, 133, 819, 131, 367, 612, 639, 707, 785, 912, 986, 999, 1007, 1069, 1228, 682, 770, 1315, 132, 369, 1134, 15, 685, 1322, 1220, 502, 159, 319, 672, 1033, 1193, 1284, 1009, 88, 249, 476, 86, 87, 103, 154, 341, 479, 574, 1005, 1076, 1176, 1208, 1209, 73, 322, 580, 857, 474, 638, 911, 157, 289, 334, 336, 396, 434, 940, 1027, 1102, 1174, 1251, 1261, 136, 213, 530, 746, 778, 850, 995, 1191, 1192, 411]. Book [957, 25, 53, 192, 506, 636, 772, 773, 813, 882, 221, 58, 475, 265, 184, 191, 702, 736, 775, 806, 841, 876, 974, 1001, 807, 579, 222, 254, 538, 541, 982, 426, 637, 843, 983, 357, 510, 1320, 437, 716, 781, 186, 293, 536, 689, 681, 706, 810, 1026, 1031, 1032, 1129, 1137, 1138, 1167, 1195, 1225, 1227, 1279, 1280, 1324, 606, 223, 500, 535, 498, 327, 1250, 855, 998, 708, 398, 958, 992, 1136, 110, 260, 1229, 643, 480, 256, 439, 564, 845, 1006, 1040, 929, 1199, 214, 705, 645, 1293, 138, 164, 165, 441, 611, 1327, 248]. Book [287, 292, 596, 769, 840, 1130, 1249, 1259, 1265, 1326, 815, 18, 328, 375, 526, 641, 1030, 1070, 1204, 1205, 1223, 363, 298, 405, 1010, 1099, 529, 576, 924, 267, 947, 427, 884, 1110, 1112, 1170, 11, 125, 377, 542, 783, 507, 509, 408, 1059, 470, 471, 404, 851, 531, 774, 16, 77, 83, 104, 469, 543, 614, 647, 1260, 1292, 370, 1002, 991, 1063, 372, 532, 440, 1074, 848, 952, 285, 428, 105, 194, 331, 463, 568, 592, 679, 853, 916, 1066, 1230, 1262, 325, 26, 49, 1235, 321, 144, 1060, 721, 575, 42, 690, 887, 1107]. Book [572, 29, 48, 109, 252, 253, 257, 270, 571, 805, 844, 894, 1029, 1106, 1257, 1282, 1295, 1296, 85, 362, 394, 406, 503, 511, 595, 656, 27, 266, 262, 340, 566, 648, 651, 1200, 893, 12, 41, 74, 211, 356, 540, 1255, 988, 890, 245, 1314, 111, 1251]. Bookstein [1263]. Boots [1196]. Boreham [710]. Boschan [297]. Boullion [360]. Bower [196]. Bowers [993]. Bowman [1138]. Box [183, 941, 1279, 449, 1314]. Bracewell [268]. Bracewell-Milnes [268]. Bradford [965, 286]. Bradley [11]. Bradshaw [383]. Brain [84]. Branching [504, 946]. Braskamp [193]. Breiman [12]. Bremner [528]. Breny [374]. Brian [713, 942, 1250, 1257]. Brinley [481, 615]. Britain [615, 943, 262, 667, 798, 1151, 1209, 1002]. British [422, 401, 481, 335, 483, 55, 410, 163, 341, 869, 492, 937, 938, 1173, 260, 1282]. Bronchitis [1108]. Brown [436, 1159, 1162, 915, 639]. Brownian [395]. Brownlee [1276]. Brumat [1205]. Brunk [528]. Bruno [671, 877]. Bry [297]. Bryan [655, 441]. Bryant [1136, 1136]. Buckland [411, 474]. Buckley [815]. Budapest [775]. Buddy [218]. Budgets [404]. Buffers [990]. Buffon [1321]. Buhler [1036]. Buhlmann [1037, 189]. Building [81, 1230, 1279]. Bulgren [290]. Bulk [459]. Bureau [162, 478, 542, 647]. Burford [222]. Burr [997]. Burrington [473]. Burstein [255]. Bury [923]. Business [1061, 467, 57, 1183, 291, 141, 1202, 639, 986, 222, 538, 541, 681, 706, 223, 998, 253, 43, 785, 254]. Buxton [1173]. Buying [226, 437]. Byrnes [612].

C [459, 401, 508, 600, 1222, 711, 76, 544, 598, 821, 1226, 399, 712, 1101, 79, 613, 299, 225, 607, 227, 749, 1158, 880, 329, 714, 892, 748, 1098, 1067, 1072, 1248, 752, 424, 754, 1163, 819, 639, 682, 369, 103, 1076, 778, 850, 636, 702, 841, 843, 437, 1031, 256, 1040, 1259, 1205, 1099, 529, 267, 11, 377, 1059, 1260, 1074, 568, 325, 690, 1029, 394, 406]. C. [622]. C.B.E [304]. Calculations [396, 988]. Calculator [537]. Calibration [45, 1216, 607]. Cam [527, 455, 392, 498, 526]. Campbell [1320, 636]. Canada [320]. Canadian [434, 1047, 1048]. Cancer [585, 1182, 151, 233, 1093, 892]. Candidates [157]. Capacity [28]. Capital [1275, 279, 203, 822, 953, 477, 267]. Car [1089]. Caradog [659]. Carbon [1216]. Carbon-14 [1216]. Carcinogenesis [151, 233]. Care [594, 482]. Carl [737, 924]. Carlo [239]. Carnap [424]. Carnegie [271]. Carnoy [540]. Carrier [368]. Carsberg [52]. Case [969, 449, 987, 540, 1099]. Case-Study [449]. Cases [1297]. Casino [280]. Cassimatis [142]. Casti [1290]. Castle [637, 983]. Categorical [1194, 672]. Caulcott [602]. Causal [410, 376]. Causation [261]. Cause [1182]. Causes [1018]. Cavalli [373]. Cavalli-Sforza [373]. Cecily [1108]. Cell [668, 800, 1294]. Census [40, 492, 588, 494, 309, 475, 83, 647]. Censuses [577, 122]. Cent [475]. Central [1183, 577, 1208]. Century [523]. Certain [258]. Chain [1149]. Chains [459, 989, 458]. Challenge [717]. Chambers [1193]. Champernowne [686]. Chance [155, 206]. Change [403, 786, 243, 224, 1263]. Characteristics [241]. Characterization [778]. Characterizations [1224]. Charles [1108, 1217, 221, 1227]. Charlotte [297]. Charnes [1264]. Chatfield [76, 450]. Chatterjee [1323]. Check [475]. Chemical [426]. Chemin [699]. Chemin-De-Fer [699]. Chemistry [883]. Chervany [707]. Chester [159]. Chesterton [612]. chez [340]. Chi [125]. Chi-Squared [125]. Chiang [1320]. Children [700, 410, 10, 176, 353, 953]. Chin [1320]. Chisholm [483]. Chistyakov [1254]. Choice [764, 773, 155]. Choudhry [434]. Chow [361, 1319, 1072]. Chris [532]. Christiansen [1166]. Chronic [1108]. Chung [770]. Cibernetice [108]. Circulation [56]. Civil [1066]. Clark [704]. Clarke [1219, 776, 910]. Classical [216]. Classification [610, 174, 586]. Classifications [1118, 1302, 1112]. Classified [1194]. Claudia [924]. Cleland [1157]. Clelland [639]. Cliff [710, 677, 706, 1200]. Clifford [1010]. Clinical [832, 374, 883]. Clive [996]. Close [255]. Clough [748]. Cluster [1245, 713, 811]. Co [680, 214]. Co-Ordinate [214]. Co-Ordinated [680]. Coates [262]. Coats [1229]. Coding [240]. Coefficients [822, 822]. Coen [153, 1039]. Cohen [921, 921, 74]. Cohort [1050, 1093, 836]. Coleman [771]. Colin [895, 432, 912]. Collected [456]. Collecting [181]. Collection [951, 1183, 655, 604, 396]. Colleges [589, 726, 468]. Collings [366]. Colloquium [478, 479, 541]. Colquhoun [358]. Combination [551]. Combinatorial [600, 330, 263]. Comment [725]. Comments [153, 450, 968, 766, 244, 1048]. Commercial [589, 726, 370]. Commission [891, 271, 29]. Commitments [68]. Committee [589, 726, 1019, 971, 301, 162, 131, 969, 873, 1272]. Commodities [652]. Commodity [225, 687]. Common [71, 1171]. Commonsense [743]. Commonwealth [271]. Communication [1017, 106]. Communities [493, 1325]. Community [1109, 454, 642, 408, 1177]. Companion [572]. Company [906]. Comparative [802, 590]. Comparing [1243]. Comparisons [1009]. Competing [1318]. Competition [1325]. Complex [812]. Comprehensive [608].

Computation [653, 227]. Computational [1290, 465, 994, 1193]. Computer [903, 1034, 753, 699, 871, 654, 297, 554, 290, 1185, 1159, 419, 222, 1327, 924, 1322, 47, 1229]. Computer-Assisted [290]. Computer-Based [554]. Computer-Oriented [222]. Computers [418, 816]. Computing [537, 609]. Concentration [124, 71, 943]. Concept [779]. Concepts [45, 127, 1128, 980, 598, 999, 884]. Conceptual [291]. Concerning [417]. Conditional [780, 595]. Conference [519, 858, 195, 430, 1041, 162, 542, 844, 1154, 1175, 822]. Confidence [255, 385]. Conflict [484, 112, 509]. Conflicting [1204]. Connor [1252]. Conolly [942]. Consequences [585]. Constance [11]. Constituency [588]. Constraints [288]. Construction [1149, 524, 142]. Constructions [330]. Consultancy [37]. Consumer [652, 6, 90, 1038, 385, 763]. Consumption [895, 279, 86, 510]. Contacts [587]. Contemporary [1132, 1104, 252, 1295]. Contingency [175, 668, 800]. Continuous [499, 321]. Contracting [1165]. Contributions [948, 1255]. Control [997, 730, 1289, 1072, 649, 1227, 917, 709, 1098, 183, 371, 645, 611, 49]. Controversy [608, 377]. Convergence [41]. Cooke [1219]. Cooley [393]. Cooper [1264]. Coping [70]. Cornish [559]. Correction [726, 1144]. Corrections [827, 766, 826, 825]. Correlation [873, 217, 921, 885]. Correlational [1258]. Correlations [1052]. Corrigenda [1302, 865, 866, 1304, 1332, 1301, 1303]. Corrigendum [233, 275]. Cost [1236, 733, 765, 265, 851, 540]. Cost-Benefit [265, 540]. Costs [788]. Cottle [679]. Council [102, 209, 355, 497, 632, 768, 909, 1057, 1190, 1313, 1022]. Countable [458]. Countries [687]. Country [1153]. Counts [156]. Course [365, 709, 21, 129, 979, 1219, 657, 708, 77, 572, 471]. Coutts [1235]. Covariation [1136, 1136]. Cowell [1176]. Cox [769, 1249, 1035, 184, 1328]. Cozzolino [913]. Craig [943]. Cramer [104, 187]. Craven [1289]. Creamer [110]. Credibility [944]. Criminology [333]. Crisis [409]. Critical [671, 877]. Critique [580]. Crockett [862]. Croft [78]. Croissance [264]. Cross [1118, 1302, 1216, 1194, 1326, 1112]. Cross-Classifications [1118, 1302, 1112]. Cross-Classified [1194]. Cross-Tabulated [1326]. Cross-Validation [1216]. Croxford [544]. Crutchfield [879]. Crystal [1322]. Csepinszky [1041]. Cullen [1002]. Cultural [244, 99]. Cumbernauld [88]. Curious [873, 1052]. Current [31, 62, 89, 113, 145, 167, 197, 230, 272, 303, 342, 378, 412, 443, 485, 545, 620]. Cushing [1197]. Cycle [1049]. Cycles [141, 541, 481]. Cyclical [402, 297]. Cytological [374].

D [246, 358, 686, 561, 320, 1094, 989, 852, 808, 711, 544, 744, 1173, 712, 1024, 1161, 483, 635, 565, 677, 683, 1233, 753, 78, 210, 652, 528, 464, 926, 1288, 190, 229, 1289, 472, 1294, 1316, 332, 461, 673, 1131, 748, 429, 1127, 226, 1319, 854, 1219, 847, 599, 941, 23, 1281, 1133, 820, 1078, 289, 265, 184, 1001, 716, 681, 1031, 1225, 998, 439, 929, 611, 769, 1249, 815, 363, 267, 1112, 440, 568, 592, 853, 1107, 1200]. Dagnelie [22, 134, 992]. Dalenius [1274]. Damaraju [330]. Daniel [110, 292, 740]. Darnley [1083]. Darwinian [526]. Data [971, 870, 1244, 644, 1271, 922, 518, 296, 1308, 1135, 179, 1297, 279, 393, 604, 180, 421, 725, 120, 243, 588, 319, 1193, 1284, 1191, 1279, 1048, 1093, 836, 1121, 1272, 269, 1161, 920, 883, 1218, 368, 1194, 847, 672, 850, 184, 736, 843, 1250, 1326, 428, 1192, 843, 507]. Daten [539]. Dating [141]. David [1109, 361, 1318, 395, 823, 993, 879, 1071, 1315, 574, 611, 1204, 1112, 1029, 1106, 359, 365, 659, 63, 1255]. Davidson [635, 599].

Davies [410, 803, 376, 426]. Davis [850, 298, 651]. Day [432]. Dean [895]. Decade [1183]. Decades [643]. DeCani [639]. Decennial [1232]. Decision [155, 212, 1222, 747, 1126, 1165, 294, 250, 1071, 912, 999, 1069, 1228, 1090, 383, 457, 467, 707, 940, 716, 571, 399, 612]. Decision-Making [747, 383, 707, 571]. Decisions [1017, 1202, 284, 1204, 105, 1070]. Decline [311]. Decomposition [466, 1140]. d’Econometrie [135]. Défaillances [340]. Deficiency [779]. del [1009]. Delaunois [605]. Delay [1197]. Delinquency [144]. Delphi [1007]. Demand [68, 578, 143, 90, 579, 266, 1038]. Demetrios [856]. Demographic [925, 368, 1065, 368]. Demography [533, 407, 684, 1233, 1035]. Dem’yanov [853]. Dennis [400, 284]. Density [1120, 1140, 943]. Department [320, 260]. Departmental [577]. Departments [816]. Dependence [241]. Dépenses [264]. Derek [55]. Derivation [1275]. Derivations [218]. Descartes [320]. Description [916]. Descriptive [1104]. Design [977, 330, 967, 1291, 642, 258, 1, 14, 534, 1144, 814, 1097, 192, 882, 1279, 641, 408, 1264, 1099, 772]. Designing [1010]. Designs [799, 1301, 812, 813]. Desk [537]. Detection [472, 368]. Determinants [173, 275, 763]. Determines [969, 873]. Determining [255]. Deterministic [917, 781]. Detroit [512]. Developing [368, 1153]. Development [784, 481, 338, 823, 337, 476, 309, 1009]. Developments [1051, 492, 1205, 48, 651, 1280, 372]. Deviates [58]. Deviations [1091]. Diagnosis [801, 556]. Diamond [377]. Dick [429]. Dickinson [986]. Dictionary [474, 752]. Dietrich [607]. Differential [535, 887]. Differentials [557]. Diffusion [395, 1168]. Diffusions [1319]. Digital [502]. Dillon [1287]. Dimensionelles [881]. Dimensions [646]. Dinamice [108]. Dinter [609]. Directional [428]. Director [655].

Dirichlet [1261]. Discret [501]. Discrete [75, 1140, 1185, 130, 1134, 919, 1287, 530, 876]. Discrete-Parameter [919]. Discriminant [1287, 846]. Discussion [1274, 244]. Diseases [696]. Dispersion [746]. Display [850]. Distances [1122]. Distinction [268]. Distributed [872]. Distribution [587, 686, 151, 233, 891, 1278, 1332, 666, 698, 731, 1266, 396, 1261, 520, 596, 11, 616, 895, 255, 125, 29, 1209, 596]. Distribution-Free [396, 11]. Distributions [1061, 175, 734, 802, 1278, 1332, 750, 1091, 1163, 669, 532, 75, 499, 1140, 187, 780, 460, 87, 213, 536, 1031, 321, 75, 499, 321]. Disturbances [1260]. Diverse [300]. Divisible [1253]. Division [687]. Dixon [1159]. Doctor [587]. Dodge [1299]. Dolezel [59]. Donald [737, 291, 918, 1265, 862, 1083]. Donna [950]. Données [881]. Doses [637, 983]. Douglas [562, 755]. Dr [722]. Draft [869]. Drago [1009]. Drain [84]. Draper [608]. Drew [1010]. Dreze [721]. Driebeek [133]. Driessche [858]. Dritte [324]. Drygas [214]. Dual [875]. Dubbey [1217]. Dubbleman [1260]. Duda [610]. Dudewicz [981]. Dudley [409, 576]. Dugdale [60]. Duncan [854, 576]. Dunn [704, 786]. Dunsmore [1096]. Durables [763]. Duran [811]. Duration [282]. Durbin [1274, 596]. Dwyer [808]. Dynamic [1175, 822, 484, 152, 281, 1288, 1290, 1072, 250, 250, 575]. Dynamics [802, 1197, 614].

E. [559]. Earl [1265]. Earley [144]. Early [1243, 1002]. Earnings [644, 667, 177]. Easily [973]. Easy [396]. Eccleston [1296]. Ecklar [647]. Ecnomic [1072]. Ecological [1325]. Econometric [269, 173, 275, 434, 439, 26, 402, 1265]. Econometrics [845, 1171, 1201, 400, 617, 818, 1073, 819, 249, 1005, 564, 104, 1006, 1074]. Economic [422, 1207, 585, 207, 653, 152, 335, 618, 747, 2, 316, 753, 615, 338, 178, 1177, 823, 162, 1266, 478, 103, 479, 1174, 254, 538, 165, 542, 181, 1107, 29, 1029, 1106, 1282, 1295, 262, 111, 784, 478, 479, 541, 614, 511]. Economics [482, 1113, 317, 1267, 734, 819, 222, 439, 554, 809, 161, 1034, 744, 741, 857, 136, 681, 706, 223, 440, 1257, 82, 142, 953]. Économie [264]. Economies [269]. Economist [377]. Economists [743, 431, 221]. Economy [730, 55, 166, 351, 401, 483, 1298, 1039, 615, 434, 783, 26, 1296]. ed [807]. Eden [912]. Edgar [786, 717]. Edgeworth [1147]. Edition [1252]. Education [68, 468, 271, 764, 241, 510, 293, 351, 540, 589, 726, 219, 880, 251, 603]. Educational [1308]. Edward [219, 1102, 469]. Edwards [685, 538]. Edwin [978]. EEC [697]. Effect [454, 510]. Effects [527, 544]. Efficiency [510]. Efficient [278]. Eggleston [1317]. Egon [695]. Ehrenberg [843, 437]. Eichhorn [1207]. Eighth [354]. Einführung [107]. Eisen [13]. Eisner [377]. Elam [253]. Elandt [299]. Elandt-Johnson [299]. Elden [11]. Eleanor [601]. Election [761]. Elections [311, 938, 937]. Electoral [243]. Electric [377]. Electrical [316]. Electricity [279]. Electronic [1297]. Elementary [290, 218, 253, 674, 1131, 980, 984, 770, 298, 257]. Elements [689, 220, 1127, 11, 470, 135]. Eli [782]. Eliasson [1298]. Elisa [258]. Elizabeth [527, 455, 392, 498, 480, 526]. Elliot [1129]. Eltis [616]. Elzey [254]. Emde [1327]. Emphysema [1108]. Empirical [879, 1241, 579, 436, 1163, 104]. Employees [798]. Employment [544, 1173, 71, 1040, 260]. Energy [1267]. Engineering [316, 1125, 667, 427]. Engineers [1124, 657]. England [1232, 892, 475, 629, 1243, 523, 309, 1093]. Engleman [1159]. English [411, 311, 752]. Enochson [502]. Enrick [218]. Enter [391]. Entropy [124, 71]. Entscheidungs [1037]. Entscheidungs- [1037]. Environment [453, 1267, 1206]. Epidemic [1276, 1148]. Epidemics [781]. Epidemiological [836]. Epidémiologie [340]. Epidemiology [507]. Epochs [462]. Equation [854, 530, 576]. Equations [1197, 152, 887, 535]. Equilibria [653]. Equilibrium [669, 721]. Erdwissenschaftler [23]. Erhardt [821]. Erling [595]. Ernest [91]. Ernst [107]. Errata [417]. Erricker [325]. Error [530]. Errors [368]. Erweiterte [324]. Essays [784, 1113, 823, 951, 482]. Essential [1059]. Essentials [128, 956, 712, 1251]. Establishment [269]. Establishments [53]. Estimates [1275, 461, 937]. Estimating [875, 451, 822]. Estimation [1150, 1304, 1125, 668, 368, 396, 530, 1138, 1223, 774, 1260, 1184, 571, 782, 1032, 214, 856]. Estimators [1201]. Ethnomethodology [580]. Eugene [645, 41]. Europe [29, 29]. European [1177, 775, 493]. Evaluation [301, 1010]. Evaluations [555]. Evans [26]. Evelyn [602]. Events [875]. Everitt [713, 1250]. Evidence [971, 1317]. Evolution [526]. Examination [157]. Examinations [8]. Example [152, 730, 785, 1153, 1323]. Examples [326, 290, 1293, 351, 77, 987, 888]. Executives [670]. Exercises [888, 468, 537, 326, 819]. Existing [670]. Expansions [779, 975]. Expectations [361, 137]. Expenditure [86, 1076, 763, 521, 1153, 406, 511]. Experience [387, 551, 643, 300]. Experiment [534]. Experimental [1161, 812, 188, 813, 192]. Experimentation [1067]. Experiments [1, 1144, 772, 1279, 330, 258, 14, 1097]. Expert [301]. Explain [453]. Explanations [255]. Exploratory [1191]. Exponential [1120, 1221, 1259]. Exports [178, 1265]. Exposition [734]. Exton [1210].

F [1175, 680, 743, 365, 600, 32, 822, 84, 373, 19, 563, 1024, 512, 635, 436, 928, 1103, 990, 607, 674, 883, 597, 20, 714, 461, 465, 1166, 1254, 742, 1000, 226, 803, 1263, 877, 193, 1266, 599, 1281, 657, 639, 1315, 1176, 1192, 192, 254, 606, 327, 1040, 363, 924, 1299, 568, 853, 1060, 270, 1282, 394, 340]. F.R.S. [1011]. Faber [409]. Factor [162, 332]. Factorial [1236, 586]. Factorielle [505]. Factors [764, 241]. Facts [407, 717]. Fairley [1100]. Fallacies [969]. Familial [696]. Families [1221, 1259, 460, 213]. Family [953, 404]. Fanger [883]. Farm [1265]. Farquharson [85]. Farr [957]. Farrell [1079]. Fatigue [1164]. Features [838]. Federal [109, 140]. Federer [534, 604]. Fein [271]. Feinstein [1076, 406]. Feller [318]. Female [838]. Fenchel [1166]. Fer [699]. Ferguson [251]. Field [690, 803, 1067]. Fienberg [876, 845, 1194]. Fifth [1311, 195]. Fiftieth [478, 479, 541]. Figures [289]. Fiji [337]. Fildes [1225, 998]. Final [544]. Finance [986]. Financial [670, 131]. Financing [271]. Find [473]. Findings [585]. Finetti [671, 877]. Finite [1288]. Finney [246, 1316]. Firm [336]. Firms [1061]. First [767, 365, 21, 979, 655, 77]. Fiscal [268, 164, 377]. Fishbein [648]. Fisher [456, 1314, 1133]. Fishman [249]. Fit [920]. Fix [451]. Fix-Point [451]. Fleischer [317]. Fleiss [570]. Fleming [917]. Fletcher [1108, 544]. Flieger [407]. Florence [335, 165]. Floud [573]. Flows [483, 872, 687]. Fogel [646]. Fogg [1008]. Folks [319]. Followed [244]. Food [86, 1265]. Foote [1206]. Forbes [1315]. Force [838]. Forecasting [313, 1175, 401, 152, 1089, 761, 1226, 553, 449, 2, 183, 941, 1241, 1049, 1322, 551, 1088, 571, 945, 1026, 1101, 1225, 998]. Forecasts [551, 225]. Foreclosure [144]. Foreign [27]. Forest [430]. Forestières [430]. Foreword [386]. Form [269]. Forms [468]. Formulation [530]. Fortieth [631]. Fortran [431, 656, 432]. Forty [767, 908, 1054, 1187, 1311]. Forty-Fifth [1311]. Forty-First [767]. Forty-Fourth [1187]. Forty-Second [908]. Forty-third [1054]. Forum [1206]. Fosback [1077]. Foundation [320, 740]. Foundations [320, 939, 423, 833, 1319, 320, 1073, 182]. Four [920]. Fourier [976]. Fourteenth [301]. Fourth [1187]. Fowlkes [258]. Fox [950, 103]. Frame [1159]. Frances [333, 475]. Francis [1147]. Frank [1108, 895]. Franklin [436, 1033]. Frankowski [1261]. Fraser [1094, 716]. Frechet [1080]. Fred [895]. Frederick [1100, 849]. Free [396, 214, 11]. Freedman [395, 458]. Freedom [1087]. Freeman [254, 569]. Freight [483]. French [26]. Frequency [460, 736]. Fresh [397]. Freund [287]. Frey [1200, 1200]. Fried [44, 985]. Frisch [486]. Fritz [477]. Front [4, 36, 66, 97, 118, 150, 172, 202, 237, 277, 308, 349, 382, 416, 448, 491, 517, 550, 584, 627, 664, 694, 728, 760, 795, 831, 868, 902, 934, 964, 1015, 1046, 1086, 1116, 1146, 1181, 1213, 1240, 1270, 1307]. Frontiers [476, 439]. Fruchter [603]. Fruit [1067]. Fuchs [482]. Full [334]. Fuller [1025]. Function [269, 729, 596]. Functions [808, 435, 1140, 218, 27]. Fund [271]. Fundamental [603]. Fundamentals [737, 913, 506]. Futures [442].

G [459, 686, 401, 809, 978, 600, 646, 1111, 822, 326, 888, 185, 467, 635, 920, 57, 1298, 1039, 1128, 1201, 217, 591, 926, 402, 1077, 30, 856, 673, 465, 748, 1000, 302, 226, 803, 156, 1224, 183, 1219, 599, 984, 657, 612, 999, 1069, 369, 341, 474, 411, 191, 1167, 1279, 606, 327, 855, 958, 110, 1040, 138, 1327, 1326, 375, 363, 405, 11, 783, 77, 440, 49, 1060, 1029, 1106, 245]. G.D.P. [152]. Gaines [377]. Gales [112]. Gallager [106]. Game [1113]. Games [310, 206, 331]. Gaming [310]. Gamma [903, 973]. Gani [775, 1123]. Gascoigne [1157]. Gauss [214]. Gauss-Estimation [214]. Gaussian [1310, 1324]. Geary [819]. Gee [475]. General [101, 100, 208, 354, 496, 495, 631, 767, 792, 908, 839, 1054, 1053, 1187, 1186, 1311, 1232, 1164, 892, 577, 914, 271, 937, 1284, 325, 1000]. Generalised [810]. Generalized [666, 698, 731, 312, 360, 329, 328]. Generating [973]. Generation [903, 904, 1185]. Genetic [718]. Genetics [373, 299, 1064, 1198, 1065]. Geoff [1008]. Geography [1200, 747, 994, 262]. Geometric [1286]. Geometrical [1321]. Geometry [635, 855]. George [1034, 1201, 251, 249, 1279, 245, 32, 621, 622]. Geostatistics [884]. Geraint [682]. Gerald [849, 914, 271, 1199]. Geraldina [1009]. Gerhard [885, 297, 221]. Germany [141, 872]. Gertrude [1328]. Getis [1196]. Gewinnung [539]. Ghana [10, 176]. Ghosal [888, 73]. Gihman [535]. Gilbert [676]. Gilchrist [1026]. Gillian [473]. Gillies [561]. Gilman [295]. Ginn [436, 512]. Giri [1158, 702, 841, 991]. Gittus [480]. Glossary [411]. Glynn [544]. Gnanadesikan [258, 1135]. Göbel [990]. Godambe [320]. Godley [1235]. Goldberger [576, 1107]. Goldfeld [400]. Goldfield [1032]. Goldsmith [426]. Goldstein [1287, 296]. Goldstine [654]. Goldthorpe [787]. Gomme [153]. Goodness [920]. Goods [1281]. Gordon [51, 751, 945]. Gore [860]. Goss [442]. Gossling [822, 1175, 822]. Gottfried [397]. Gottinger [940]. Gottlieb [322]. Goult [657]. Government [1169, 389, 109, 383, 335]. Graded [669]. Gradient [1120]. Grading [787]. Granger [256, 225, 725, 314]. Grant [645]. Grants [969, 873, 970]. Graphical [1017, 1250]. Graphs [719]. Gray [475]. Graybill [1033]. Great [615, 943, 798, 361]. Greek [752]. Green [1004, 1161]. Greensted [712]. Greenwood [1247, 1278, 1332]. Gregory [512, 1072]. Greville [810]. Griliches [269, 403]. Gronau [143]. Gross [737]. Grossman [579]. Group [430]. Grouped [774]. Groups [719, 465, 1136]. Growth [782, 84, 152, 1039, 615, 1142, 567, 338, 906, 29, 27, 111, 616]. Grundlagen [107]. Grundriss [539]. Gudzinowicz [883]. Guenther [598, 1098]. Guiasu [108]. Guide [634, 883, 371]. Guidebook [408]. Guilford [603]. Gunnar [345, 169]. Gunter [539]. Gupta [913, 1071, 593, 673]. Guttman [427, 216]. Guttmann [1160].

H [359, 212, 917, 1222, 43, 189, 317, 544, 45, 126, 1061, 878, 1318, 436, 787, 55, 537, 613, 528, 82, 291, 374, 567, 654, 229, 338, 883, 30, 597, 433, 714, 461, 1286, 1131, 814, 1028, 24, 187, 1319, 1252, 1037, 196, 1206, 1281, 1097, 820, 333, 28, 367, 1007, 1069, 1220, 341, 1076, 638, 396, 136, 1320, 1031, 1136, 214, 1327, 1070, 1223, 11, 125, 471, 404, 1292, 1230, 721, 48, 406, 648, 211, 540, 1255]. Habakkuk [338]. Haberman [736]. Hadley [112]. Haggett [1200]. Haines [1257]. Haitovsky [774]. Hajek [129]. Haley [1003]. Hall [1109, 816, 1292]. Hamburg [739, 1126]. Hampel [461]. Hanbury [1228]. Handbook [750, 574, 1210, 1031, 109]. Handbuch [947]. Handling [507]. Haney [543]. Hans [940, 540, 960]. Hansen [653]. Hansluwka [1320]. Harary [600]. Harcourt [270]. Harding [635, 599]. Harmonizing [421]. Harold [633, 1210, 812, 28, 806, 705, 531, 91, 1299]. Harper [270]. Harris [737, 912, 42]. Harrison [1209, 893]. Harrod [614]. Harry [508, 55, 322]. Hart [610]. Harter [820]. Hartley [433, 1255]. Harvey [105]. Hastings [750, 575]. Hatton [340]. Hawkins [267]. Hawtrey [861]. Haycocks [367]. Heady [1265]. Heal [809]. Health [339, 67, 482, 594, 587, 889, 585, 1109, 158, 301, 1117, 1177, 46, 717, 740, 25, 53, 579, 527]. Hector [959]. Hedging [225]. Heermann [193]. Heights [700, 10, 353]. Heinz [609, 539, 947]. Held [320, 1170]. Henderson [270]. Henn [1113]. Henri [466, 1038]. Henry [128, 1082, 939, 533, 956, 1102, 469, 897, 1155, 890]. Henschke [924]. Henshaw [999]. Herbert [653, 361, 261, 1156]. Herman [255]. Heron [63]. Herzog [144]. Heterogeneity [1119]. Hey [744]. Heyde [1248]. Heyworth [658]. Hickman [1039, 402]. High [555, 1092]. Higher [68, 271, 764]. Hildebrand [1112]. Hill [965, 286]. Hillier [849]. Hindess [580]. Hine [920]. Hinkley [769, 1249]. Hints [468]. Hinweisen [947]. Historians [573]. Historical [1276, 715, 341, 260, 363, 608]. History [1111, 1108, 191, 1001, 646]. Hobcraft [368]. Hoblyn [560]. Hodder [996]. Hodges [127, 879]. Hodson [363]. Hoel [984]. Hogben [1011]. Hogg [1129]. Hoinville [1199]. Holland [876, 507]. Hollander [562, 1162]. Holmes [710]. Home [144]. Homogeneity [175]. Homogeneous [553]. Honor [1113, 399, 845, 1255]. Honorary [1056, 1189, 1312]. Honour [784, 978, 1123]. Hope [787]. Horton [1284]. Hoskins [657]. Hospital [53]. Household [907, 577, 86]. Housing [173, 275, 480, 894]. Houthakker [404]. Howard [250, 1204, 1112, 250]. Hsiung [1064]. Huang [706]. Huber [461]. Hughes [809, 568, 725, 314]. Human [593, 373, 1018, 567, 953, 478, 1200]. Hundred [101, 208, 354, 496, 631, 767, 908, 1054, 1187, 1311]. Hungarian [1041]. Hunter [1279, 1279, 427]. Huntsberger [1127]. Hypergeometric [1210]. Hypothesis [120, 311, 1260].

I. [1248]. Ian [1068, 996, 410, 376]. Ibern [505]. Ibragimov [1324]. Ideal [972]. Ideas [1030, 366]. Identification [385]. Ihre [539]. II [160, 455, 841, 651, 871, 698, 1111, 364, 605, 402, 1071, 775, 372]. III [498, 731]. Ijiri [1061]. Illustrated [629]. Ilse [141]. Implications [151, 233, 1201]. Importance [1017]. Imports [178]. Improved [1223]. Improvements [207]. Incentive [555]. Incidence [454, 836]. Including [333]. Income [686, 907, 895, 891, 802, 90, 302, 701, 1076, 405, 406, 87]. Incomes [29]. Incomplete [1121]. Independence [1319, 597]. Independent [915]. Index [1236, 972, 1149, 524, 242, 851, 1172, 926]. Indexes [403]. India [1020]. Indian [353]. Indicators [301]. Indices [733, 1207]. Indirect [1119]. Individuals [177]. Inductive [424]. Industrial [234, 607, 138, 611, 1235]. Industries [1275]. Industry [335, 798, 644, 142, 426, 316]. Inequalities [190]. Inequality [618, 1176]. Infectiousness [1276]. Inference [320, 457, 528, 518, 564, 529, 595, 215, 633, 939, 598, 185, 210, 1158, 590, 673, 1127, 283, 1129, 705, 716]. Inferences [922]. Inferring [1182, 1018]. Infinitely [1253]. Inflation [895, 542]. Influence [418]. Influencing [764]. Informatiei [108]. Information [3, 96, 117, 201, 236, 348, 381, 490, 516, 626, 663, 759, 794, 901, 933, 1045, 1180, 1306, 1109, 82, 70, 786, 176, 334, 1087, 594, 674, 1221, 106]. Informationstheorie [107]. Ingram [512]. Initial [891]. Innes [785]. Input [1175, 822, 224, 300, 1041, 822, 1041, 890]. Input-Output [1175, 822, 224, 300, 1041, 822, 1041, 890]. Inspection [869, 1098, 49, 1303, 1214]. Institute [1203, 1062, 131]. Insurance [9]. Integral [855]. Integrals [1210]. Integrated [79]. Integration [435]. Integrodifferential [1197]. Inter [316, 577, 798]. Inter-Departmental [577]. Inter-Industry [798]. Inter-Plant [316]. Interaction [865, 796, 1206]. Interactive [295, 1226]. Interchangeability [1319]. Interdependent [451]. Interdisciplinary [1206, 708]. Interfaces [1327]. International [195, 619, 824, 687, 162, 242, 952, 844, 477, 1009, 952]. Internationale [430]. Interpretation [837, 714, 604, 676].

Interpreting [1120, 843]. Interviewer [408]. Intriligator [439]. Introduction [358, 247, 155, 457, 573, 1034, 296, 913, 993, 180, 879, 318, 914, 638, 974, 1279, 564, 11, 1292, 532, 976, 1162, 254, 1262, 508, 981, 563, 849, 323, 1025, 878, 44, 985, 397, 52, 880, 14, 51, 13, 1285, 854, 609, 103, 1102, 136, 702, 806, 841, 807, 705, 840, 1010, 924, 531, 469, 853, 1257]. Introductory [671, 877, 577, 745, 1103, 1220, 154, 427, 471, 1029]. Invariance [991]. Invariant [278]. Inventory [649]. Inverse [360, 1277, 329, 328, 1310]. Inverses [810]. Investigation [1241, 369, 579]. Investment [895, 337, 203, 173, 275, 164, 267, 27, 1074, 163]. Invited [439]. Involving [990]. Iosifescu [566]. Iran [30]. Iris [755]. Irish [166]. Ironmonger [652]. Irving [997]. Irwin [216, 427, 429]. Isaacson [989]. Isard [300]. Isotonic [528]. Israel [810]. Issues [377]. Itzhak [372]. IV [928, 1137]. Iyer [58].

J [1197, 459, 1036, 246, 889, 680, 1264, 978, 600, 364, 1222, 782, 1109, 981, 317, 544, 195, 948, 821, 56, 19, 849, 712, 505, 1161, 688, 47, 512, 878, 920, 436, 928, 127, 677, 683, 787, 1096, 950, 537, 528, 922, 818, 263, 1125, 1165, 1288, 374, 567, 1290, 338, 883, 30, 1123, 880, 294, 1294, 20, 258, 295, 1316, 1075, 1160, 1217, 823, 461, 777, 748, 750, 879, 220, 803, 1256, 741, 1248, 1253, 1252, 539, 142, 344, 745, 943, 460, 881, 1097, 1159, 1321, 603, 657, 133, 819, 639, 986, 682, 1315, 15, 1322]. J [1220, 574, 1209, 289, 334, 850, 773, 813, 882, 736, 775, 806, 706, 1279, 643, 256, 845, 1040, 705, 1293, 138, 164, 1327, 596, 1326, 815, 1223, 1010, 267, 427, 11, 377, 1059, 404, 531, 16, 104, 1002, 1074, 592, 679, 916, 575, 1107, 270, 1295, 394, 27, 74, 1079, 988]. J.-A [334]. J.-M [881]. Jack [749]. Jacob [533, 399, 921, 1237]. Jacquard [718]. Jacques [501, 919, 721]. Jagdish [1031]. Jagers [946]. Jahn [345, 169]. James [594, 563, 436, 1125, 166, 298, 1112, 377, 144]. Jan [784]. January [589, 726]. Jardine [712]. Jaroslav [129]. Jati [463]. Jean [704, 608, 109]. Jeffrey [424]. Jeffreys [633]. Jenkins [941, 449, 183]. Jennrich [1159]. Jerry [530]. Jerzy [527, 455, 392, 498, 526]. Jessen [1256]. Jessie [755]. Jeux [264]. Joan [431, 1314]. Job [110]. Joginder [650]. Johan [465]. Johansen [435]. John [155, 703, 710, 1004, 744, 678, 512, 436, 227, 913, 368, 751, 196, 816, 1097, 912, 1193, 434, 850, 1191, 1192, 293, 1195, 110, 287, 144, 1276, 1157]. Johnson [75, 499, 895, 1134, 321, 48, 1128, 299, 1278, 1332, 1027]. Joint [158, 589, 726]. Jolliffe [743]. Jon [1005]. Jones [1068, 659, 341]. Josas [430]. Joseph [570]. Joslin [823]. Journals [762, 866]. Jouy [430]. Jouy-en-Josas [430]. Jowell [1199]. Juan [140]. Judge [1201]. Judith [634, 682]. Julius [1329]. June [1170]. Juster [1266]. Juvancz [160].

Kabe [673]. Kac [738]. Kagan [778]. Kahn [944]. Kai [770]. Kain [512, 436]. Kali [1236, 851]. Kallenberg [1259]. Kalton [369]. Kan [1327]. Kapadia [1031]. Karl [923, 103]. Karol [782]. Kaser [161]. Kasius [717]. Kathleen [112]. Katz [27]. Kaufman [1069]. Keeney [1070, 1204]. Keersten [1253]. Kellerer [960]. Kempthorne [319]. Kendall [977, 1111, 1095, 599, 474, 411, 363, 153, 450, 1139, 635, 217, 591, 191, 356, 245]. Kenneth [1125, 1065, 164, 840]. Kenya [540]. Key [480]. Keyfitz [407, 1233]. Keynes [895, 1296]. Kilpatrick [594]. Kindahl [138]. Kinetics [374, 1294]. Kingaby [405]. Kingdom [928, 803, 406, 152, 196, 493, 122, 1110]. Kingman [394]. Kingsley [651]. Kleijnen [749]. Klein [377]. Klevmarken [644]. Klugh [128, 956]. Kohout [680]. Kolchin [1254]. Koller [1320]. Kolman [465]. Konferenz [160]. Konijn [814]. Kossach [924]. Kotowitz [434]. Kotz [75, 499, 715, 1134, 411, 321]. Koutsoyiannis [617]. Kraft [11]. Kramer [46]. Krarup [679]. Krech [879]. Kres [947]. Kreyszig [154]. Krickeberg [211]. Krishna [58]. Krishnaiah [1137, 1280, 1170]. Krotki [782]. Krutchkoff [185]. Kshirsagar [425]. Ku [45]. Kuebler [1028]. Kumar [650]. Kurz [164]. Kuznets [479]. Kyburg [939].

L [527, 247, 215, 457, 989, 1111, 43, 360, 373, 821, 605, 1318, 928, 127, 811, 57, 75, 570, 499, 1290, 229, 910, 955, 1291, 1168, 295, 1075, 748, 879, 687, 1263, 455, 715, 1252, 193, 218, 918, 392, 324, 675, 1281, 1133, 1159, 820, 707, 1134, 672, 1284, 434, 1027, 1102, 506, 772, 244, 222, 426, 1138, 223, 498, 645, 840, 526, 1070, 1204, 469, 194, 321, 887, 48, 257]. L. [344]. L.S.E [112]. Lab [879]. Laboratory [883]. Labour [281, 895, 838, 260, 952, 1310]. Labys [225]. Lachenbruch [846]. Lag [872]. Lagged [2]. Lags [90]. Lai [770]. Laing [1112]. Lainiotis [856]. Lamperti [1195]. Lancaster [24, 638, 125]. Lancelot [1011]. Land [650, 436, 656]. Langford [300]. Language [80, 293]. Lapin [1202]. Large [1150, 1304, 1291, 239, 1151, 606]. Large-Scale [239]. Larson [950, 1290, 742, 806, 705, 531]. Lass [322]. László [525]. Latent [1107]. Latin [753]. Law [1224]. Lawley [332]. Lawrence [565, 1202, 377, 195]. Leabo [429]. Leak [959]. Learning [680, 293]. Least [904]. Leavenworth [645]. Lecture [942]. Lectures [358]. Lee [291, 258, 1110]. Left [268]. Legislation [1271]. Lehmann [127, 955]. Leicester [895]. Leif [435]. Leisure [928]. l’Enfant [340]. Length [1020, 520]. Lennard [844]. Lenstra [1327]. Leo [12]. Leonard [845, 294]. Leontief [1174]. Leroy [319]. Lesley [1109]. Leslie [534]. L’Esperance [223]. Less [904]. Let [289]. Letters [525]. Level [844, 137].

Levels [158, 1009]. Levy [140, 558]. Lewellen [302]. Lewes [928]. Lewis [748, 1231, 1218, 112, 500]. Li [1064]. Library [422, 34, 64, 94, 115, 148, 170, 199, 232, 274, 306, 346, 414, 445, 488, 514, 547, 582, 624, 661, 692, 724, 757, 791, 829, 864, 899, 931, 962, 1013, 1043]. Lickorish [928]. Lie [465]. Lieberman [849, 252]. Liesner [895]. Life [60, 9, 838, 987, 1020, 847, 1314, 1320]. Likelihood [1138, 1259]. Limit [800, 954]. Limits [255]. Lindgren [982]. Lindley [284]. Lindman [812]. Lindsey [922]. Linéaire [264]. Linear [703, 809, 729, 14, 51, 799, 1301, 133, 312, 1033, 936, 1284, 882, 1223, 1260, 848, 916, 630, 1032, 1092, 809, 248, 529, 1060]. Lingstone [1007]. LINK [1075]. Linked [375]. Linking [351]. Linnik [778, 1058]. Listening [817]. Literature [554]. Living [765, 851, 1236, 745]. Local [1243, 173, 275]. Locality [480]. Location [747, 461]. Locational [1200, 1200, 1200]. Loci [375]. Lock [1000]. Locus [1141]. Loeffel [1037]. Loftsgaarden [1131]. Logic [335, 1077, 424]. Logical [939]. Logit [464]. Lohnes [393]. London [1175, 643, 554, 405]. Long [435, 481, 1320, 1088, 1322]. Long-Term [1088]. Longitudinal [1308, 1121, 229]. Loo [888]. Look [179, 725, 289]. Lord [658]. Loren [502]. Lorie [477]. Lothar [17]. Lotte [340]. Louvish [853]. Loveday [572]. Lowes [441]. Lubomir [59]. Lucas [126]. Lucein [498]. Lucien [527, 455, 392, 526]. Lucy [431]. Ludwig [852]. Lukacs [41]. Lung [585, 1182, 1093]. Lutte [340]. Lyman [742, 257]. Lynette [261]. Lynx [1047, 1048, 1244].

M [1197, 527, 889, 401, 215, 1111, 1109, 544, 945, 483, 737, 951, 1318, 928, 137, 1039, 634, 1201, 217, 528, 400, 591, 885, 374, 296, 578, 944, 224, 1065, 1123, 913, 738, 1168, 294, 1294, 425, 1217, 823, 1166, 1258, 987, 1224, 183, 455, 877, 1219, 392, 914, 1163, 881, 994, 1097, 1159, 1162, 1287, 657, 999, 1007, 1069, 1193, 396, 1261, 530, 778, 995, 411, 191, 876, 637, 983, 1032, 498, 1130, 1259, 1265, 18, 328, 526, 405, 377, 105, 194, 1262, 1295, 27, 262, 566, 648, 356, 111]. M. [1079]. M.C. [304]. Macdonald [1294]. MacFarlane [712]. Machi [671, 877]. Machlup [477]. MacKay [1173]. Mackenzie [1047]. MacLagan [789]. Macro [435, 1298]. Macroeconomic [905]. Madansky [1073]. Made [987]. Madge [60]. Madsen [989]. Mahlon [436]. Main [838]. Maistrov [715]. Making [747, 383, 1126, 294, 707, 1228, 571, 284]. Makridakis [1226, 1101]. Male [798]. Malozemov [853]. Man [696, 438]. Management [391, 1034, 683, 78, 296, 913, 741, 163, 203, 707, 682, 575, 1101, 302, 612, 441, 1264, 912]. Manager [119, 966]. Managerial [105, 1227]. Managing [1296, 1004]. Mandel [1320]. Mann [847]. Manpower [313, 1264, 1283, 553, 316, 228, 1315, 1092, 683, 748, 1066]. Manual [368, 640, 642, 1320]. Manufacturing [269, 1275, 27]. many [77]. Mappings [136]. March [320]. Mardia [213, 428]. Margerison [1161]. Marginal [175]. Market [1077, 71, 336, 256, 334]. Marketing [712]. Markets [670]. Markov [459, 1165, 458, 250, 1319, 989, 1222]. Markovian [351, 1092, 212]. Marks [782]. Marriage [953]. Marriott [714]. Marsal [23]. Marschak [1237, 399]. Marshall [1156]. Martin [883, 13, 682, 540]. Martingales [1319, 919, 1319, 501]. Maruyama [1198]. Mary [1328]. Mason [681]. Masoretic [205]. Massey [226]. Material [1234]. Materials [533, 1267]. Materials/ [1267]. Mathai [1224, 1130]. Mathematical [446, 527, 1203, 508, 678, 374, 52, 1217, 13, 455, 752, 392, 423, 820, 1062, 819, 154, 857, 1261, 778, 974, 498, 840, 526, 194, 1230, 656, 776, 1036, 858, 189, 1113, 1233, 567, 1289, 1291, 777, 287, 916]. Mathematically [1004]. Mathematics [537, 655, 657, 572, 218, 132, 1283, 682, 221, 363]. Mathématiques [738]. Mathematischen [107]. Mather [994]. Matheron [855]. Matrices [1150, 1304, 224, 329, 328, 360, 720]. Matter [35, 65, 95, 116, 4, 36, 66, 97, 149, 171, 200, 235, 118, 150, 172, 202, 276, 307, 347, 380, 237, 277, 308, 349, 415, 447, 489, 515, 382, 416, 448, 491, 549, 583, 625, 662, 517, 550, 584, 627, 693, 727, 758, 793, 664, 694, 728, 760, 830, 867, 900, 932, 795, 831, 868, 902, 963, 1014, 1044, 1085, 934, 964, 1015, 1046, 1115, 1145, 1179, 1212, 1086, 1116, 1146, 1181, 1239, 1269, 1305, 1331, 1213, 1240, 1270, 1307]. Matthes [1253]. Matthews [293]. Maunder [928, 803, 1281]. Maurice [977, 1139, 1095, 474, 245, 1082, 54, 1080, 897]. Maximum [1138]. Maxwell [332, 1023]. Mayes [1029, 1029, 1106]. McCammon [884]. McCullagh [850]. McElroy [253]. McGhee [859]. McGuire [399]. McLean [772]. McMillan [376, 410]. McNicol [472]. Mean [836]. Means [1297, 1091]. Measles [454]. Measure [1285]. Measurement [45, 268, 907, 1276, 384, 279, 420, 1263, 1038, 385, 403, 607]. Measures [124, 243, 697]. Measuring [595, 1166, 1176, 110]. Mechanics [690]. Mechanisms [271]. Mecke [1253]. Medgyessy [1140]. Medical [286, 482, 1021, 801, 1164, 1103, 220, 271, 638, 911, 285]. Medicine [1109, 566, 1036, 358, 777]. Medium [1175]. Medium-Term [1175]. Meenan [166]. Meeting [100, 495, 792, 839, 1053, 1186, 775, 101, 208, 354, 496, 631, 767, 908, 1054, 1187, 1311, 158]. Meetings [439]. Megalithic [797]. Melbourne [270]. Melvin [574]. Memorial [957]. Memory [635, 823, 599, 938]. Men [333, 1093]. Mendel [530]. Mendenhall [742, 741, 257, 323, 14]. Menges [539]. Mental [46]. Merrill [103]. Mervyn [957].

Merwe [754]. Method

[1316, 451, 1007, 239, 181, 332, 356]. Methoden [23, 675]. Méthodes [505, 22, 134]. Methodology [492, 385]. Methods [459, 776, 832, 585, 573, 189, 205, 407, 644, 801, 533, 1101, 299, 570, 400, 403, 955, 141, 1050, 994, 333, 369, 1193, 249, 396, 309, 426, 706, 1246, 998, 1136, 929, 285, 463, 1262, 503, 1200, 997, 317, 1226, 126, 1024, 54, 1128, 217, 1290, 913, 465, 562, 1205, 408, 679, 339, 1135, 847, 958]. Michael [161, 484, 140, 483, 1171, 337, 842, 682, 850, 579, 293, 439, 509, 1002, 26, 266, 510]. Michaely [266]. Micro [435, 1298]. Micro-to-macro [1298]. Microeconomics [571]. Middlebrooks [988]. Migration [872, 481, 615]. Mike [409]. Miles [1321]. Millham [221]. Mills [51]. Milner [657]. Milnes [268]. Milton [895, 227, 1059, 1296]. Mine [212]. Minimax [991, 853]. Mining [1050]. Minium [219]. Mintz [141]. Missing [668, 800]. Mitchell [341, 1265]. Mitra [329]. Mitrinović [190]. Mixed [1244, 732, 1121]. Moakes [537]. Mobility [798, 477]. Model [38, 593, 281, 1298, 1039, 732, 282, 451, 834, 1284, 434, 1279, 1040, 1265, 351, 1260, 938, 1092, 26, 1310, 630, 512, 1033, 81, 1230]. Modèles [505]. Modelling [837, 935, 707, 905, 1063, 783]. Models [1197, 1036, 1231, 729, 436, 696, 151, 233, 872, 299, 6, 990, 121, 567, 1308, 402, 1064, 188, 1075, 14, 777, 923, 226, 156, 906, 1134, 312, 936, 781, 1167, 1246, 1229, 292, 576, 592, 595, 822, 703, 47, 748, 250, 854, 999, 882, 248, 1223, 1107, 503, 1200, 228, 1196]. Modern [739, 1202, 612, 1134, 136, 441, 1069, 223]. Moeschlberger [1318]. Moeschlin [1113]. Moment [734]. Monetary [688, 377, 578, 56]. Monitoring [1109]. Monnaie [264]. Monte [239]. Montgomery [226]. Moore [467, 1071, 77]. Morales [334]. Morbidity [821]. Mordecai [164]. Morgenstern [256, 1113].

Morishima [578, 111]. Morphometrics [362]. Morrell [1252]. Morris [739, 1126, 1228]. Morrison [1024, 226]. Mortality [1232, 158, 821, 1243, 1119, 453, 367, 1320, 835, 836]. Mortgage [144]. Morton [46]. Moser [369]. Mosteller [1192, 1100]. Motion [395]. Motives [586]. Mould [1103]. Movement [477, 1002]. Multi [28, 881]. Multi-Dimensionelles [881]. Multi-Product [28]. Multidimensional [175]. Multifactorial [696]. Multiple [921, 714, 1070, 1034, 1297]. Multiplicative [732, 503]. Multiresponse [258]. Multiserver [5]. Multivariate [832, 870, 1226, 499, 8, 1135, 993, 994, 876, 1250, 1136, 1024, 1158, 425, 393, 673, 1023, 706, 1137, 1136, 362]. Multivariaten [947]. Munby [1281]. Murdoch [1293]. Murray [735, 803]. Murthy [1164]. My [60]. Myers [218, 110, 1124]. Myles [562].

N [1172, 365, 600, 317, 366, 1173, 888, 1233, 468, 75, 619, 824, 1077, 975, 258, 534, 332, 777, 750, 133, 707, 506, 560, 882, 810, 1130, 1110, 991, 440, 853, 575, 690, 48]. Nagylaki [1141]. Namias [109]. Nancy [753, 847]. Nanda [434]. Nandi [536]. Narayan [1158, 702, 841]. Nathan [407]. National [1022, 162, 242, 478, 701, 1076, 542, 648, 587, 54, 1177, 406]. Natural [1108, 1166]. Nature [57, 552]. NBER [512]. Negative [1120, 720]. Neil [676]. Neill [537]. Nekrasov [317]. Nelder [227]. Neo [526]. Neo-Darwinian [526]. Neter [703]. Network [1231]. Neubearbeitete [324]. Neuer [324]. Neumann [654]. Neurons [1167]. Neutronics [239]. Neveu [501, 919]. Newell [606, 327]. Newson [293]. Ney [886, 504]. Neyman [527, 455, 392, 498, 526]. Nicholson [484, 509]. Niehaus [1264]. Nielsen [1221]. Nievergelt [1037]. Nigeria [838]. Nigerian [1153]. Night [761]. Nikaido [136]. Ninth [496]. No [891]. Nobay [1295]. Noether [11, 397]. Noise [949]. Noisy [730]. Non [709, 432, 9, 1032, 526, 1092, 630, 720]. Non-Darwinian [526]. Non-Life [9]. Non-Linear [630, 1032, 1092]. Non-Negative [720]. Non-Numerical [432]. Non-Parametric [709]. Nonlinear [288, 848, 50, 400]. Nonparametric [129, 11, 215, 1132, 562]. Nonparametrics. [955]. Norbert [218]. Nordhaus [1235]. Normal [903, 1224, 58, 883, 975]. Norman [889, 499, 368, 1134, 321]. Norwegian [269]. Norwich [822]. Note [313, 417, 1089, 907, 1184]. Notes [31, 62, 89, 113, 145, 167, 197, 230, 272, 303, 342, 378, 412, 443, 485, 446, 545, 620, 942, 300]. Nozer [847]. Null [120]. Number [972, 1149, 524, 462]. Numbers [926, 242, 851]. Numeracy [1016]. Numerate [390]. Numerical [432, 459, 537, 1262]. Nunally [880]. Nunen [1165].

O

[646, 1034, 1113, 610, 433, 892, 1131, 24, 1221, 854, 941, 638, 1027, 1138, 1265, 125, 1255]. O.B.E. [560]. O.R. [613]. Objective [561]. Objectives [436, 1070, 1204]. Observations [314, 1135, 714, 774]. Observed [241]. Occupational [835, 1232]. Occupations [787]. Occurrence [261]. O’Connor [890]. Odeh [1133]. Odell [360, 811]. O’Donnell [544]. Office [577, 1208, 952]. Official [98, 207, 493, 580]. Ogawa [813]. Oil [929]. Okum [377]. Old [179, 725, 205]. Oligopoly [484]. Olive [704]. Oliver [748]. One [101, 208, 354, 496, 631, 767, 908, 1054, 1187, 1311, 904, 920, 1141]. One- [1141]. Ontario [320]. Open [709, 708]. Operational [317, 195, 1117, 613]. Operations [849, 888, 913, 1327, 372, 105, 1003]. Opinion [1309]. Opportunities [391]. Optima [799, 1301]. Optimal [917, 730, 361, 28, 164]. Optimality [1259, 721]. Optimisation [679]. Optimization [1262, 958, 288]. Optimum [1227]. Oral [339]. Ord [677, 460]. Order [528, 1092, 359]. Ordering [870]. Ordinate [214]. Ordinated [680]. Organisation [747]. Organismes [430]. Organization [1264, 553, 399, 158, 301, 25, 53, 436]. Organizations [803]. Oriented [222]. Orton [996]. Osaki [212]. Oscar [319]. Oskar [1113, 256]. O’Sullivan [483]. Oswald [622]. Other [178, 124, 1067, 367, 390, 351]. Otis [576]. Otnes [502]. Ott [742, 257]. Ottestad [188]. Outliers [1218]. Outline [687]. Output [1175, 822, 224, 300, 1041, 1076, 1040, 406, 822, 1041, 890]. Overs [336]. Owen [1133, 426, 820, 1001, 1031, 929, 375]. Ownership [1089, 302, 1088]. Oxton [658].

P [459, 1036, 320, 317, 544, 504, 56, 335, 976, 747, 1035, 22, 134, 749, 944, 569, 954, 1132, 1177, 886, 1294, 188, 892, 846, 1105, 461, 673, 946, 1254, 1127, 803, 987, 1104, 156, 183, 752, 788, 881, 984, 994, 603, 28, 986, 911, 58, 536, 1137, 1279, 1280, 643, 1040, 165, 1327, 1130, 363, 1110, 1170, 1059, 77, 1063, 532, 1074, 285, 1230, 144, 566, 1200]. Paelinck [317]. Page [398]. Pal [1140]. Paper [153, 450, 244]. Papers [446, 1274, 978, 1123, 456, 717, 439, 1255, 245, 192]. Parameter [904, 530, 1125, 919]. Parametric [709]. Parker [928]. Parkin [1229]. Part [1290, 702, 841, 666, 698, 731]. Parthasarathy [331, 1285]. Participant [293]. Particular [585]. Parts [566]. Pascal [654]. Passenger [143, 1281]. Patel [1031]. Patient [587]. Patil [156]. Patricia [921]. Patrick [360, 483, 811]. Pattern [508, 995, 610]. Patterns [266, 643]. Paul [808, 1065, 393, 476, 876, 781, 377, 558]. Pauline [858]. Paulus [374]. Pay [557]. Payments [266, 1008]. Peacock [750, 894]. Pearce [1040, 1067, 265, 267]. Pearson [433, 191, 695]. Pederzoli [1224]. Peltier [81]. People [270]. Percentage [1247]. Perceptions [670]. Percy [722, 475]. Perdrizet [340]. Perennial [1067]. Performance [316, 8, 670]. Pericope [315]. Periodicals [33, 61, 93, 114, 146, 168, 198, 231, 273, 305, 343, 379, 413, 444, 487, 513, 546, 581, 623, 660, 691, 723, 756, 790, 828, 863, 898, 930, 961, 1012, 1042, 1114, 1143, 1178, 1211, 1238, 1268, 1300, 1330]. Perks [147]. Person [331]. Personal [665, 628, 803, 1209, 405]. Personnel [834, 53]. Persons [686]. Perspective [1271]. Perspectives [629, 1123]. Peter [467, 610, 288, 142, 322, 426, 500, 1006]. Peto [1108]. Petriccione [1291]. Petrov [915]. Pharmacology [605]. Phasing [28]. Phenomena [394]. Philadelphia [300]. Philip [859]. Phillips [565]. Physical [1291, 994]. Physique [738]. Piccioni [1291]. Pierre [992]. Pilling [1081]. Piper [568]. Pitman [978]. Plackett [247, 1111, 672]. Planes [568]. Planning [784, 1264, 1109, 301, 670, 642, 748, 1227, 611, 408, 1066, 889, 1283, 263, 409, 131, 1315, 1225, 132]. Plant [316]. Plants [1151, 1067]. Platelet [374]. Plusieurs [992]. Pocket [1133]. Point [1164, 1253, 918, 451, 500]. Points [1247]. Policies [7, 271, 1265, 29, 266, 165]. Policy [268, 377, 1282, 1100, 1174, 164]. Political [422, 268]. Politics [859]. Pollard [567, 471]. Polls [1309]. Pollution [527, 678]. Polymorphisms [375]. Pontoon [280]. Population [1197, 1231, 852, 40, 629, 1064, 1198, 492, 577, 588, 494, 1246, 83, 407, 782, 650, 84, 1169, 338, 943]. Populations [373, 567, 718, 1166]. Port [886]. Portfolio [163]. Poser [1297]. Positivism [580]. Post [387, 1215]. Post-Experience [387]. Postal [408]. Postwar [141, 29, 266]. Poverty [907, 270]. Powell [405, 656]. Power [1017]. Practical [1149, 218, 945, 468]. Practice [19, 1252, 926, 876, 1199]. Practitioners [750]. Prais [404]. Pratt [678, 657]. Pre [1201, 176, 534]. Pre-1968 [534]. Pre-School [176]. Pre-Test [1201]. Precision [45]. Predict [453]. Predictability [256]. Predicting [522]. Prediction [629, 729, 1096, 333, 773, 333, 1223, 1112, 503]. Preferences [1070]. Prekopa [649]. Preliminary [585, 385]. Preparation [446]. Prepared [320, 271]. Presentation [1017, 439]. President [1055, 1188]. Pressat [684]. Prevention [261]. Price [1172, 137, 225, 1149, 179, 90, 725, 1323, 137, 403]. Prices [1040, 949, 256, 138]. Pricing [543, 1235]. Primarily [655]. Primer [472]. Principle [1267]. Principles [594, 19, 674, 441, 286, 1290, 77, 105]. Pringle [328]. Prior [176]. Priorities [1022]. Prithwis [593]. Probabilistic [250]. Probabilities [937, 1184]. Probability [527, 978, 1111, 323, 127, 468, 607, 70, 121, 674, 954, 1123, 384, 886, 1131, 1091, 1224, 187, 13, 318, 455, 671, 780, 877, 1285, 392, 424, 1242, 1321, 770, 1134, 940, 1102, 702, 841, 974, 186, 536, 498, 705, 526, 469, 532, 561, 981, 72, 979, 1286, 182, 525, 423, 1317, 191, 702, 1001, 1094, 185, 951, 299, 1124, 130, 715, 1319, 842, 319, 322, 1129, 292, 1130, 1059, 532, 12, 211]. Probation [333]. Probit [246]. Problem [1182, 689, 552]. Problèmes [738]. Problems [39, 858, 140, 449, 330, 518, 1148, 1198, 69, 15, 479, 778, 239, 689, 37, 252, 571, 678, 290, 1293, 988, 1058, 1249, 848]. Procedure [875]. Procedures [44, 297, 1258, 278, 45]. Proceedings [101, 100, 208, 354, 496, 495, 631, 767, 792, 908, 839, 1054, 1053, 1187, 1186, 1311, 1175, 320, 822, 1041, 1170, 844, 527, 195, 455, 392, 498, 526]. Process [1164]. Processes [1168, 946, 1319, 500, 566, 12, 250, 212, 504, 1165, 771, 735, 1253, 918, 1196, 770, 1195, 1324, 592]. Processing [786, 1297]. Product [28]. Production [269, 895, 687, 426, 122, 435, 27]. Professions [390]. Professor [1274, 486].

Profits [1077]. Programme [293, 158].

Programmed [973, 920, 254].

Programmes [301]. Programming

[50, 263, 1290, 52, 1289, 1291, 194, 463, 575, 1288, 288, 133, 924, 848, 1230, 656, 1160]. Programs [1159, 431, 656, 297]. Progress [1021, 775]. Prohorov [211]. Project [139, 30, 655, 1075, 764, 1246]. Projecting [822]. Projections [1169]. Projective [568]. Projects [1106]. Promises [140]. Promotion [1092, 630]. Proneness [261, 1151]. Proof [1317]. Properties [5]. Proportional [729, 311]. Proportions [241, 570]. Prose [520]. Prospect [478, 479, 541, 511]. Protection [971, 1271, 1272]. Protein [552]. Prothero [450]. Prototype [512]. Providing [589, 726, 1236]. Proxima [799, 1301]. Psychology [219, 880, 251, 603]. Public [1100, 301, 717, 1309, 511, 164]. Publications [163]. Purchases [6]. Pure [619, 824]. Puri [215]. Purposes [242].

Qualified [316]. Quality

[709, 997, 403, 1098, 475, 645, 49, 79]. Quandt [400, 1032]. Quantitative [573, 646, 913, 258, 479, 439, 1257]. Quarterly [872, 701]. Quelques [738]. Quenouille [1082, 1097, 396]. Questionnaire [642]. Queue [74]. Queueing [776, 737, 942, 73, 327, 990, 398, 1293]. Queues [5, 459].

R [1036, 482, 247, 1231, 809, 743, 989, 1264, 600, 646, 1111, 264, 917, 822, 1113, 544, 1234, 19, 563, 185, 927, 1161, 436, 928, 683, 137, 1103, 1096, 1039, 1128, 1035, 528, 818, 926, 812, 291, 374, 1290, 1135, 674, 569, 143, 771, 975, 20, 258, 329, 456, 393, 1105, 461, 777, 673, 742, 228, 980, 162, 1256, 1124, 1028, 1104, 112, 1285, 640, 1252, 847, 1133, 1159, 1287, 1321, 657, 819, 639, 707, 1220, 672, 1284, 88, 341, 474, 1261, 778, 411, 636, 772, 184, 807, 222, 1137, 1138, 1225, 1280, 998, 958, 1136, 1229, 248, 769]. R [1249, 328, 375, 1070, 363, 405, 1170, 377, 1074, 568, 679, 1066, 690, 270, 894, 1295, 1296, 85, 362, 503, 890]. R. [32, 304, 1314]. Rade [844]. Radhakrishna [529]. Radioisotopic [374]. Radner [399]. Raffle [643]. Raghavan [331]. Raghavarao [330]. Ragnar [486]. Rahman [468]. Raiffa [1070, 1204]. Raj [641]. Ralph [861, 1204]. Rama [885]. Ramachandran [778]. Random [903, 904, 973, 915, 1185, 58, 1324, 910, 1254, 156, 187, 735, 918, 855]. Ranga [975]. Range [1322]. Rank [597, 1163, 278, 217]. Ranks [955]. Rao [600, 329, 999, 778, 975, 529]. Rapid [396]. Rashi [271]. Rate [164]. Rates [570, 1243, 196]. Rathie [1130]. Ratio [1259]. Ratios [1119]. Raveendra [619, 824]. Rawstron [262]. Ray [847, 441]. Rayner [21, 328]. Readings [163, 370, 252, 1069, 193]. Real [950, 578, 987]. Real-Life [987]. Realistic [335, 772]. Reappraisal [481, 30]. Reasoning [219, 601, 1131]. Recherches [430]. Recognition [508]. Recollections [300]. Recommendation [350]. Recommendations [952]. Reconciliation [1242, 122]. Record [875]. Recurrent [873, 1153]. Reddaway [544]. Redpath [405]. Reduction [368, 843]. Reference [585, 464, 432, 426, 891]. Reflections [300]. Reform [786]. Reformulation [936]. Regenerative [394]. Regina [299]. Regional [1275, 196, 1246, 196, 317, 300, 1208, 262]. Regions [216, 88]. Registrar [1232, 892]. Registrar-General [892]. Regression [1034, 729, 920, 921, 733, 123, 1060, 528, 704, 1192, 1323, 774]. Regression/Correlation [921]. Regulations [234, 548]. Reid [1083]. Reinhardt [1131]. Reinmuth [741]. Related [1168, 886, 1071]. Relation [152]. Relationship [390]. Relationships [2, 651, 650]. Relevance [336]. Relevant [987]. Reliability [79, 555, 452, 847, 893]. Reliable [106]. Religion [480]. Remarks [801]. Rene [320, 135]. Renton [783]. Rényi [525, 92, 182]. Repeat [437]. Repeat-Buying [437]. Repeated [120]. Repges [1036]. Reply [1052]. Report [102, 209, 355, 497, 632, 519, 589, 768, 726, 909, 1057, 1056, 1055, 1190, 1189, 1188, 1313, 1312, 158, 301, 1272, 1154, 891, 577, 271, 204, 1022, 544, 80]. Reports [655, 957]. Representing [520]. Reproduction [593]. Requirements [553]. Research [1273, 646, 1022, 195, 436, 753, 1117, 613, 913, 261, 1003, 754, 478, 588, 479, 541, 426, 204, 480, 1199, 1010, 408, 372, 105, 511, 548, 642, 478, 542, 408, 128, 19, 849, 888, 1164, 956, 1258, 162, 1023, 1327, 292, 285, 648, 370]. Resources [478, 1267]. Responsiveness [266]. Ressources [264]. Restrictions [528]. Results [385]. Retrospect [478, 479, 541, 511]. Return [164]. Reuben [1084]. Reutlinger [139]. Revenue [140]. Revesz [954]. Review [1197, 817, 1172, 459, 776, 859, 1207, 1236, 286, 339, 1036, 1175, 246, 358, 482, 527, 594, 889, 686, 128, 247, 858, 1203, 139, 915, 1231, 401, 784, 561, 155, 215, 320, 359, 457, 633, 703, 939, 977, 1094, 593, 809, 466, 601, 680, 743, 989, 1264, 1283, 852, 978, 365, 508, 573, 600, 646, 808, 1111, 264, 212, 364, 917, 1222, 50, 43, 189, 1022, 161, 407, 709, 710, 711, 782, 1100, 1108, 1109, 1232, 822, 650, 360, 435, 981, 1034, 653, 1113, 76, 997, 1068, 484, 317, 544, 644, 450, 140, 84, 158]. Review [366, 598, 442, 361, 504, 45, 195, 1234, 373, 948, 821, 326, 1004, 616, 56, 1226, 744, 678, 481, 19, 21, 126, 563, 335, 533, 1173, 269, 849, 888, 945, 399, 618, 1061, 185, 323, 602, 712, 927, 976, 1025, 1101, 1139, 1169, 860, 467, 330, 505, 610, 1024, 1095, 1161, 688, 47, 483, 512, 747, 635, 737, 951, 739, 878, 605, 1210, 44, 565, 920, 54, 1318, 436, 928, 127, 268, 174, 677, 713, 811, 316, 683, 684, 1233, 753, 787, 78, 942, 137, 57, 301, 79, 1164, 1103, 1298, 468, 1096, 55, 473]. Review [1039, 1325, 80, 634, 950, 615, 1128, 1035, 395, 537, 613, 210, 299, 652, 72, 1126, 75, 219, 1142, 225, 1171, 1201, 129, 217, 1008, 528, 570, 922, 956, 22, 134, 160, 464, 81, 82, 400, 499, 591, 617, 818, 885, 926, 1073, 263, 949, 1125, 619, 824, 403, 990, 607, 812, 985, 996, 1165, 1288, 291, 895, 259, 704, 374, 567, 227, 749, 60, 925, 1290, 654, 397, 190, 229, 296, 410, 786, 921, 1323, 432, 720, 1135, 1158, 297, 402, 578, 674, 944, 979, 338, 224, 1297, 1077, 569, 883, 590, 30, 143, 597, 771]. Review [910, 954, 955, 975, 1064, 1132, 1198, 1065, 1123, 52, 288, 913, 1289, 1291, 1177, 46, 719, 738, 1168, 880, 294, 300, 655, 642, 856, 1267, 891, 472, 886, 1294, 20, 188, 258, 295, 329, 425, 433, 456, 714, 1218, 1316, 1075, 393, 1, 14, 534, 1144, 717, 892, 1160, 1217, 141, 409, 823, 332, 846, 993, 1105, 461, 777, 673, 430, 462, 465, 718, 946, 1141, 1166, 1254, 1286, 368, 51, 1131, 604, 742, 1194, 228, 748, 980, 1000, 1258, 302, 750, 879, 923, 1098, 438, 687, 429, 220, 740, 251, 1127, 1067, 1072, 226, 1041]. Review [162, 803, 376, 814, 1256, 987, 1124, 751, 1028, 1104, 112, 156, 261, 290, 676, 741, 1140, 1224, 24, 182, 183, 187, 458, 525, 1248, 1263, 13, 130, 250, 318, 455, 501, 671, 715, 735, 779, 780, 877, 919, 1058, 1221, 1253, 1285, 1319, 640, 1202, 1252, 193, 218, 562, 752, 59, 854, 609, 1003, 804, 1219, 918, 163, 106, 107, 108, 17, 135, 166, 539, 1037, 1266, 1038, 196, 255, 284, 577, 142, 788, 1206, 216, 283, 392, 423, 424, 608, 1071, 1317, 337, 649, 847, 914, 371, 599, 842, 941, 745, 943, 23, 324, 675, 816, 1281]. Review [754, 460, 1163, 881, 984, 994, 1023, 1097, 1133, 1159, 1162, 1196, 1287, 1321, 953, 820, 1062, 333, 603, 431, 271, 477, 478, 657, 28, 133, 819, 131, 367, 612, 639, 707, 785, 912, 986, 999, 1007, 1069, 1228, 682, 770, 1315, 132, 369, 1134, 15, 685, 1322, 1220, 502, 159, 319, 672, 1033, 1193, 1284, 1009, 88, 249, 476, 86, 87, 103, 154, 341, 479, 574, 1005, 1076, 1176, 1208, 1209, 73, 322, 580, 857, 474, 638, 911, 157, 289, 334, 336, 396, 434, 940, 1027, 1102, 1174, 1251, 1261, 136, 213, 530, 746, 778, 850]. Review [995, 1191, 1192, 411, 957, 25, 53, 192, 506, 636, 772, 773, 813, 882, 221, 58, 475, 265, 184, 191, 702, 736, 775, 806, 841, 876, 974, 1001, 807, 579, 222, 254, 538, 541, 982, 426, 637, 843, 983, 357, 510, 1320, 437, 716, 781, 186, 293, 536, 689, 681, 706, 810, 1026, 1031, 1032, 1129, 1137, 1138, 1167, 1195, 1225, 1227, 1279, 1280, 1324, 606, 223, 500, 535, 498, 327, 1250, 855, 998, 708, 398, 958, 992, 1136, 110, 260, 1229, 643, 480, 256, 439, 564, 845, 1006, 1040, 929, 1199, 214, 705, 645, 1293, 138, 164, 165]. Review [441, 611, 1327, 248, 287, 292, 596, 769, 840, 1130, 1249, 1259, 1265, 1326, 815, 18, 328, 375, 526, 641, 833, 1030, 1070, 1204, 1205, 1223, 363, 298, 405, 1010, 1099, 529, 576, 924, 267, 947, 427, 884, 1110, 1112, 1170, 11, 125, 377, 542, 783, 507, 509, 408, 1059, 470, 471, 404, 851, 531, 774, 16, 77, 83, 104, 469, 543, 614, 647, 1260, 1292, 370, 1002, 991, 1063, 372, 532, 440, 1074, 848, 952, 285, 428, 105, 194, 331, 463, 568, 592, 679, 853, 916, 1066, 1230, 1262, 325, 26, 49, 1235, 321, 144, 1060, 721]. Review [575, 42, 690, 887, 1107, 572, 29, 48, 109, 252, 253, 257, 270, 571, 805, 844, 894, 1029, 1106, 1257, 1282, 1295, 1296, 85, 362, 394, 406, 503, 511, 595, 656, 27, 266, 262, 340, 566, 648, 651, 1200, 893, 12, 41, 74, 211, 356, 540, 1255, 988, 890, 245, 1314, 111]. Reviews [827, 826, 1281, 825, 928, 803, 928, 803, 1281]. Revised [875]. Revisions [701]. Revisited [383]. Revolution [857]. Reyment [362]. Reynolds [1265]. Ricciardi [1168]. Richard [1108, 610, 950, 400, 925, 1132, 717, 879, 429, 59, 1206, 424, 1317, 999, 1032, 645, 884, 257, 571]. Richardson [55]. Right [268]. Riley [862]. Ringstad [269]. Rinnooy [1327]. Rishel [917]. Risk [43, 189, 467, 773]. Risks [1318]. Rita [54].

River [1047]. Riverflow [935]. Road [1019, 30, 261, 1281, 1088]. Robbins [361, 878]. Robert [817, 457, 563, 44, 985, 1267, 1258, 106, 502, 510, 186, 681, 1129, 377, 543, 572, 789]. Roberts [682]. Robson [438]. Robust [461]. Roderick [573]. Roessler [1102, 469]. Roger [81, 1199, 893]. Rogers [461, 746]. Rohatgi [974]. Rohlf [20, 19, 563]. Roland [684]. Role [753, 337]. Rollo [635, 599]. Ronald [250, 270]. Rosander [1099]. Rose [295]. Rosenblatt [735]. Rosenthal [1112]. Roskill [350]. Ross [647, 979]. Rota [600]. Round [844]. Rowley [1074]. Rows [1122]. Roy [1109, 399, 227, 258, 614, 264, 135]. Royal [101, 208, 354, 496, 495, 631, 548, 767, 792, 908, 839, 971, 1054, 1053, 1187, 1186, 1311, 352, 891]. Royce [512, 436]. Rozanov [211, 1324]. Rudiments [1016]. Rudolf [424]. Ruggles [753]. Rugh [916]. Rule [1201]. Run [435]. Runyon [1132, 1105, 1104]. Russian [411, 411]. Ruth [356]. Ruymgaart [597]. Ryzin [878].

S [1172, 1236, 594, 1094, 1283, 600, 808, 212, 782, 1068, 544, 361, 1226, 533, 849, 888, 712, 1101, 951, 928, 811, 57, 79, 210, 652, 75, 1171, 190, 296, 786, 1323, 1297, 1123, 1291, 886, 258, 329, 433, 879, 1067, 814, 1319, 1071, 477, 639, 132, 249, 1027, 995, 58, 191, 843, 1320, 437, 716, 1167, 1227, 1081, 845, 645, 248, 18, 405, 576, 427, 1059, 404, 851, 104, 848, 331, 144, 1107, 656, 340]. S. [892]. Sachs [17, 324, 675]. Salant [477]. Salaries [644]. Sampath [1167]. Sample [814, 130, 799, 1301, 475, 596, 641, 1099, 408, 109]. Samples [1138]. Sampling [1245, 640, 255, 309, 833, 1214, 1255, 869, 1303, 1030, 48, 257, 1098, 49]. Samuel [499, 715, 1134, 411, 321]. Samuelson [377]. Sard [259]. Sargant [335]. Sargent [165]. Sarkadi [775]. Sastry [885]. Sato [136]. Sauvy [1142]. Savage [344, 845]. Sawyer [434]. Scale [269, 239, 787]. Scarf [653].

Scene [610]. Schafer [847]. Scheaffer [257]. Scheerer [130]. Scheffe [1155]. Scheuer [987]. Schmalensee [571]. Schmetterer [840]. Schmidt [1006]. School [38, 655, 554, 176, 241, 844]. Schools [77]. Schultz [953]. Schultze [107]. Science [1264, 1152, 522, 816, 204, 422, 683, 1125, 923, 682]. Science-Based [522]. Sciences [1327, 466, 44, 81, 229, 296, 921, 993, 742, 438, 740, 193, 1284, 293, 1229, 363, 576, 252]. Scientific [607, 1030, 633, 356]. Scientist [1314]. Scientists [680, 657, 565, 51, 1124]. Sclerosis [1297]. Scott [527, 263, 455, 392, 1322, 498, 526, 543]. Scott-Armstrong [1322]. Seal [43]. Search [929]. Searl [248]. Seasonal [449, 732, 123]. Seasonality [927]. Seber [1060]. Second [908, 1041, 572, 341]. Secondary [38]. Secretariat [29]. Section [234, 548]. Sector [1265, 27]. Secular [542]. Security [494]. Seers [409]. Selected [1247, 44, 297, 717, 1069, 192, 109, 1203, 820, 1062, 1261]. Selection [1166, 408, 1141]. Selections [957]. Selective [544]. Sellekaerts [784]. Seltzer [782]. Semi [250]. Semi-Markov [250]. Seminar [42]. Seminars [1109]. Sen [690, 618]. Sendov [777]. Seneta [720, 1248]. Sengupta [463]. Sensitivity [313]. Sentence [520]. Sentence-Length [520]. Septembre [430]. Sequential [911]. Series [1047, 976, 1051, 179, 297, 837, 178, 183, 935, 725, 941, 701, 314, 502, 551, 905, 977, 1025, 1139, 591, 1159, 357, 42]. Serra [1321]. Server [606, 74]. Service [803, 389, 471, 587, 1066]. Services [67, 585, 1117, 410, 628, 376, 25, 803]. Sets [136, 855]. Setting [1022]. Settlements [788]. Sevast’yanov [1254]. Seven [88]. Seventh [208]. Sex [557]. Sforza [373]. Shackle [57]. Shahani [536]. Shanti [1071]. Shape [1263]. Sharing [140]. Sharpe [695]. Shaw [261]. Shelby [736]. Sheldon [979]. Shelemyahu [283]. Shenton [1138]. Shipping [181]. Shiskin [1329]. Shiv [913].

Shlomo [139]. Short [435, 314]. Should [874]. Shrikhande [600]. Shryock [533]. Sichel [261]. Sickness [643]. Siegel [533]. Siegmund [361]. Signal [472]. Significance [701, 120, 709, 602]. Silvey [817, 210]. Silviu [108]. Simha [58]. Simon [479, 1061, 333]. Simple [800]. Simplified [1303, 1214]. Simulation [47, 512, 314, 1265, 749, 1063]. Sinai [315]. since [338, 166]. Singapore [835]. Singer [574]. Singh [888, 336]. Single [1167, 74]. Singpurwalla [847]. Sir [965, 861]. Sirc [687]. Sisteme [108]. Situation [555]. Sixth [101, 527, 468, 455, 392, 498, 526]. Size [1061, 907, 255, 788, 552]. Skala [539]. Sketch [715]. Skew [1061]. Skill [206]. Skipworth [326, 751]. Skorohod [535]. Slater [431]. Slopes [836]. Small [637, 983, 1138, 574]. Smallwood [400]. Smart [220]. Smeed [1084]. Smith [1233, 291, 47, 683, 79, 228, 1028, 671, 877, 1066, 48]. Smoking [585, 1182, 586, 1110, 1093]. Snap [370]. Snedecor [621]. Snowdon [16]. Snyder [918]. Sobel [1261]. Social [466, 680, 1150, 1304, 565, 753, 787, 1117, 81, 925, 229, 296, 642, 993, 51, 742, 438, 628, 803, 196, 816, 369, 685, 1284, 241, 293, 204, 1229, 480, 576, 351, 408, 370, 592, 1106, 262, 786]. Societies [368]. Society [604, 439, 101, 208, 354, 496, 495, 631, 548, 767, 792, 908, 839, 971, 1054, 1053, 1187, 1186, 1311, 131, 352]. Socio [1107]. Socio-Economic [1107]. Sociological [922, 754]. Sociology [580]. Sokal [19, 563, 20]. Sol [259]. Solomon [1286]. Solow [137]. Solutions [1249, 468, 77]. Solve [689, 988]. Solving [239, 689, 15]. Som [640]. Some [153, 450, 1051, 5, 972, 1148, 1308, 179, 492, 725, 452, 906, 701, 385, 929, 37, 1048, 801, 244, 73, 331]. Sophisticated [1077]. Source [218]. Sources [54, 928, 895, 1000, 803, 1281, 928, 685, 538]. Soviet [857, 161]. Space [1325]. Spaces [130]. Spatial [483, 263, 1196, 746, 850, 995, 677, 996]. Special [100, 495, 792, 839, 1053, 1186, 644, 464, 432, 426]. Specific [836]. Specification [451]. Spectral [42, 314, 249]. Spectrum [1244]. Speculation [225]. Spencer [993, 819]. Spieltheorie [1037]. Spiezer [1108]. Spike [1167]. Splines [259]. Spoilt [180]. Sponsored [577]. Sprott [320]. Squared [125]. Sreedharan [28]. Srinivasan [1167]. Srivastava [600, 258, 882]. Staff [643, 630]. Staffing [389]. Stage [834]. Staging [28]. Standard [869, 1091]. Standardized [1119]. Standing [891]. Stanley [366]. Stat [879]. State [1170]. States [140, 821]. Statics [802]. Stationary [7]. Statistic [1278, 1332]. Statistical [39, 101, 208, 354, 496, 495, 631, 548, 767, 792, 762, 866, 908, 839, 971, 1054, 1053, 1187, 1186, 1311, 969, 155, 215, 320, 703, 939, 601, 205, 315, 1022, 711, 1232, 598, 801, 45, 1047, 185, 1025, 1024, 967, 928, 299, 1201, 786, 1135, 1158, 883, 590, 955, 1218, 892, 332, 673, 1131, 980, 1098, 1127, 554, 421, 803, 24, 1248, 1221, 562, 59, 577, 283, 452, 847, 1281, 1133, 389, 574, 1208, 474, 396, 995, 411, 882, 521, 843, 357, 706, 1031, 1129, 500, 1136, 929, 705, 37, 529, 11, 556, 77, 352]. Statistical [1002, 991, 1153, 648, 356, 988, 1087, 594, 466, 997, 644, 678, 126, 1161, 301, 1096, 1128, 210, 1126, 219, 528, 570, 291, 895, 227, 749, 20, 188, 1316, 750, 923, 251, 814, 1256, 218, 216, 1071, 707, 1228, 1315, 1076, 746, 813, 982, 426, 681, 1026, 645, 285, 690, 572, 988, 245]. Statistician [422, 119, 388, 1147, 1271, 60]. Statisticians [98, 391, 387, 430, 390, 775, 433, 751]. Statistics [98, 589, 726, 1273, 358, 527, 1203, 238, 320, 977, 1094, 743, 978, 1111, 709, 1021, 207, 981, 629, 366, 1247, 19, 712, 951, 739, 44, 565, 920, 874, 54, 78, 665, 468, 1272, 766, 418, 75, 1008, 1183, 499, 607, 704, 70, 60, 397, 1132, 1123, 880, 1152, 9, 628, 1124, 1104, 290, 455, 193, 69, 392, 745, 1017, 820, 1062, 603, 419, 639, 493, 319, 580, 1261, 957, 53, 221, 191, 702, 775, 841, 1001, 807, 222, 254, 293, 223, 498, 204, 260, 845, 292, 526, 924, 1170, 471, 181, 99, 321, 572, 252, 253]. Statistics [1296, 383, 1255, 1203, 1062, 244, 11, 286, 247, 359, 365, 808, 710, 948, 326, 1173, 323, 860, 1095, 878, 127, 301, 1103, 473, 129, 985, 925, 46, 1105, 1000, 879, 429, 220, 779, 1058, 1219, 392, 1163, 984, 367, 785, 15, 685, 1220, 103, 154, 341, 322, 638, 1102, 1251, 778, 806, 841, 974, 538, 1280, 287, 769, 840, 1130, 1249, 328, 427, 11, 470, 531, 16, 469, 1292, 952, 325, 844, 1029, 1106, 859, 128, 680, 364, 1100, 76, 744, 634, 950, 956, 294, 604, 742, 987, 1028, 741, 1202, 1252, 1206, 1162]. Statistics [986, 159, 1027, 25, 636, 637, 983, 536, 1279, 708, 1110, 428, 844, 893]. Statistik [539]. Statistique [992, 738]. Statistiques [22, 134]. Statistische [17, 23, 324, 675, 947]. Stein [1201]. Stein-Rule [1201]. Stephen [400, 1194, 1251, 876, 1032]. Stewart [1005]. Stigler [138]. Stochastic [593, 917, 282, 935, 770, 1227, 606, 12, 852, 43, 635, 771, 1064, 1198, 226, 599, 1167, 1195, 500, 535, 463, 592, 887, 566, 41]. Stock [1275, 371, 336, 256, 1077]. Stocks [722]. Stone [1122, 925, 788]. Stones [1122]. Stoodley [711]. Stopping [361]. Storage [73, 649]. Straszheim [436]. Strategies [612]. Stratification [1215]. Stratified [309]. Street [1077]. Streeten [476]. Strike [282, 1151]. Strike-Proneness [1151]. Stromback [1040]. Stroup [950]. Structural [1164, 854, 334, 576]. Structure [335, 1325, 718, 788, 586, 165, 1174]. Structures [1321, 156]. Struggles [966]. Stuart [1279, 427, 245, 977, 1095, 1030]. Student [8, 131]. Students [750, 293, 21, 745, 112]. Studies [699, 871, 403, 229, 1099, 262, 1121, 78, 476, 978, 1111, 885, 619, 424, 191, 1032, 845]. Study [969, 205, 315, 269, 449, 316, 615, 1276, 8, 300, 522, 333, 314, 88, 334, 244, 1093, 99, 29, 27, 540]. Stylistics [59]. Subject [68]. Subjective [384, 555, 1091]. Success [57, 462]. Successful [350]. Sugden [544]. Suggestions [436]. Summarising [920]. Sums [915]. Sunspot [1049]. Superpositions [1140]. Supplement [1232, 892]. Surfaces [719]. Surrey [401]. Surrogate [588]. Survey [600, 240, 1047, 922, 1177, 461, 465, 814, 1256, 577, 882, 833, 298, 408, 48, 257, 1255, 811, 270, 369, 1199]. Surveys [339, 109, 577, 641, 405]. Susser [957]. Swan [896]. Sweden [1271]. Swedish [644, 1298]. Switching [904]. Sydenstricker [717, 717]. Sydney [885]. Symmetric [808]. Symposium [527, 320, 455, 392, 498, 526, 1170]. System [38, 688, 967, 925, 242, 1174, 241, 680, 543]. Système [264]. Systems [268, 786, 875, 1177, 250, 914, 451, 669, 606, 611, 1092, 942, 1291, 1141, 228, 1072, 73, 916, 889].

T [459, 889, 719, 258, 1218, 534, 777, 1141, 1166, 604, 106, 953, 1220, 1027, 357, 510, 1040, 331, 1282]. T. [560, 1081]. Table [175, 668, 800, 838, 1320, 844, 569]. Tables [1203, 433, 255, 820, 1062, 1076, 157, 1261, 20, 1133, 751, 58]. Tabulated [1326]. Tafeln [947]. Tak [30]. Take [336]. Take-Overs [336]. Takeo [1198]. Tanis [1129]. Tanur [634]. Tapiero [1227]. Tarling [809]. Tarshis [477]. Tautu [1036, 363, 566]. Tax [544, 268]. Taxation [511]. Taylor [163]. Teach [874]. Teacher [557]. Teaching [589, 726, 766, 418, 419, 844]. Team [709, 708]. Technical [589, 726, 468]. Techniques [215, 508, 749, 432, 993, 368, 1041, 1007, 1315, 530, 681, 1250, 556, 77, 711, 1041, 1256, 640, 929, 194, 139]. Technology [162, 76]. Teekens [503]. Teicher [1319]. Teil [539]. Templeton [459]. Temps [501]. Ten [655, 475]. Teoriei [108]. Terje [653]. Term [1088, 1175]. Terminology [524, 754]. Terms [268, 474, 411]. Terry [1004]. Tessa [112]. Test [920, 1201, 278]. Testament [205]. Testing [709, 396, 1260]. Tests [920, 597, 120, 596, 1259, 11, 602, 991]. Text [205, 920, 745, 222]. Theil [466, 82, 1038]. Their [88, 336, 188, 141, 1134, 25]. Thelma [895]. Theorems [954]. Theoretical [762, 866, 579, 1249, 366, 769]. Theorie [22, 134, 264]. Theories [852]. Theory [247, 561, 977, 1094, 989, 43, 240, 1113, 442, 361, 1234, 1095, 920, 1318, 528, 1276, 926, 619, 824, 1148, 944, 597, 288, 462, 814, 1248, 318, 715, 1319, 1252, 106, 283, 392, 1071, 770, 336, 813, 806, 876, 974, 437, 781, 689, 810, 500, 398, 705, 1293, 596, 1059, 531, 532, 887, 776, 155, 600, 1222, 189, 737, 82, 954, 1289, 856, 472, 13, 455, 1221, 1134, 982, 186, 498, 327, 815, 690, 211, 1207, 617, 578, 671, 877, 1038, 1033, 1063, 85, 111]. Thesaurus [754]. Thias [540]. Thinking [914]. Third [844, 1054, 430]. Thirty [101, 208, 354, 496]. Thirty-Eighth [354]. Thirty-Ninth [496]. Thirty-Seventh [208]. Thirty-Sixth [101]. Thomas [360, 300, 1266, 810, 1265, 894, 1296, 481, 615, 1069, 88]. Thompson [889, 72]. Thorndike [1258]. Three [920, 892]. Threshold [781]. Throughput [822]. Thwaites [655]. Tijms [1222]. Tilford [377]. Time [977, 976, 1025, 1051, 951, 297, 837, 178, 143, 935, 701, 502, 551, 357, 905, 42, 1139, 591, 183, 941]. Time-Series [1051, 178]. Tinbergen [784]. Tinker [1108]. Tintner [221, 885]. Tjur [780, 423]. Today [541]. Tolerance [216]. Too [762, 866]. Tool [742]. Tools [612]. Top [391, 612]. Topics [1168, 1071, 331, 886]. Torrens [505]. Torrens-Ibern [505]. Torres [140]. Tourism [928, 927]. Toutenburg [1223]. Towns [88]. Toyne [747]. Trace [434]. Tracey [808]. Trade [824, 687, 162, 619, 1008]. Tradeoffs [1070]. Trading [442]. Traffic [815, 1088]. Training [391, 387]. Trains [1167]. Transformation [464]. Transition [937, 241, 1184]. Transitions [938].

Transport [1068, 1281, 643, 1281]. Transportation [143, 815]. Trappings [1047]. Travel [143]. Treasurer [1056, 1189, 1312]. Treatment [39, 1161, 534, 671, 877]. Trees [1067]. Trends [650, 158, 523, 479, 651]. Trials [462, 911]. Tribute [965, 635, 599, 695]. Trivedi [1040, 1074]. True [1236, 120]. Trunk [1019]. Truth [356]. Tsanov [777]. Tsokos [1059, 532]. Tuberculose [340]. Tue [780, 423]. Tukey [461, 1191, 1192]. Tummula [999]. Turnover [281, 1310]. Turoff [1007]. Twentieth [523]. Two [920, 1141, 834, 643, 331]. Two-Locus [1141]. Two-Person [331]. Two-Stage [834]. Type [1261]. Tzelekis [752].

U [1267, 1281, 405, 613]. U-Boat [613]. U.K [688, 1076, 56, 1040, 1235]. U.K. [279, 701]. U.N [158]. U.N./W.H.O [158]. U.S [1022, 1039, 1265]. Udny [245]. UK [730, 1275]. Ultimate [1245]. Unbalanced [1118, 1302]. Uncertainty [966, 721, 139, 824, 607]. Unconstrained [1262]. Underlying [451]. Understanding [601, 1005]. Unemployment [152]. Ungarischen [160]. Unified [50, 922]. Union [430]. Unit [895]. United [821, 928, 803, 196, 1110, 406, 152, 493, 122]. Units [907]. Unity [238]. Univariate [1226, 551, 321]. Universities [665, 468, 162, 542, 77, 1153]. University [969, 320, 709, 8, 970, 522, 816, 521, 708, 1170, 969, 873]. Unknown [634]. untitled [804, 805]. Upper [38, 800]. Uppuluri [1261]. Upton [1326]. Urban [317, 481, 512, 436, 410, 788, 1068]. Urbanization [651]. Urmitzer [211]. Urn [1134]. Usage [279]. Use [436, 176, 816, 419, 580, 157, 239, 204, 836]. Users [1034]. Uses [67, 464, 440]. Using [453, 279, 243, 1296]. Utility [420]. Utting [816].

V [915, 320, 269, 927, 1164, 885, 1125, 52, 1218, 717, 1254, 923, 220, 1058, 196, 28, 999, 1261, 213, 778, 772, 58, 974, 1129, 535, 769, 1249, 375, 11, 428, 853, 211]. Vacancies [110]. Vaccination [454]. Vajda [1283, 132, 848]. Valerie [376]. Validation [1216]. Valleron [1294]. Valuation [1090]. Value [800, 143, 1070]. Values [883]. Variables [903, 904, 973, 869, 178, 1185, 187, 480, 1107, 915, 910, 992]. Variance [1034, 920, 812, 704]. Variation [1018, 123]. Variations [835, 177, 262, 410, 376]. Varieties [388]. Vehicle [1088]. Vekerdi [525]. Ventilatory [729]. Ventures [612]. Verarbeitung [539]. Verification [1234]. Vernon [162]. Vervaat [462]. VI [478]. Vic [590]. Victor [482, 541]. Victoria [858]. Victorian [1002]. View [665, 12]. Views [377]. VII [1281, 479]. Vincze [775]. Virginia [704]. Vital [60, 875, 710, 860, 957]. Vocational [244, 99]. Vol [977, 1111, 1109, 1319, 820, 53, 526]. Vols [533, 775]. Volume [3, 96, 117, 201, 236, 348, 417, 381, 490, 516, 626, 663, 759, 794, 901, 933, 1045, 1180, 1306, 527, 1203, 399, 928, 22, 134, 456, 803, 250, 455, 392, 424, 1281, 1062, 1261, 957, 498, 480, 651, 1200]. Volumes [605, 402, 1227, 372]. Voluntary [803]. Vorträge [160]. Voter [937, 1184]. Voting [588, 938, 85].

W [1207, 1036, 1175, 784, 601, 989, 1264, 646, 917, 50, 822, 997, 544, 373, 948, 616, 678, 1025, 621, 928, 55, 72, 219, 225, 229, 300, 393, 14, 461, 462, 302, 1098, 740, 226, 803, 112, 741, 318, 779, 59, 1281, 1133, 1159, 1162, 1287, 953, 367, 612, 434, 1191, 1192, 411, 773, 265, 876, 982, 357, 706, 223, 500, 1136, 256, 1259, 267, 507, 470, 679, 1235, 74, 890]. W.H.O [158]. Waddington [613]. Wadsworth [688]. Waelbroeck [1075]. Wage [87]. Wages [574, 1040]. Wagner [105]. Wales [1232, 892, 629, 1243, 523, 309, 1093, 475]. Wall [229, 1077]. Walpole [1124, 980, 807]. Walsh [958]. Walter [300, 534, 604, 477]. Walters [949]. Waltman [781]. War [613, 574, 823]. Ward [337]. Waring [666, 698, 731]. Warren [1026]. Wasserman [703]. Wassily [1174]. Water [649]. Waterloo [320]. Watson [1281]. Wayne [740, 689]. Wealth [891, 1209]. Weber [608, 271]. Webster [1004]. Wedderburn [789]. Wednesday [589, 726]. Weights [700, 10, 353, 697]. Weil [690]. Weille [30]. Wein [28]. Weinberg [914]. Weintraub [259]. Weiss [1065]. Welch [244]. Welham [56]. Well [1266]. Well-Being [1266]. Wen [1064]. Wen-Hsiung [1064]. Werner [940]. Wesolowsky [1034]. Wessels [1222]. West [745, 872, 353]. Western [141]. Wetherill [920, 920, 49]. Wheat [179, 725]. Wheelwright [1101, 1226]. which [244]. Whilhelm [407]. White [751, 1288, 719]. Whittle [288]. WHO [301, 817, 67, 874]. Whose [765]. Wickelgren [689]. Wickwire [535, 840]. Wijngaarden [80]. Wild [441]. Wilfred [147]. Wilks [427]. Willemsen [601]. William [703, 782, 1100, 598, 323, 393, 742, 438, 290, 789, 103, 474, 957, 1279, 257]. Williams [978, 291, 229, 1319, 1230]. Williamson [376]. Willis [707]. Wilson [916]. Winder [434]. Winifred [464, 637, 983]. Winkler [457]. Winning [1105]. Winter [439, 823]. Within [1136]. Wolfe [562, 1262]. Women [523, 1093]. Wonnacott [1220, 1220]. Wood [998, 1225]. Woodroofe [842]. Woodward [196]. Work [1047, 523, 60, 1, 1144, 1217]. Workbook [684]. Worked [326]. Working [838, 1020, 1293]. Workloads [553]. World [650, 158, 301, 613, 25, 53, 950, 53, 651]. Worswick [440]. Wright [1170]. Wrighton [674]. Writings [957]. Written [520].

Yard [797]. Yates [192]. Year [177].

Year-to-Year [177]. Years [1047, 655, 892].

Yeats [751]. Yehuda [434]. Yodzis [1325]. Youden [773]. Young [333]. Ysidro [1147]. Yu [778, 211]. Yule [245]. Yvonne [876].

Z [269, 1133]. Zacks [283]. Zangwill [50]. Zaranowitz [541]. Zauberman [857]. Zeigler [1063]. Zellner [564, 845]. Zero [1142]. Ziegler [948]. Zuijlen [1163]. zur [947]. Zvi [403].

References

Herzberg:1969:RWD

[1] Agnes M. Herzberg and D. R. Cox. Recent work on the design of experiments: A bibliography and a review. Journal of the Royal Statistical Society. Series A (General), 132(1):29–67, ???? 1969. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343755. See correction

[1144].

Coen:1969:LRE

[2] P. J. Coen, E. D. Gomme, and M. G. Kendall. Lagged relationships in economic forecasting. Journal of the Royal Statistical Society. Series A (General), 132(2):133–163, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/

stable/2343782. See comments [153].

Anonymous:1970:VIa

[3] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 133(1):??, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343806.

Anonymous:1970:FMa

[4] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 133(1):??, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343807.

Cox:1970:SPM

[5] D. R. Cox and D. V. Hinkley. Some properties of multiserver queues with appointments. Journal of the Royal Statistical Society. Series A (General), 133(1):1–13, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343808.

Emshoff:1970:AMC

[6] James R. Emshoff and Alan Mercer. Aggregate models of consumer purchases. Journal of the Royal Statistical Society. Series A (General), 133(1):14– 32, ???? 1970. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343809.

Blackwell:1970:SP

[7] David Blackwell. On stationary policies. Journal of the Royal Statistical Society. Series A (General), 133(1):33– 37, ???? 1970. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343810.

Freeman:1970:MSS

[8] P. R. Freeman. A multivariate study of students’ performance in university examinations. Journal of the Royal Statistical Society. Series A (General), 133 (1):38–55, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343811.

Hey:1970:SNL

[9] G. B. Hey. Statistics in non-life insurance. Journal of the Royal Statistical Society. Series A (General), 133(1):56– 85, ???? 1970. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343812.

Kpedekpo:1970:HWC

[10] G. M. K. Kpedekpo. Heights and weights of children in Ghana. Journal of the Royal Statistical Society. Series A (General), 133(1):86–93, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343813.

Stuart:1970:BRB

[11] Alan Stuart. Book review: Elements of Nonparametric Statistics, by G. E. Noether; A Nonparametric Introduction to Statistics, by C. H. Kraft; Constance van Elden; Distribution-free Statistical Tests, by J. V. Bradley. Journal of the Royal Statistical Society. Series A (General), 133(1):94–97, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343814.

Wishart:1970:BRBa

[12] D. M. G. Wishart. Book review: Probability and Stochastic Processes: With a View Toward Applications, by Leo Breiman. Journal of the Royal Statistical Society. Series A (General), 133 (1):97–98, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343815.

Kingman:1970:BRB

[13] J. F. C. Kingman. Book review: Introduction to Mathematical Probability Theory, by Martin Eisen. Journal of the Royal Statistical Society. Series A (General), 133(1):98, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343816.

Herzberg:1970:BRB

[14] Agnes M. Herzberg. Book review: Introduction to Linear Models and the Design and Analysis of Experiments, by W. Mendenhall. Journal of the Royal Statistical Society. Series A (General), 133(1):98–99, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343817.

Murphy:1970:BRB

[15] B. P. Murphy. Book review: Solving Problems in Statistics, by J. Aitchison. Journal of the Royal Statistical Society. Series A (General), 133(1):99, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343818.

Thomas:1970:BRBa

[16] C. J. Thomas. Book review: Basic Statistics, by J. Snowdon. Journal of the Royal Statistical Society. Series A (General), 133(1):99–100, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343819.

Leser:1970:BRB

[17] C. E. V. Leser. Book review: Statistische Auswertungsmethoden, by Lothar Sachs. Journal of the Royal Statistical Society. Series A (General), 133(1): 100–101, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343820.

Smith:1970:BRB

[18] C. A. B. Smith. Book review: Biomathematics, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 133(1):101–102, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343821.

Campbell:1970:BRBa

[19] R. C. Campbell. Book review: Biometry: The Principles and Practice of Statistics in Biological Research, by R. R. Sokal; F. J. Rohlf. Journal of the Royal Statistical Society. Series A (General), 133(1):102, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343822.

Healy:1970:BRB

[20] M. J. R. Healy. Book review: Statistical Tables, by F. J. Rohlf; R. R. Sokal. Journal of the Royal Statistical Society. Series A (General), 133(1):103, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343823.

Campbell:1970:BRBb

[21] R. C. Campbell. Book review: A First Course in Biometry for Agriculture Students, by A. A. Rayner. Journal of the Royal Statistical Society. Series A (General), 133(1):103–104, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343824.

Finney:1970:BRB

[22] D. J. Finney. Book review: Theorie et Méthodes Statistiques, Applications Agronomiques. Volume 1, by P. Dagnelie. Journal of the Royal Statistical Society. Series A (General), 133(1): 104, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343825.

Lutz:1970:BRB

[23] W. Lutz. Book review: Statistische Methoden fur Erdwissenschaftler, by D. Marsal. Journal of the Royal Statistical Society. Series A (General), 133 (1):104–105, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343826.

Kendall:1970:BRB

[24] M. G. Kendall. Book review: Bibliography of Statistical Bibliographies, by H. O. Lancaster. Journal of the Royal Statistical Society. Series A (General), 133(1):105, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343827.

Palmer:1970:BRBa

[25] Cynthia Palmer. Book review: Statistics of Health Services and their Activities, by World Health Organization. Journal of the Royal Statistical Society. Series A (General), 133(1):105– 106, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343828.

Wallis:1970:BRB

[26] K. F. Wallis. Book review: An Econometric Model of the French Economy, by Michael K. Evans. Journal of the Royal Statistical Society. Series A (General), 133(1):106–107, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343829.

Williamson:1970:BRB

[27] J. H. Williamson. Book review: Production Functions, Foreign Investment and Growth: A Study Based on the Argentine Manufacturing Sector 1946–1961, by J. M. Katz. Journal of the Royal Statistical Society. Series A (General), 133(1):107, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343830.

Mills:1970:BRB

[28] G. Mills. Book review: The Optimal Staging and Phasing of Multi-Product Capacity, by Harold H. Wein; V. P. Sreedharan. Journal of the Royal Statistical Society. Series A (General), 133 (1):108, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343831.

Wilkinson:1970:BRBa

[29] R. K. Wilkinson. Book review: Incomes in Postwar Europe: A Study of Policies, Growth and Distribution, by Secretariat of the Economic Commission for Europe. Journal of the Royal Statistical Society. Series A (General), 133(1):109, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343832.

Gwilliam:1970:BRB

[30] K. M. Gwilliam. Book review: Reappraisal of a Road Project in Iran, by H. G. van der Tak; J. de Weille. Journal of the Royal Statistical Society. Series A (General), 133(1):109–110, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343833.

Anonymous:1970:CNa

[31] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 133(1):111–127, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343834.

Benjamin:1970:RFG

[32] Bernard Benjamin and Iris Douglas. R. F. George. Journal of the Royal Statistical Society. Series A (General), 133 (1):128–129, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343835.

Anonymous:1970:RPa

[33] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 133(1):130–136, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343836.

Anonymous:1970:LAa

[34] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 133(1):137–138, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343837.

Anonymous:1970:BMa

[35] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 133(1):i–vi, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343838.

Anonymous:1970:FMb

[36] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 133(2):??, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343687.

Sprent:1970:SPS

[37] P. Sprent. Some problems of statistical consultancy. Journal of the Royal Statistical Society. Series A (General), 133(2):139–165, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343688.

Armitage:1970:TMU

[38] Peter H. Armitage, Celia M. Phillips, and Judith Davies. Towards a model of the upper secondary school system. Journal of the Royal Statistical Society. Series A (General), 133(2):166– 205, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343689.

Aitchison:1970:SPT

[39] J. Aitchison. Statistical problems of treatment allocation. Journal of the Royal Statistical Society. Series A (General), 133(2):206–239, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343690.

Benjamin:1970:PCA

[40] Bernard Benjamin. The 1971 population census and after. Journal of the Royal Statistical Society. Series A (General), 133(2):240–256, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343691.

Wishart:1970:BRBb

[41] D. M. G. Wishart. Book review: Stochastic Convergence, by Eugene Lukacs. Journal of the Royal Statistical Society. Series A (General), 133 (2):257, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343692.

Whittle:1970:BRB

[42] P. Whittle. Book review: Advanced Seminar on Spectral Analysis of Time Series, by Bernard Harris. Journal of the Royal Statistical Society. Series A (General), 133(2):257–258, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343693.

Beard:1970:BRB

[43] R. E. Beard. Book review: Stochastic Theory of a Risk Business, by H. L. Seal. Journal of the Royal Statistical Society. Series A (General), 133 (2):259–260, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343694.

Conway:1970:BRB

[44] Freda Conway. Book review: Introduction to Statistics: Selected Procedures for the Behavioural Sciences, by Robert Fried. Journal of the Royal Statistical Society. Series A (General), 133 (2):260–261, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343695.

Buckland:1970:BRB

[45] W. R. Buckland. Book review: Precision Measurement and Calibration: Statistical Concepts and Procedures, by H. H. Ku. Journal of the Royal Statistical Society. Series A (General), 133 (2):261, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343696.

Hamilton:1970:BRB

[46] M. Hamilton. Book review: Applications of Mental Health Statistics, by Morton Kramer. Journal of the Royal Statistical Society. Series A (General), 133(2):261–262, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343697.

Clementson:1970:BRB

[47] A. T. Clementson. Book review: Computer Simulation Models, by J. Smith. Journal of the Royal Statistical Society. Series A (General), 133(2):262, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343698.

Wilkinson:1970:BRBb

[48] R. K. Wilkinson. Book review: New Developments in Survey Sampling, by N. L. Johnson; H. Smith. Journal of the Royal Statistical Society. Series A (General), 133(2):262–263, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343699.

Ward:1970:BRB

[49] Denis H. Ward. Book review: Sampling Inspection and Quality Control, by G. Barrie Wetherill. Journal of the Royal Statistical Society. Series A (General), 133(2):263–264, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343700.

Beale:1970:BRB

[50] E. M. L. Beale. Book review: Nonlinear Programming: A Unified Approach, by W. Zangwill. Journal of the Royal Statistical Society. Series A (General), 133(2):264–265, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343701.

Holmes:1970:BRB

[51] R. A. Holmes. Book review: Introduction to Linear Algebra for Social Scientists, by Gordon Mills. Journal of the Royal Statistical Society. Series A (General), 133(2):265–266, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343702.

Haley:1970:BRB

[52] K. B. Haley. Book review: Introduction to Mathematical Programming for Accountants, by B. V. Carsberg. Journal of the Royal Statistical Society. Series A (General), 133(2):266, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343703.

Palmer:1970:BRBb

[53] Cynthia Palmer. Book review: World Health Statistics Annual 1963: Vol. 3. Health Personnel and Hospital Establishments, by World Health Organization. Journal of the Royal Statistical Society. Series A (General), 133(2): 266–267, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343704.

Cooper:1970:BRB

[54] R. Cooper. Book review: National Accounts Statistics: Sources and Methods, by Rita Maurice. Journal of the Royal Statistical Society. Series A (General), 133(2):267, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343705.

Deane:1970:BRB

[55] Marjorie Deane. Book review: The British Economy 1870–1939, by Derek H. Aldcroft; Harry W. Richardson. Journal of the Royal Statistical Society. Series A (General), 133(2):268, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343706.

Burman:1970:BRB

[56] J. P. Burman. Book review: Monetary Circulation in the U.K, by P. J. Welham. Journal of the Royal Statistical Society. Series A (General), 133 (2):268–269, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343707.

Cunningham:1970:BRB

[57] N. J. Cunningham. Book review: On the Nature of Business Success, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 133 (2):269–270, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343708.

Pearson:1970:BRB

[58] E. S. Pearson. Book review: Tables of Bivariate Random Normal Deviates, by P. V. Krishna Iyer; P. S. Simha. Journal of the Royal Statistical Society. Series A (General), 133(2):270, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343709.

Lancaster:1970:BRB

[59] H. O. Lancaster. Book review: An Annotated Bibliography of Statistical Stylistics, by Richard W. Bailey; Lubomir Dolezel. Journal of the Royal Statistical Society. Series A (General), 133(2):271, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343710.

George:1970:BRB

[60] C. Oswald George. Book review: My Statistics are Vital: Life and Work as a Statistician, by Madge Dugdale. Journal of the Royal Statistical Society. Series A (General), 133(2):271, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343711.

Anonymous:1970:RPb

[61] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 133(2):272–283, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343712.

Anonymous:1970:CNb

[62] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 133(2):284–286, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343713.

Pearson:1970:DH

[63] E. S. Pearson. David Heron, 1881– 1969. Journal of the Royal Statistical Society. Series A (General), 133(2): 287–291, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343714.

Anonymous:1970:LAb

[64] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 133(2):292–294, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343715.

Anonymous:1970:BMb

[65] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 133(2):i–vi, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343716.

Anonymous:1970:FMc

[66] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 133(3):??, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343558.

Ashford:1970:WUH

[67] J. R. Ashford and N. G. Pearson. Who uses the health services and why? Journal of the Royal Statistical Society. Series A (General), 133(3):295– 357, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343559.

Barnard:1970:SCD

[68] G. A. Barnard and M. D. McCreath. Subject commitments and the demand for higher education. Journal of the Royal Statistical Society. Series A (General), 133(3):358–408, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343560.

Lancaster:1970:PBS

[69] H. O. Lancaster. Problems in the bibliography of statistics. Journal of the Royal Statistical Society. Series A (General), 133(3):409–441, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343561.

Gani:1970:CNI

[70] J. Gani. On coping with new information in probability and statistics. Journal of the Royal Statistical Society. Series A (General), 133(3):442–462, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343562.

Horowitz:1970:ECC

[71] Ira Horowitz. Employment concentration in the common market: An entropy approach. Journal of the Royal Statistical Society. Series A (General), 133(3):463–479, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343563.

Evans:1970:BRB

[72] I. G. Evans. Book review: Applied Probability, by W. A. Thompson. Journal of the Royal Statistical Society. Series A (General), 133(3):480–481, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343564.

Nixon:1970:BRB

[73] Carol M. Nixon. Book review: Some Aspects of Queueing and Storage Systems, by A. Ghosal. Journal of the Royal Statistical Society. Series A (General), 133(3):481, ????

1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343565.

Wishart:1970:BRBc

[74] D. M. G. Wishart. Book review: The Single Server Queue, by J. W. Cohen. Journal of the Royal Statistical Society. Series A (General), 133(3):482, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343566.

Everitt:1970:BRB

[75] B. S. Everitt. Book review: Distributions in Statistics: Discrete Distributions, by N. L. Johnson; S. Kotz. Journal of the Royal Statistical Society. Series A (General), 133(3):482–483, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343567.

Bissell:1970:BRB

[76] A. F. Bissell. Book review: Statistics for Technology, by C. Chatfield. Journal of the Royal Statistical Society. Series A (General), 133(3):483–484, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343568.

Thomas:1970:BRBb

[77] C. J. Thomas. Book review: Principles of Statistical Techniques: A First Course, from the Beginnings, for Schools and Universities, with many Examples and Solutions, by P. G. Moore. Journal of the Royal Statistical Society. Series A (General), 133 (3):484–485, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343569.

Coyle:1970:BRB

[78] R. G. Coyle. Book review: Applied Statistics for Management Studies, by D. Croft. Journal of the Royal Statistical Society. Series A (General), 133 (3):485, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343570.

Daisley:1970:BRB

[79] P. A. Daisley. Book review: Quality and Reliability: An Integrated Approach, by C. S. Smith. Journal of the Royal Statistical Society. Series A (General), 133(3):485–486, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343571.

Douglas:1970:BRB

[80] A. S. Douglas. Book review: Report on the Algorithmic Language, Algol 68, by A. van Wijngaarden. Journal of the Royal Statistical Society. Series A (General), 133(3):486–487, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343572.

Fisher:1970:BRBa

[81] G. R. Fisher. Book review: Model Building in the Social Sciences, by Roger Peltier. Journal of the Royal Statistical Society. Series A (General), 133(3):488, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343573.

Fisher:1970:BRBb

[82] G. R. Fisher. Book review: Economics and Information Theory, by H. Theil. Journal of the Royal Statistical Society.

Series A (General), 133(3):489, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343574.

Thomas:1970:BRBc

[83] C. J. Thomas. Book review: The Population Census, by B. Benjamin. Journal of the Royal Statistical Society. Series A (General), 133(3):490, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343575.

Brass:1970:BRB

[84] W. Brass. Book review: Population Growth and the Brain Drain, by F. Bechhofer. Journal of the Royal Statistical Society. Series A (General), 133 (3):490–491, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343576.

Williams:1970:BRB

[85] G. L. Williams. Book review: Theory of Voting, by R. Farquharson. Journal of the Royal Statistical Society. Series A (General), 133(3):491–492, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343577.

Nicholson:1970:BRBa

[86] R. J. Nicholson. Book review: Household Food Consumption and Expenditure: 1967. Journal of the Royal Statistical Society. Series A (General), 133 (3):492, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343578.

Nicholson:1970:BRBb

[87] R. J. Nicholson. Book review: Income and Wage Distributions, by K. Bjerke.

Journal of the Royal Statistical Society. Series A (General), 133(3):493, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343579.

Nevitt:1970:BRB

[88] Della A. Nevitt. Book review: Aycliffe to Cumbernauld; A Study of Seven New Towns in Their Regions, by R. Thomas. Journal of the Royal Statistical Society. Series A (General), 133 (3):494, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343580.

Anonymous:1970:CNc

[89] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 133(3):495–497, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343581.

Heien:1970:AIP

[90] Dale M. Heien. Addendum: Income and price lags in consumer demand analysis. Journal of the Royal Statistical Society. Series A (General), 133 (3):497, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343582.

Campion:1970:HEB

[91] Harry Campion. Harold Ernest Bishop. Journal of the Royal Statistical Society. Series A (General), 133(3):498, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343583.

Kendall:1970:AR

[92] D. G. Kendall. Alfréd Rényi. Journal of the Royal Statistical Society.

Series A (General), 133(3):499, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343584.

Anonymous:1970:RPc

[93] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 133(3):500–505, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343585.

Anonymous:1970:LAc

[94] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 133(3):506–508, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343586.

Anonymous:1970:BMc

[95] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 133(3):i–iv, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343587.

Anonymous:1970:VIb

[96] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 133(4):605– 607, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344295.

Anonymous:1970:FMd

[97] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 133(4):??, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344296.

Allen:1970:OSO

[98] Roy Allen. On official statistics and official statisticians. Journal of the Royal Statistical Society. Series A (General), 133(4):509–530, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344297.

Welch:1970:SVC

[99] B. L. Welch. Statistics — a vocational or a cultural study? Journal of the Royal Statistical Society. Series A (General), 133(4):531–555, ???? 1970. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2344298. See comments [244].

Anonymous:1970:PSG

[100] Anonymous. Proceedings of a special general meeting. Journal of the Royal Statistical Society. Series A (General), 133(4):556, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344299.

Anonymous:1970:POH

[101] Anonymous. Proceedings of the one hundred and thirty-sixth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 133 (4):557, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344300.

Anonymous:1970:ARC

[102] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 133 (4):558–581, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344301.

Nicholson:1970:BRBc

[103] R. J. Nicholson. Book review: Introduction to Economic Statistics, by William C. Merrill; Karl A. Fox. Journal of the Royal Statistical Society. Series A (General), 133(4):582–583, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344302.

Thomas:1970:BRBd

[104] J. J. Thomas. Book review: Empirical Econometrics, by J. S. Cramer. Journal of the Royal Statistical Society. Series A (General), 133(4):583–584, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344303.

Vajda:1970:BRB

[105] S. Vajda. Book review: Principles of Operations Research, with Applications to Managerial Decisions, by Harvey M. Wagner. Journal of the Royal Statistical Society. Series A (General), 133 (4):584–585, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344304.

Lee:1970:BRBa

[106] P. M. Lee. Book review: Information Theory and Reliable Communication, by Robert T. Gallager. Journal of the Royal Statistical Society. Series A (General), 133(4):585, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344305.

Lee:1970:BRBb

[107] P. M. Lee. Book review: Einführung in die Mathematischen Grundlagen der Informationstheorie, by Ernst Schultze.

Journal of the Royal Statistical Society. Series A (General), 133(4):586, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344306.

Lee:1970:BRBc

[108] P. M. Lee. Book review: Aplicaţii ale Teoriei Informaţiei: Sisteme Dinamice, Sisteme Cibernetice, by Silviu Guiaşu. Journal of the Royal Statistical Society. Series A (General), 133 (4):586, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344307.

Wilkinson:1970:BRBc

[109] R. Wilkinson. Book review: Handbook of Selected Sample Surveys in the Federal Government with Annotated Bibliography, 1960–1968, by Jean Namias. Journal of the Royal Statistical Society. Series A (General), 133(4):587, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344308.

Routh:1970:BRB

[110] Guy Routh. Book review: Measuring Job Vacancies, by John G. Myers; Daniel Creamer. Journal of the Royal Statistical Society. Series A (General), 133(4):587–588, ???? 1970. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344309.

Zauberman:1970:BRB

[111] Alfred Zauberman. Book review: Theory of Economic Growth, by M. Morishima. Journal of the Royal Statistical Society. Series A (General), 133 (4):588–590, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344310.

Kelsall:1970:BRB

[112] R. K. Kelsall. Book review: Students in Conflict: L.S.E. in 1967, by Tessa Blackstone; Kathleen Gales; R. Hadley; W. Lewis. Journal of the Royal Statistical Society. Series A (General), 133 (4):590–591, ???? 1970. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344311.

Anonymous:1970:CNd

[113] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 133(4):591, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344312.

Anonymous:1970:RPd

[114] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 133(4):592–601, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344313.

Anonymous:1970:LAd

[115] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 133(4):602–604, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344314.

Anonymous:1970:BMd

[116] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 133(4):i–iv, ???? 1970. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344315.

Anonymous:1971:VIa

[117] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 134(1):??, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343968.

Anonymous:1971:FMa

[118] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 134(1):??, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343969.

Benjamin:1971:SM

[119] B. Benjamin. The statistician and the manager. Journal of the Royal Statistical Society. Series A (General), 134(1): 1–14, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343970.

McPherson:1971:RST

[120] C. K. McPherson and P. Armitage. Repeated significance tests on accumulating data when the null hypothesis is not true. Journal of the Royal Statistical Society. Series A (General), 134(1):15– 25, ???? 1971. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343971.

Gani:1971:MAP

[121] J. Gani. Models in applied probability. Journal of the Royal Statistical Society. Series A (General), 134(1):26–38, ???? 1971. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343972.

Primorac:1971:CPU

[122] Emile Primorac. Censuses of production in the United Kingdom 1924–58, an attempt at reconciliation. Journal of the Royal Statistical Society. Series A (General), 134(1):39–56, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343973.

Thomas:1971:SVR

[123] J. J. Thomas and Kenneth F. Wallis. Seasonal variation in regression analysis. Journal of the Royal Statistical Society. Series A (General), 134(1):57– 72, ???? 1971. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343974.

Hart:1971:EOM

[124] P. E. Hart. Entropy and other measures of concentration. Journal of the Royal Statistical Society. Series A (General), 134(1):73–85, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343975.

Stuart:1971:BRB

[125] Alan Stuart. Book review: The Chi-Squared Distribution, by H. O. Lancaster. Journal of the Royal Statistical Society. Series A (General), 134(1):86– 87, ???? 1971. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343976.

Campbell:1971:BRB

[126] R. C. Campbell. Book review: Statistical Methods, by H. Lucas. Journal of the Royal Statistical Society. Series A (General), 134(1):87–88, ????

1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343977.

Corlett:1971:BRB

[127] W. J. Corlett. Book review: Basic Concepts of Probability and Statistics, by J. L. Hodges,; E. L. Lehmann. Journal of the Royal Statistical Society. Series A (General), 134(1):88, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343978.

Bailey:1971:BRB

[128] B. J. R. Bailey. Book review: Statistics: The Essentials for Research, by Henry E. Klugh. Journal of the Royal Statistical Society. Series A (General), 134(1):89–90, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343979.

Farlie:1971:BRBa

[129] D. J. G. Farlie. Book review: A Course in Nonparametric Statistics, by Jaroslav Hajek. Journal of the Royal Statistical Society. Series A (General), 134(1):90, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343980.

Kingman:1971:BRB

[130] J. F. C. Kingman. Book review: Probability on Discrete Sample Spaces with Applications, by Anne E. Scheerer. Journal of the Royal Statistical Society. Series A (General), 134(1):91, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343981.

Moore:1971:BRB

[131] P. G. Moore. Book review: Actuaries and Financial Planning, by Committee of the Institute of Actuaries Students’ Society. Journal of the Royal Statistical Society. Series A (General), 134(1):91– 92, ???? 1971. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343982.

Morton:1971:BRB

[132] G. Morton. Book review: Planning by Mathematics, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 134(1):92, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343983.

Mills:1971:BRB

[133] G. Mills. Book review: Applied Linear Programming, by N. J. Driebeek. Journal of the Royal Statistical Society. Series A (General), 134(1):92–93, ???? 1971. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343984.

Finney:1971:BRBa

[134] D. J. Finney. Book review: Theorie et Méthodes Statistiques, Applications Agronomiques. Volume 2, by P. Dagnelie. Journal of the Royal Statistical Society. Series A (General), 134(1):93– 94, ???? 1971. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343985.

Leser:1971:BRBa

[135] C. E. V. Leser. Book review: Eléments d’Econometrie, by René Roy. Journal of the Royal Statistical Society.

Series A (General), 134(1):94, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343986.

Ord:1971:BRBa

[136] Keith Ord. Book review: Introduction to Sets and Mappings in Modern Economics, by H. Nikaido; K. Sato. Journal of the Royal Statistical Society. Series A (General), 134(1):94–95, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343987.

Crossley:1971:BRB

[137] J. R. Crossley. Book review: Price Expectations and the Behaviour of the Price Level, by R. M. Solow. Journal of the Royal Statistical Society. Series A (General), 134(1):95–97, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343988.

Shaw:1971:BRBa

[138] R. W. Shaw. Book review: The Behavior of Industrial Prices, by G. J. Stigler; J. K. Kindahl. Journal of the Royal Statistical Society. Series A (General), 134(1):97–98, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343989.

Ball:1971:BRB

[139] R. J. Ball. Book review: Techniques for Project Appraisal Under Uncertainty, by Shlomo Reutlinger. Journal of the Royal Statistical Society. Series A (General), 134(1):98–99, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343990.

Boyle:1971:BRB

[140] Lawrence Boyle. Book review: Federal Revenue Sharing with the States: Problems and Promises, by Michael E. Levy; Juan de Torres. Journal of the Royal Statistical Society. Series A (General), 134(1):99–100, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343991.

Hillmann:1971:BRB

[141] H. C. Hillmann. Book review: Dating Postwar Business Cycles: Methods and their Application to Western Germany, 1950–67, by Ilse Mintz. Journal of the Royal Statistical Society. Series A (General), 134(1):100, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343992.

Lewis:1971:BRB

[142] J. Parry Lewis. Book review: Economics of the Construction Industry, by Peter J. Cassimatis. Journal of the Royal Statistical Society. Series A (General), 134(1):101, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343993.

Gwilliam:1971:BRB

[143] K. M. Gwilliam. Book review: The Value of Time in Passenger Transportation: The Demand for Air Travel, by R. Gronau. Journal of the Royal Statistical Society. Series A (General), 134(1):101–102, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343994.

Welham:1971:BRB

[144] P. J. Welham. Book review: Home Mortgage Delinquency and Foreclosure, by John P. Herzog; James S. Earley. Journal of the Royal Statistical Society. Series A (General), 134(1):102– 103, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343995.

Anonymous:1971:CNa

[145] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 134(1):104–108, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343996.

Anonymous:1971:RPa

[146] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 134(1):109–123, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343997.

Good:1971:WP

[147] I. J. Good. Wilfred perks. Journal of the Royal Statistical Society. Series A (General), 134(1):124, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343998.

Anonymous:1971:LAa

[148] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 134(1):125–129, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343999.

Anonymous:1971:BMa

[149] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 134(1):130–iv, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344000.

Anonymous:1971:FMb

[150] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 134(2):??, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343870.

Doll:1971:ADC

[151] Richard Doll. The age distribution of cancer: Implications for models of carcinogenesis. Journal of the Royal Statistical Society. Series A (General), 134 (2):133–166, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343871. See

corrigendum [233].

Bray:1971:DEE

[152] Jeremy Bray. Dynamic equations for economic forecasting with the G.D.P. — unemployment relation and the growth of G.D.P. in the United Kingdom as an example. Journal of the Royal Statistical Society. Series A (General), 134(2):167–228, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343872.

Box:1971:SCP

[153] George E. P. Box and Paul Newbold. Some comments on a paper of Coen, Gomme and Kendall. Journal of the Royal Statistical Society. Series A (General), 134(2):229–240, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343873. See [2].

Nicholson:1971:BRB

[154] R. J. Nicholson. Book review: Introductory Mathematical Statistics, by E. Kreyszig. Journal of the Royal Statistical Society. Series A (General), 134 (2):241–242, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343874.

Barnett:1971:BRBa

[155] V. D. Barnett. Book review: Choice Against Chance: An Introduction to Statistical Decision Theory, by John Aitchison. Journal of the Royal Statistical Society. Series A (General), 134 (2):242–243, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343875.

Kemp:1971:BRB

[156] C. D. Kemp. Book review: Random Counts in Models and Structures, by G. P. Patil. Journal of the Royal Statistical Society. Series A (General), 134 (2):243–244, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343876.

Oliver:1971:BRB

[157] F. R. Oliver. Book review: Abridged Tables for Use of Examination Candidates. Journal of the Royal Statistical Society. Series A (General), 134 (2):245, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343877.

Brass:1971:BRB

[158] W. Brass. Book review: Programme of Analysis of Mortality Trends and Levels: Report of a Joint U.N./W.H.O. Meeting, by World Health Organization. Journal of the Royal Statistical Society. Series A (General), 134(2): 245–246, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343878.

Nelder:1971:BRB

[159] J. A. Nelder. Book review: Statistics In Biology, by Chester I. Bliss. Journal of the Royal Statistical Society. Series A (General), 134(2):246–247, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343879.

Finney:1971:BRBb

[160] D. J. Finney. Book review: Vorträge der II. Ungarischen Biometrischen Konferenz, by I. Juvancz. Journal of the Royal Statistical Society. Series A (General), 134(2):247, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343880.

Bellamy:1971:BRB

[161] R. Bellamy. Book review: Soviet Economics, by Michael Kaser. Journal of the Royal Statistical Society. Series A (General), 134(2):247–248, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343881.

Jones:1971:BRB

[162] A. J. Jones. Book review: The Technology Factor in International Trade. A Conference of the Universities — National Bureau Committee for Economic Research, by R. Vernon. Journal of the Royal Statistical Society. Series A (General), 134(2):248–249, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343882.

Lawson:1971:BRB

[163] G. H. Lawson. Book review: Investment Analysis and Portfolio Management: Readings from British Publications, by Basil Taylor. Journal of the Royal Statistical Society. Series A (General), 134(2):249–250, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343883.

Shaw:1971:BRBb

[164] G. K. Shaw. Book review: Public Investment, The Rate of Return, and Optimal Fiscal Policy, by Kenneth J. Arrow; Mordecai Kurz. Journal of the Royal Statistical Society. Series A (General), 134(2):250–251, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343884.

Shaw:1971:BRBc

[165] R. W. Shaw. Book review: Atlas of Economic Structure and Policies, by P. Sargent Florence. Journal of the Royal Statistical Society. Series A (General), 134(2):251, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343885.

Leser:1971:BRBb

[166] G. E. V. Leser. Book review: The Irish Economy since 1922, by James Meenan. Journal of the Royal Statistical Society. Series A (General), 134 (2):252, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343886.

Anonymous:1971:CNb

[167] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 134(2):253–254, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343887.

Anonymous:1971:RPb

[168] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 134(2):255–268, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343888.

Aukrust:1971:GJ

[169] Odd Aukrust. Gunnar Jahn. Journal of the Royal Statistical Society. Series A (General), 134(2):269, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343889.

Anonymous:1971:LAb

[170] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 134(2):270–271, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343890.

Anonymous:1971:BMb

[171] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 134(2):i–vii, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343891.

Anonymous:1971:FMc

[172] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 134(3):??, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344235.

Nicholson:1971:DIH

[173] R. J. Nicholson and N. Topham. The determinants of investment in housing by local authorities: An econometric approach. Journal of the Royal Statistical Society. Series A (General), 134 (3):273–320, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344236. See

corrigendum [275].

Cormack:1971:RC

[174] R. M. Cormack. A review of classification. Journal of the Royal Statistical Society. Series A (General), 134 (3):321–367, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344237.

Fryer:1971:HMD

[175] J. G. Fryer. On the homogeneity of the marginal distributions of a multidimensional contingency table. Journal of the Royal Statistical Society. Series A (General), 134(3):368–371, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344238.

Kpedekpo:1971:PSC

[176] G. M. K. Kpedekpo. Pre-school children in Ghana — the use of prior information. Journal of the Royal Statistical Society. Series A (General), 134 (3):372–373, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344239.

Thatcher:1971:YYV

[177] A. R. Thatcher. Year-to-year variations in the earnings of individuals. Journal of the Royal Statistical Society. Series A (General), 134(3):374–382, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344240.

Gudmundsson:1971:TSA

[178] G. Gudmundsson. Time-series analysis of imports, exports and other economic variables. Journal of the Royal Statistical Society. Series A (General), 134 (3):383–412, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344241.

Granger:1971:NLS

[179] C. W. J. Granger and A. O. Hughes. A new look at some old data: The Beveridge wheat price series. Journal of the Royal Statistical Society. Series A (General), 134(3):413–428, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/

stable/2344242. See comments [725, 314].

Hoyle:1971:SDI

[180] M. H. Hoyle. Spoilt data — an introduction and bibliography. Journal of the Royal Statistical Society. Series A (General), 134(3):429–439, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344243.

Walters:1971:MCE

[181] A. A. Walters. A method for collecting economic statistics of shipping. Journal of the Royal Statistical Society. Series A (General), 134(3):440–448, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344244.

Kendall:1971:BRBa

[182] David Kendall. Book review: Foundations of Probability, by A. Rényi. Journal of the Royal Statistical Society. Series A (General), 134(3):449–450, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344245.

Kendall:1971:BRBb

[183] M. G. Kendall. Book review: Time Series Analysis, Forecasting and Control, by G. E. P. Box; G. M. Jenkins. Journal of the Royal Statistical Society. Series A (General), 134(3):450–453, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344246.

Plackett:1971:BRBa

[184] R. L. Plackett. Book review: The Analysis of Binary Data, by D. R. Cox. Journal of the Royal Statistical Society. Series A (General), 134(3):453, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344247.

Chatfield:1971:BRB

[185] C. Chatfield. Book review: Probability and Statistical Inference, by R. G. Krutchkoff. Journal of the Royal Statistical Society. Series A (General), 134 (3):453–454, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344248.

Rahman:1971:BRB

[186] N. A. Rahman. Book review: Basic Probability Theory, by Robert B. Ash. Journal of the Royal Statistical Society. Series A (General), 134(3):454– 455, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344249.

Kendall:1971:BRBc

[187] D. G. Kendall. Book review: Random Variables and Probability Distributions, by H. Cramér. Journal of the Royal Statistical Society. Series A (General), 134(3):455–456, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344250.

Healy:1971:BRB

[188] M. J. R. Healy. Book review: Statistical Models and their Experimental Application, by P. Ottestad. Journal of the Royal Statistical Society. Series A (General), 134(3):456, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344251.

Beard:1971:BRB

[189] R. E. Beard. Book review: Mathematical Methods in Risk Theory, by H. Buhlmann. Journal of the Royal Statistical Society. Series A (General), 134 (3):456–458, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344252.

Godwin:1971:BRB

[190] H. J. Godwin. Book review: Analytic Inequalities, by D. S. Mitrinović. Journal of the Royal Statistical Society. Series A (General), 134(3):458, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344253.

Plackett:1971:BRBb

[191] R. L. Plackett. Book review: Studies in the History of Statistics and Probability, by E. S. Pearson; M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 134(3):459, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344254.

Pearce:1971:BRB

[192] S. C. Pearce. Book review: Experimental Design: Selected Papers, by F. Yates. Journal of the Royal Statistical Society. Series A (General), 134(3): 459–460, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344255.

Knott:1971:BRBa

[193] M. Knott. Book review: Readings in Statistics for the Behavioral Sciences, by E. F. Heermann; L. A. Braskamp. Journal of the Royal Statistical Society. Series A (General), 134(3):460– 461, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344256.

Vajda:1971:BRB

[194] S. Vajda. Book review: Applications of Mathematical Programming Techniques, by E. M. L. Beale. Journal of the Royal Statistical Society.

Series A (General), 134(3):461, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344257.

Buckland:1971:BRB

[195] W. R. Buckland. Book review: Proceedings of the Fifth International Conference on Operational Research. OR 69, by J. Lawrence. Journal of the Royal Statistical Society. Series A (General), 134(3):462, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344258.

Lewes:1971:BRB

[196] F. M. M. Lewes. Book review: The Anatomy of Regional Activity Rates, by John Bower; Regional Social Accounts for the United Kingdom, by V. H. Woodward. Journal of the Royal Statistical Society. Series A (General), 134(3):462–463, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344259.

Anonymous:1971:CNc

[197] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 134(3):464–466, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344260.

Anonymous:1971:RPc

[198] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 134(3):467–479, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344261.

Anonymous:1971:LAc

[199] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 134(3):480–483, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344262.

Anonymous:1971:BMc

[200] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 134(3):i–vi, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344263.

Anonymous:1971:VIb

[201] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 134(4):719– 721, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343652.

Anonymous:1971:FMd

[202] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 134(4):??, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343653.

Luck:1971:MCI

[203] G. M. Luck, J. R. Morgan, P. J. Farmer, and J. Stringer. The management of capital investment. Journal of the Royal Statistical Society. Series A (General), 134(4):485–533, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343654.

Rosenbaum:1971:RUS

[204] S. Rosenbaum. A report on the use of statistics in social science research. Journal of the Royal Statistical Society. Series A (General), 134(4):534– 610, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343655.

Bee:1971:SMS

[205] Ronald E. Bee. Statistical methods in the study of the Masoretic text of the Old Testament. Journal of the Royal Statistical Society. Series A (General), 134(4):611–622, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343656.

Reep:1971:SCB

[206] C. Reep, R. Pollard, and B. Benjamin. Skill and chance in ball games. Journal of the Royal Statistical Society. Series A (General), 134(4):623–629, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343657.

Berman:1971:RIO

[207] Lawrence S. Berman. Recent improvements in official economic statistics. Journal of the Royal Statistical Society. Series A (General), 134(4):630– 651, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343658.

Anonymous:1971:POH

[208] Anonymous. Proceedings of the one hundred and thirty-seventh Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 134 (4):652, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343659.

Anonymous:1971:ARC

[209] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 134 (4):653–675, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343660.

Edwards:1971:BRB

[210] A. W. F. Edwards. Book review: Statistical Inference, by S. D. Silvey. Journal of the Royal Statistical Society. Series A (General), 134(4):676–677, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343661.

Wishart:1971:BRB

[211] D. M. G. Wishart. Book review: Probability Theory, by Yu. V. Prohorov; Yu. A. Rozanov; K. Krickeberg; H. Urmitzer. Journal of the Royal Statistical Society. Series A (General), 134 (4):678, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343662.

Bather:1971:BRB

[212] J. A. Bather. Book review: Markovian Decision Processes, by H. Mine; S. Osaki. Journal of the Royal Statistical Society. Series A (General), 134 (4):678–679, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343663.

Ord:1971:BRBb

[213] J. K. Ord. Book review: Families of Bivariate Distributions, by K. V. Mardia. Journal of the Royal Statistical Society. Series A (General), 134(4):679– 680, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343664.

Seber:1971:BRB

[214] G. A. F. Seber. Book review: The Co-Ordinate Free Approach to Gauss-Estimation, by H. Drygas. Journal of the Royal Statistical Society. Series A (General), 134(4):680–681, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343665.

Barnett:1971:BRBb

[215] V. D. Barnett. Book review: Nonparametric Techniques in Statistical Inference, by M. L. Puri. Journal of the Royal Statistical Society. Series A (General), 134(4):681–682, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343666.

Lindley:1971:BRB

[216] D. V. Lindley. Book review: Statistical Tolerance Regions: Classical and Bayesian, by Irwin Guttman. Journal of the Royal Statistical Society. Series A (General), 134(4):682, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343667.

Farlie:1971:BRBb

[217] D. J. G. Farlie. Book review: Rank Correlation Methods, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 134 (4):682, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343668.

Knott:1971:BRBb

[218] Martin Knott. Book review: Statistical Functions: A Source of Practical Derivations Based on Elementary Mathematics, by Buddy L. Myers; Norbert L. Enrick. Journal of the Royal Statistical Society. Series A (General), 134(4):683, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343669.

Everitt:1971:BRB

[219] B. S. Everitt. Book review: Statistical Reasoning in Psychology and Education, by Edward W. Minium. Journal of the Royal Statistical Society. Series A (General), 134(4):683–684, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343670.

Irving:1971:BRB

[220] Doreen Irving. Book review: Elements of Medical Statistics, by J. V. Smart. Journal of the Royal Statistical Society. Series A (General), 134(4):684, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343671.

Peare:1971:BRB

[221] C. W. Derek Peare. Book review: Mathematics and Statistics for Economists, by Gerhard Tintner; Charles B. Millham. Journal of the Royal Statistical Society. Series A (General), 134(4):684–685, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343672.

Prais:1971:BRB

[222] Z. Prais. Book review: Basic Statistics for Business and Economics: A Computer-Oriented Text, by R. L. Burford. Journal of the Royal Statistical Society. Series A (General), 134(4): 685, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343673.

Reid:1971:BRB

[223] David J. Reid. Book review: Modern Statistics for Business and Economics, by W. L. L’Esperance. Journal of the Royal Statistical Society. Series A (General), 134(4):686, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343674.

Green:1971:BRB

[224] Michael Green. Book review: Biproportional Matrices and Input-Output Change, by M. Bacharach. Journal of the Royal Statistical Society. Series A (General), 134(4):686–688, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343675.

Fagan:1971:BRB

[225] Paul B. Fagan. Book review: Speculation, Hedging and Commodity Price Forecasts, by W. C. Labys; C. Granger. Journal of the Royal Statistical Society. Series A (General), 134(4):688– 689, ???? 1971. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343676.

Jephcott:1971:BRB

[226] J. St. G. Jephcott. Book review: Stochastic Models of Buying Behavior, by W. F. Massey; D. B. Montgomery; D. G. Morrison. Journal of the Royal Statistical Society. Series A (General), 134(4):689, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343677.

Garside:1971:BRB

[227] M. J. Garside. Book review: Statistical Computation, by Roy C. Milton; John A. Nelder. Journal of the Royal Statistical Society. Series A (General), 134 (4):690–691, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343678.

Hopes:1971:BRB

[228] R. F. A. Hopes. Book review: Models of Manpower Systems, by A. R. Smith. Journal of the Royal Statistical Society. Series A (General), 134 (4):691–692, ???? 1971. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343679.

Goldstein:1971:BRB

[229] H. Goldstein. Book review: Longitudinal Studies and the Social Sciences, by W. D. Wall; H. L. Williams. Journal of the Royal Statistical Society. Series A (General), 134(4):692–693, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343680.

Anonymous:1971:CNd

[230] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 134(4):694–698, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343681.

Anonymous:1971:RPd

[231] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 134(4):699–713, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343682.

Anonymous:1971:LAd

[232] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 134(4):714–716, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343683.

Doll:1971:CAD

[233] R. Doll. Corrigendum: The age distribution of cancer: Implications for models of carcinogenesis. Journal of the Royal Statistical Society. Series A (General), 134(4):716, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343684. See [151].

Anonymous:1971:RIA

[234] Anonymous. Regulations of the industrial applications section. Journal of the Royal Statistical Society. Series A (General), 134(4):717–718, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343685.

Anonymous:1971:BMd

[235] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 134(4):i–iii, ???? 1971. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343686.

Anonymous:1972:VIa

[236] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 135(1):??, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345035.

Anonymous:1972:FMa

[237] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 135(1):??, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345036.

Barnard:1972:US

[238] G. A. Barnard. The unity of statistics. Journal of the Royal Statistical Society. Series A (General), 135(1):1–15, ???? 1972. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345037.

Parker:1972:UMC

[239] J. B. Parker. The use of the Monte Carlo method for solving largescale problems in neutronics. Journal of the Royal Statistical Society. Series A (General), 135(1):16–43, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345038.

Berlekamp:1972:SCT

[240] E. R. Berlekamp. A survey of coding theory. Journal of the Royal Statistical Society. Series A (General), 135(1):44– 73, ???? 1972. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345039.

Orr:1972:DTP

[241] Lea Orr. The dependence of transition proportions in the education system on observed social factors and school characteristics. Journal of the Royal Statistical Society. Series A (General), 135 (1):74–95, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345040.

Khamis:1972:NSI

[242] Salem H. Khamis. A new system of index numbers for national and international purposes. Journal of the Royal Statistical Society. Series A (General), 135(1):96–121, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345041.

Miller:1972:MEC

[243] W. L. Miller. Measures of electoral change using aggregate data. Journal of the Royal Statistical Society. Series A (General), 135(1):122–142, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345042.

Pearson:1972:SVC

[244] E. S. Pearson. “Statistics — a vocational or a cultural study?” Some comments on B. L. Welch’s paper and the discussion which followed. Journal of the Royal Statistical Society. Series A (General), 135(1):143–146, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345043. See [99].

Yates:1972:BRB

[245] F. Yates. Book review: Statistical Papers of George Udny Yule (1871–1951), by Alan Stuart; Maurice G. Kendall. Journal of the Royal Statistical Society. Series A (General), 135(1):147– 148, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345044.

Ashford:1972:BRBa

[246] J. R. Ashford. Book review: Probit Analysis, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 135(1):148, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345045.

Bailey:1972:BRB

[247] B. J. R. Bailey. Book review: An Introduction to the Theory of Statistics, by R. L. Plackett. Journal of the Royal Statistical Society. Series A (General), 135(1):149–150, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345046.

Silvey:1972:BRBa

[248] S. R. Silvey. Book review: Linear Models, by S. R. Searl. Journal of the Royal Statistical Society. Series A (General), 135(1):150–151, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345047.

Newbold:1972:BRB

[249] P. Newbold. Book review: Spectral Methods in Econometrics, by George S. Fishman. Journal of the Royal Statistical Society. Series A (General), 135 (1):151–152, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345048.

Kingman:1972:BRBa

[250] J. F. C. Kingman. Book review: Dynamic Probabilistic Systems. Volume 1: Markov Models, by Ronald A. Howard; Dynamic Probabilistic Systems. Volume 2: Semi-Markov and Decision Processes, by Ronald A. Howard. Journal of the Royal Statistical Society. Series A (General), 135(1):152–153, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345049.

Jackson:1972:BRB

[251] P. H. Jackson. Book review: Statistical Analysis in Psychology and Education, by George A. Ferguson. Journal of the Royal Statistical Society. Series A (General), 135(1):153–154, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345050.

Wilkinson:1972:BRBa

[252] R. K. Wilkinson. Book review: Contemporary Problems in Statistics: A Book of Readings for the Behavioural Sciences, by Lieberman. Journal of the Royal Statistical Society. Series A (General), 135(1):154, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345051.

Wilkinson:1972:BRBb

[253] R. K. Wilkinson. Book review: Applied Business Statistics — An Elementary Approach, by Elam E. McElroy. Journal of the Royal Statistical Society. Series A (General), 135(1):154–155, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345052.

Prais:1972:BRB

[254] Z. Prais. Book review: Business and Economic Statistics: A Programmed Introduction, by Freeman F. Elzey. Journal of the Royal Statistical Society. Series A (General), 135(1):155– 156, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345053.

Lewes:1972:BRBa

[255] F. M. M. Lewes. Book review: Attribute Sampling: Tables and Explanations. Tables for Determining Confidence Limits and Sampling Size Based on Close Approximations of the Binomial Distribution, by Herman Burstein. Journal of the Royal Statistical Society. Series A (General), 135(1):156, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345054.

Sargan:1972:BRB

[256] J. D. Sargan. Book review: Predictability of Stock Market Prices, by C. W. J. Granger; Oskar Morgenstern. Journal of the Royal Statistical Society. Series A (General), 135(1):156–157, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345055.

Wilkinson:1972:BRBc

[257] R. K. Wilkinson. Book review: Elementary Survey Sampling, by William Mendenhall; Lyman Ott; Richard L. Scheaffer. Journal of the Royal Statistical Society. Series A (General), 135 (1):157, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345056.

Healy:1972:BRBa

[258] M. J. R. Healy. Book review: Analysis and Design of Certain Quantitative Multiresponse Experiments, by S. N. Roy; R. Gnanadesikan; J. N. Srivastava; E. B. Fowlkes,; Elisa T. Lee. Journal of the Royal Statistical Society. Series A (General), 135(1):157–158, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345057.

Fyfe:1972:BRB

[259] D. J. Fyfe. Book review: A Book of Splines, by Arthur Sard; Sol Weintraub. Journal of the Royal Statistical Society. Series A (General), 135(1): 158–159, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345058.

Routh:1972:BRB

[260] Guy Routh. Book review: British Labour Statistics: Historical Abstract 1886–1968, by Department of Employment. Journal of the Royal Statistical Society. Series A (General), 135(1): 159–161, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345059.

Kemp:1972:BRBa

[261] C. D. Kemp. Book review: Accident Proneness: Research in the Occurrence, Causation and Prevention of Road Accidents, by Lynette Shaw; Herbert Sichel. Journal of the Royal Statistical Society. Series A (General), 135 (1):162–163, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345060.

Wilson:1972:BRBa

[262] Alan G. Wilson. Book review: Regional Variations in Britain: Studies in Economic and Social Geography, by B. E. Coates; E. M. Rawstron. Journal of the Royal Statistical Society. Series A (General), 135(1):163–164, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345061.

Flowerdew:1972:BRB

[263] A. D. J. Flowerdew. Book review: Combinatorial Programming, Spatial Analysis and Planning, by Allen J. Scott. Journal of the Royal Statistical Society. Series A (General), 135(1): 164–165, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345062.

Basilevsky:1972:BRB

[264] A. Basilevsky. Book review: Théorie des Jeux et Allocation des Ressources, Monnaie et Croissance. Système Linéaire de Dépenses. Économie Agricole, by R. Roy. Journal of the Royal Statistical Society. Series A (General), 135(1):165–166, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2345063.

Peters:1972:BRB

[265] G. H. Peters. Book review: Cost-Benefit Analysis, by D. W. Pearce. Journal of the Royal Statistical Society. Series A (General), 135(1):166– 167, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345064.

Williamson:1972:BRB

[266] John Williamson. Book review: The Responsiveness of Demand Policies to Balance of Payments: Postwar Patterns, by Michael Michaely. Journal of the Royal Statistical Society. Series A (General), 135(1):167–168, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345065.

Stapleton:1972:BRB

[267] R. C. Stapleton. Book review: Capital Investment Appraisal, by C. J. Hawkins; D. W. Pearce. Journal of the Royal Statistical Society. Series A (General), 135(1):168, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345066.

Corley:1972:BRB

[268] T. A. B. Corley. Book review: The Measurement of Fiscal Policy: An Analysis of Tax Systems in Terms of the Political Distinction between “Right” and “Left”, by Barry Bracewell-Milnes. Journal of the Royal Statistical Society. Series A (General), 135(1):169, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345067.

Casson:1972:BRB

[269] M. C. Casson. Book review: Economies of Scale and the Form of the Production Function: An Econometric Study of Norwegian Manufacturing Establishment Data, by Z. Griliches; V. Ringstad. Journal of the Royal Statistical Society. Series A (General), 135 (1):170–171, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345068.

Wilkinson:1972:BRBd

[270] R. K. Wilkinson. Book review: People in Poverty: A Melbourne Survey, by Ronald F. Henderson; Alison Harcourt; R. J. A. Harper. Journal of the Royal Statistical Society. Series A (General), 135(1):171, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345069.

Maynard:1972:BRB

[271] Alan Maynard. Book review: Financing Medical Education. An Analysis of Alternative Policies and Mechanisms. A General Report Prepared for the Carnegie Commission on Higher Education and the Commonwealth Fund, by Rashi Fein; Gerald I. Weber. Journal of the Royal Statistical Society. Series A (General), 135(1):172, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345070.

Anonymous:1972:CNa

[272] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 135(1):173–176, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345071.

Anonymous:1972:RPa

[273] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 135(1):177–181, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345072.

Anonymous:1972:LAa

[274] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 135(1):182–183, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345073.

Nicholson:1972:CDI

[275] R. J. Nicholson and N. Topham. Corrigendum: The determinants of investment in housing by local authorities: An econometric approach. Journal of the Royal Statistical Society. Series A (General), 135(1):183, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345074. See [173].

Anonymous:1972:BMa

[276] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 135(1):i–ii, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345075.

Anonymous:1972:FMb

[277] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 135(2):??, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344316.

Peto:1972:AER

[278] Richard Peto and Julian Peto. Asymptotically efficient rank invariant test procedures. Journal of the Royal Statistical Society. Series A (General), 135 (2):185–207, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344317.

Heathfield:1972:MCU

[279] David F. Heathfield. The measurement of capital usage using electricity consumption data for the U.K. Journal of the Royal Statistical Society. Series A (General), 135(2):208–220, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344318.

Holder:1972:CP

[280] R. L. Holder and F. Downton. Casino pontoon. Journal of the Royal Statistical Society. Series A (General), 135 (2):221–241, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344319.

Clowes:1972:DMA

[281] G. A. Clowes. A dynamic model for the analysis of labour turnover. Journal of the Royal Statistical Society. Series A (General), 135(2):242–256, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344320.

Lancaster:1972:SMD

[282] Tony Lancaster. A stochastic model for the duration of a strike. Journal of the Royal Statistical Society. Series A (General), 135(2):257–271, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344321.

Lindley:1972:BRB

[283] D. V. Lindley. Book review: The Theory of Statistical Inference, by Shelemyahu Zacks. Journal of the Royal Statistical Society. Series A (General), 135 (2):272–273, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344322.

Lewes:1972:BRBb

[284] F. M. M. Lewes. Book review: Making Decisions, by Dennis Lindley. Journal of the Royal Statistical Society. Series A (General), 135(2):273–274, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344323.

Tweedie:1972:BRB

[285] M. C. K. Tweedie. Book review: Statistical Methods in Medical Research, by P. Armitage. Journal of the Royal Statistical Society. Series A (General), 135 (2):274–275, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344324.

Armitage:1972:BRBa

[286] P. Armitage. Book review: Principles of Medical Statistics, by Austin Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 135 (2):275, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344325.

Silvey:1972:BRBb

[287] S. D. Silvey. Book review: Mathematical Statistics, by John E. Freund. Journal of the Royal Statistical Society. Series A (General), 135(2):276, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344326.

Haley:1972:BRB

[288] K. B. Haley. Book review: Optimization Under Constraints: Theory and Applications of Nonlinear Programming, by Peter Whittle. Journal of the Royal Statistical Society. Series A (General), 135(2):276, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344327.

Oliver:1972:BRBa

[289] F. R. Oliver. Book review: Let’s Look at the Figures, by D. J. Bartholomew; E. E. Bassett. Journal of the Royal Statistical Society. Series A (General), 135(2):277, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344328.

Kemp:1972:BRBb

[290] D. J. Kemp. Book review: A Computer-Assisted Approach to Elementary Statistics: Examples and Problems, by William Bulgren. Journal of the Royal Statistical Society. Series A (General), 135(2):277–278, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344329.

Friend:1972:BRB

[291] D. V. Friend. Book review: Statistical Analysis for Business: A Conceptual Approach, by Lee H. Smith; Donald R. Williams. Journal of the Royal Statistical Society. Series A (General), 135(2):278, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344330.

Silvey:1972:BRBc

[292] S. D. Silvey. Book review: Probability and Statistics: Models for Research, by Daniel E. Bailey. Journal of the Royal Statistical Society. Series A (General), 135(2):279, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344331.

Rahman:1972:BRB

[293] N. A. Rahman. Book review: The Language of Basic Statistics. A Participant Learning Programme for Students in Education and the Social Sciences, by John Newson; Michael Matthews. Journal of the Royal Statistical Society. Series A (General), 135(2):279– 281, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344332.

Hansford-Miller:1972:BRBa

[294] F. H. Hansford-Miller. Book review: Statistics: The Arithmetic of Decision Making, by J. M. Leonard. Journal of the Royal Statistical Society. Series A (General), 135(2):281–282, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344333.

Healy:1972:BRBb

[295] M. J. R. Healy. Book review: APL/ 360 — An Interactive Approach, by L. Gilman; A. J. Rose. Journal of the Royal Statistical Society. Series A (General), 135(2):282–283, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344334.

Goldstein:1972:BRB

[296] H. Goldstein. Book review: An Introduction to Data Management in the Behavioral and Social Sciences, by S. Blackman; K. M. Goldstein. Journal of the Royal Statistical Society. Series A (General), 135(2):283, ????

1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344335.

Granger:1972:BRB

[297] C. W. J. Granger. Book review: Cyclical Analysis of Time Series: Selected Procedures and Computer Programs, by Gerhard Bry; Charlotte Boschan. Journal of the Royal Statistical Society. Series A (General), 135(2):284, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344336.

Spitz:1972:BRB

[298] Kathleen E. Spitz. Book review: Elementary Survey Analysis, by James A. Davis. Journal of the Royal Statistical Society. Series A (General), 135(2): 285, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344337.

Edwards:1972:BRB

[299] A. W. F. Edwards. Book review: Probability Models and Statistical Methods in Genetics, by Regina C. Elandt-Johnson. Journal of the Royal Statistical Society. Series A (General), 135 (2):285–286, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344338.

Hansford-Miller:1972:BRBb

[300] F. H. Hansford-Miller. Book review: Regional Input-Output Study: Recollections, Reflections and Diverse Notes on The Philadelphia Experience, by Walter Isard; Thomas W. Langford. Journal of the Royal Statistical Society. Series A (General), 135(2):286–287, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344339.

Curnow:1972:BRB

[301] R. N. Curnow. Book review: Statistical Indicators for the Planning and Evaluation of Public Health Programmes. Fourteenth Report of the WHO Expert Committee on Health Statistics, by World Health Organization. Journal of the Royal Statistical Society. Series A (General), 135(2):287–288, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344340.

Howe:1972:BRB

[302] M. Howe. Book review: The Ownership Income of Management, by W. G. Lewellen. Journal of the Royal Statistical Society. Series A (General), 135 (2):288–290, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344341.

Anonymous:1972:CNb

[303] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 135(2):291–294, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344342.

Fowler:1972:RBA

[304] R. F. Fowler. R. B. Ainsworth, C.B.E., M.C. Journal of the Royal Statistical Society. Series A (General), 135(2): 295, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344343.

Anonymous:1972:RPb

[305] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 135(2):296–304, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344344.

Anonymous:1972:LAb

[306] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 135(2):305–306, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344345.

Anonymous:1972:BMb

[307] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 135(2):i–ii, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344346.

Anonymous:1972:FMc

[308] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 135(3):??, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344610.

Orton:1972:DSS

[309] C. R. Orton. The development of stratified sampling methods for the agricultural census of England and Wales. Journal of the Royal Statistical Society. Series A (General), 135(3):307– 335, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344611.

Downton:1972:BGG

[310] F. Downton and R. L. Holder. Banker’s games and the Gaming Act 1968. Journal of the Royal Statistical Society. Series A (General), 135(3):336–364, ????

1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344612.

Taylor:1972:PDH

[311] A. H. Taylor. The proportional decline hypothesis in English elections. Journal of the Royal Statistical Society. Series A (General), 135(3):365–369, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344613.

Nelder:1972:GLM

[312] J. A. Nelder and R. W. M. Wedderburn. Generalized linear models. Journal of the Royal Statistical Society. Series A (General), 135(3):370–384, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344614.

Ahamad:1972:NSA

[313] B. Ahamad and K. F. N. Scott. A note on sensitivity analysis in manpower forecasting. Journal of the Royal Statistical Society. Series A (General), 135(3):385–392, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344615.

Neave:1972:OSA

[314] Henry R. Neave. Observations on “Spectral analysis of short series — a simulation study” by Granger and Hughes. Journal of the Royal Statistical Society. Series A (General), 135 (3):393–405, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344616. See

Bee:1972:SSS

[315] Ronald E. Bee. A statistical study of the Sinai Pericope. Journal of the Royal Statistical Society. Series A (General), 135(3):406–421, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344617.

Cox:1972:BRB

[316] Joan G. Cox. Book review: Qualified Manpower and Economic Performance: An Inter-Plant Study in the Electrical Engineering Industry. Journal of the Royal Statistical Society. Series A (General), 135(3):422–427, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344618.

Bowers:1972:BRB

[317] J. K. Bowers. Book review: Regional and Urban Economics: Operational Methods, by J. H. P. Paelinck; A. Fleischer; N. N. Nekrasov. Journal of the Royal Statistical Society. Series A (General), 135(3):428–430, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344619.

Kingman:1972:BRBb

[318] J. F. C. Kingman. Book review: An Introduction to Probability Theory and its Applications, by W. Feller. Journal of the Royal Statistical Society. Series A (General), 135(3):430, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344620.

Nelder:1972:BRB

[319] J. A. Nelder. Book review: Probability, Statistics and Data Analysis, by Oscar Kempthorne; Leroy Folks. Journal of the Royal Statistical Society. Series A (General), 135(3):430–431, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344621.

Barnett:1972:BRBa

[320] V. D. Barnett. Book review: Foundations of Statistical Inference. Proceedings of the Symposium on the Foundations of Statistical Inference Prepared Under the Auspices of the Rene Descartes Foundation and Held at The Department of Statistics, University of Waterloo, Ontario, Canada, from March 31 to April 9, 1970, by V. P. Godambe; D. A. Sprott. Journal of the Royal Statistical Society. Series A (General), 135(3):431–433, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344622.

Welch:1972:BRB

[321] B. L. Welch. Book review: Distributions in Statistics: Continuous Univariate Distributions, by Norman L. Johnson; Samuel Kotz. Journal of the Royal Statistical Society. Series A (General), 135(3):433, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344623.

Nixon:1972:BRB

[322] Carol M. Nixon. Book review: Probability and Statistics, by Harry Lass; Peter Gottlieb. Journal of the Royal Statistical Society. Series A (General), 135(3):433–434, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344624.

Chatfield:1972:BRB

[323] C. Chatfield. Book review: Introduction to Probability and Statistics, by William Mendenhall. Journal of the Royal Statistical Society. Series A (General), 135(3):434–435, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344625.

Lutz:1972:BRB

[324] W. Lutz. Book review: Statistische Auswertungsmethoden. Dritte, Neubearbeitete und Erweiterte Auflage mit Neuer Bibliographie, by L. Sachs. Journal of the Royal Statistical Society. Series A (General), 135(3):435– 436, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344626.

Vesselo:1972:BRB

[325] I. R. Vesselo. Book review: Advanced General Statistics, by B. C. Erricker. Journal of the Royal Statistical Society. Series A (General), 135(3):436– 437, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344627.

Bunday:1972:BRB

[326] B. D. Bunday. Book review: Exercises and Worked Examples in Statistics, by G. E. Skipworth. Journal of the Royal Statistical Society. Series A (General), 135(3):437, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344628.

Reynolds:1972:BRB

[327] John F. Reynolds. Book review: Applications of Queueing Theory, by G.

F. Newell. Journal of the Royal Statistical Society. Series A (General), 135 (3):437–438, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344629.

Smith:1972:BRBa

[328] T. M. F. Smith. Book review: Generalized Inverse Matrices with Applications to Statistics, by R. M. Pringle; A. A. Rayner. Journal of the Royal Statistical Society. Series A (General), 135 (3):438–439, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344630.

Healy:1972:BRBc

[329] M. J. R. Healy. Book review: Generalized Inverse of Matrices and its Applications, by C. R. Rao; S. K. Mitra. Journal of the Royal Statistical Society. Series A (General), 135(3):439, ???? 1972. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344631.

Clarke:1972:BRB

[330] G. M. Clarke. Book review: Constructions and Combinatorial Problems in Design of Experiments, by Damaraju Raghavarao. Journal of the Royal Statistical Society. Series A (General), 135 (3):440, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344632.

Vajda:1972:BRB

[331] S. Vajda. Book review: Some Topics in Two-Person Games, by T. Parthasarathy; T. E. S. Raghavan. Journal of the Royal Statistical Society. Series A (General), 135(3):440– 441, ???? 1972. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344633.

Hills:1972:BRB

[332] M. Hills. Book review: Factor Analysis as A Statistical Method, by D. N. Lawley; A. E. Maxwell. Journal of the Royal Statistical Society. Series A (General), 135(3):441–442, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344634.

Maxwell:1972:BRB

[333] A. E. Maxwell. Book review: Prediction Methods in Criminology Including a Prediction Study of Young Men on Probation, by Frances H. Simon. Journal of the Royal Statistical Society. Series A (General), 135(3):442–443, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344635.

Oliver:1972:BRBb

[334] F. R. Oliver. Book review: Bayesian Full Information Structural Analysis with an Application to the Study of the Belgian Beef Market, by J.-A. Morales. Journal of the Royal Statistical Society. Series A (General), 135(3):443– 444, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344636.

Cannon:1972:BRB

[335] C. M. Cannon. Book review: The Logic of British and American Industry. A Realistic Analysis of Economic Structure and Government, by P. Sargant Florence. Journal of the Royal Statistical Society. Series A (General), 135 (3):444–445, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344637.

Oliver:1972:BRBc

[336] F. R. Oliver. Book review: Take-Overs: Their Relevance to the Stock Market and the Theory of the Firm, by Ajit Singh. Journal of the Royal Statistical Society. Series A (General), 135(3): 445–446, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344638.

Livingstone:1972:BRB

[337] Ian Livingstone. Book review: The Role of Investment in the Development of Fiji, by Michael Ward. Journal of the Royal Statistical Society. Series A (General), 135(3):446–448, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344639.

Grebenik:1972:BRB

[338] E. Grebenik. Book review: Population Growth and Economic Development since 1750, by H. J. Habakkuk. Journal of the Royal Statistical Society. Series A (General), 135(3):448– 449, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344640.

Armitage:1972:BRBb

[339] P. Armitage. Book review: Oral Health Surveys: Basic Methods. Journal of the Royal Statistical Society. Series A (General), 135(3):449, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344641.

Wilson:1972:BRBb

[340] W. A. Wilson. Book review: Epidémiologie de la Tuberculose, et Défaillances de la Lutte Antituberculeuse chez l’Enfant, by A. Lotte; S. Perdrizet; F. Hatton. Journal of the Royal Statistical Society. Series A (General), 135(3):449–450, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344642.

Nicholson:1972:BRB

[341] R. J. Nicholson. Book review: Second Abstract of British Historical Statistics, by B. R. Mitchell; H. G. Jones. Journal of the Royal Statistical Society. Series A (General), 135(3):450, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344643.

Anonymous:1972:CNc

[342] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 135(3):451–454, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344644.

Anonymous:1972:RPc

[343] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 135(3):455–461, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344645.

Lindley:1972:JS

[344] D. V. Lindley. L. J. Savage. Journal of the Royal Statistical Society. Series A (General), 135(3):462–463, ????

1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344646.

Anonymous:1972:GJ

[345] Anonymous. Gunnar Jahn. Journal of the Royal Statistical Society. Series A (General), 135(3):463, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344647.

Anonymous:1972:LAc

[346] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 135(3):464–466, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344648.

Anonymous:1972:BMc

[347] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 135(3):i–iv, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344649.

Anonymous:1972:VIb

[348] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 135(4):635– 638, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344677.

Anonymous:1972:FMd

[349] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 135(4):??, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344678.

Abelson:1972:RSR

[350] P. W. Abelson and A. D. J. Flowerdewe. Roskill’s successful recommendation. Journal of the Royal Statistical Society. Series A (General), 135 (4):467–510, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344679.

Stone:1972:MEM

[351] Richard Stone. A Markovian education model and other examples linking social behaviour to the economy. Journal of the Royal Statistical Society. Series A (General), 135(4):511–544, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344680.

Tippett:1972:ARS

[352] L. H. C. Tippett. Annals of the Royal Statistical Society 1934–71. Journal of the Royal Statistical Society. Series A (General), 135(4):545–568, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344681.

Lovell:1972:HWW

[353] H. G. Lovell. The heights and weights of West Indian children. Journal of the Royal Statistical Society. Series A (General), 135(4):569–578, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344682.

Anonymous:1972:POH

[354] Anonymous. Proceedings of the one hundred and thirty-eighth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 135 (4):579, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344683.

Anonymous:1972:ARC

[355] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 135 (4):580–603, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344684.

Wold:1972:BRB

[356] H. Wold. Book review: Scientific Truth and Statistical Method, by M. Boldrini; Ruth Kendall. Journal of the Royal Statistical Society. Series A (General), 135(4):604–605, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344685.

Priestley:1972:BRB

[357] M. B. Priestley. Book review: The Statistical Analysis of Time Series, by T. W. Anderson. Journal of the Royal Statistical Society. Series A (General), 135 (4):605–606, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344686.

Ashford:1972:BRBb

[358] J. R. Ashford. Book review: Lectures on Biostatistics: An Introduction to Statistics with Applications in Biology and Medicine, by D. Colquhoun. Journal of the Royal Statistical Society. Series A (General), 135(4):606, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344687.

Barnett:1972:BRBb

[359] V. D. Barnett. Book review: Order Statistics, by H. A. David. Journal of the Royal Statistical Society. Series A (General), 135(4):607, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344688.

Bibby:1972:BRB

[360] J. Bibby. Book review: Generalized Inverse Matrices, by Thomas L. Boullion; Patrick L. Odell. Journal of the Royal Statistical Society. Series A (General), 135(4):608–609, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344689.

Brown:1972:BRB

[361] B. M. Brown. Book review: Great Expectations: The Theory of Optimal Stopping, by Y. S. Chow; Herbert Robbins; David Siegmund. Journal of the Royal Statistical Society. Series A (General), 135(4):610, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344690.

Williams:1972:BRB

[362] D. A. Williams. Book review: Multivariate Morphometrics, by R. E. Blackith; R. A. Reyment. Journal of the Royal Statistical Society. Series A (General), 135(4):610–611, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344691.

Sneath:1972:BRB

[363] P. H. A. Sneath. Book review: Mathematics in the Archaeological and Historical Sciences, by F. R. Hodson; D. G. Kendall; P. Tautu. Journal of the Royal Statistical Society. Series A (General), 135(4):611–612, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344692.

Bather:1972:BRB

[364] J. A. Bather. Book review: Statistics II, by J. Aitchison. Journal of the Royal Statistical Society. Series A (General), 135(4):612, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344693.

Barton:1972:BRB

[365] D. E. Barton. Book review: A First Course in Statistics, by F. N. David. Journal of the Royal Statistical Society. Series A (General), 135(4):612– 613, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344694.

Brook:1972:BRB

[366] D. Brook. Book review: Theoretical Statistics — Basic Ideas, by Stanley N. Collings. Journal of the Royal Statistical Society. Series A (General), 135 (4):613, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344695.

Moore:1972:BRB

[367] P. G. Moore. Book review: The Analysis of Mortality and Other Acturial Statistics, by B. Benjamin; H. W. Haycocks. Journal of the Royal Statistical Society. Series A (General), 135(4): 613–614, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344696.

Hollingsworth:1972:BRB

[368] T. H. Hollingsworth. Book review: Demographic Estimation for Developing Societies. A Manual of Techniques for the Detection and Reduction of Errors in Demographic Data, by Norman Carrier; John Hobcraft. Journal of the Royal Statistical Society. Series A (General), 135(4):614–616, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344697.

Moss:1972:BRB

[369] L. Moss. Book review: Survey Methods in Social Investigation, by C. A. Moser; G. Kalton. Journal of the Royal Statistical Society. Series A (General), 135 (4):616–617, ???? 1972. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344698.

Tippett:1972:BRB

[370] L. H. C. Tippett. Book review: Applications of “Tippett” Snap Readings in Social and Commercial Research. Journal of the Royal Statistical Society. Series A (General), 135(4):617–618, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344699.

Lowe:1972:BRB

[371] E. A. Lowe. Book review: A Guide to Stock Control, by A. Battersby. Journal of the Royal Statistical Society. Series A (General), 135(4):618–619, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344700.

Tomlinson:1972:BRB

[372] R. C. Tomlinson. Book review: Developments in Operations Research, Volumes I and II, by Benjamin Avi-Itzhak. Journal of the Royal Statistical Society. Series A (General), 135(4):619, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344701.

Bulmer:1972:BRB

[373] M. G. Bulmer. Book review: The Genetics of Human Populations, by L. L. Cavalli-Sforza; W. F. Bodmer. Journal of the Royal Statistical Society. Series A (General), 135(4):619–621, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344702.

Gani:1972:BRB

[374] J. Gani. Book review: Platelet Kinetics: Radioisotopic, Cytological, Mathematical and Clinical Aspects, by J. M. Paulus; R. H. Aster; H. Breny. Journal of the Royal Statistical Society. Series A (General), 135(4):621, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344703.

Smith:1972:BRBb

[375] Cedric A. B. Smith. Book review: Polymorphisms with Linked Loci, by V. Arunachalam; A. R. G. Owen. Journal of the Royal Statistical Society. Series A (General), 135(4):621–622, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344704.

Kalton:1972:BRB

[376] G. Kalton. Book review: Variations in Services for the Aged: A Causal Analysis, by Bleddyn Davies; Andrew Barton; Ian McMillan; Valerie Williamson. Journal of the Royal Statistical Society. Series A (General), 135(4):622– 623, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344705.

Sumner:1972:BRB

[377] M. T. Sumner. Book review: Issues in Fiscal and Monetary Policy: The Electric Economist Views the Controversy, by James J. Diamond; Paul A. Samuelson; Tilford C. Gaines; Lawrence R. Klein; Arthur M. Okum; Robert Eisner. Journal of the Royal Statistical Society. Series A (General), 135(4):623– 624, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344706.

Anonymous:1972:CNd

[378] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 135(4):625–630, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344707.

Anonymous:1972:RPd

[379] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 135(4):631–634, ???? 1972. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344708.

Anonymous:1972:BMd

[380] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series A (General), 135(4):i–ii, ???? 1972. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344709.

Anonymous:1973:VIa

[381] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 136(1):??, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344416.

Anonymous:1973:FMa

[382] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 136(1):??, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344417.

Wilson:1973:SDM

[383] J. Harold Wilson. Statistics and decision-making in government — Bradshaw revisited. Journal of the Royal Statistical Society. Series A (General), 136(1):1–20, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344418.

Hampton:1973:SPM

[384] J. M. Hampton, P. G. Moore, and H. Thomas. Subjective probability and its measurement. Journal of the Royal Statistical Society. Series A (General), 136(1):21–42, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344419.

Pickering:1973:IMC

[385] J. F. Pickering, J. A. Harrison, and C. D. Cohen. Identification and measurement of consumer confidence: Methodology and some preliminary results. Journal of the Royal Statistical Society. Series A (General), 136(1):43– 63, ???? 1973. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344420.

Anonymous:1973:F

[386] Anonymous. Foreword. Journal of the Royal Statistical Society. Series A (General), 136(1):64, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344421.

Bartholomew:1973:PET

[387] D. J. Bartholomew. Post-experience training for statisticians. Journal of the Royal Statistical Society. Series A (General), 136(1):65–70, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344422.

Healy:1973:VS

[388] M. J. R. Healy. The varieties of statistician. Journal of the Royal Statistical Society. Series A (General), 136(1):71– 74, ???? 1973. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344423.

Moser:1973:SGS

[389] C. A. Moser. Staffing in the government statistical service. Journal of the Royal Statistical Society. Series A (General), 136(1):75–88, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344424.

Moore:1973:RBS

[390] P. G. Moore. Relationship between statisticians and other numerate professions. Journal of the Royal Statistical Society. Series A (General), 136 (1):89–94, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344425.

Armstrong:1973:MTS

[391] William Armstrong. Management training for statisticians and opportunities to enter top management. Journal of the Royal Statistical Society. Series A (General), 136(1):95–99, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344426.

Lindley:1973:BRBa

[392] D. V. Lindley. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability. Volume I Theory of Statistics, by Lucien M. Le Cam; Jerzy Neyman; Elizabeth L. Scott. Journal of the Royal Statistical Society. Series A (General), 136(1):100–101, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344427.

Herne:1973:BRB

[393] H. Herne. Book review: Multivariate Data Analysis, by William W. Cooley; Paul R. Lohnes. Journal of the Royal Statistical Society. Series A (General), 136(1):101–103, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344428.

Williams:1973:BRBa

[394] D. Williams. Book review: Regenerative Phenomena, by J. F. C. Kingman. Journal of the Royal Statistical Society. Series A (General), 136(1):

103–105, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344429.

Eagleson:1973:BRB

[395] G. K. Eagleson. Book review: Brownian Motion and Diffusion, by David Freedman. Journal of the Royal Statistical Society. Series A (General), 136 (1):105–106, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344430.

Oliver:1973:BRBa

[396] F. R. Oliver. Book review: Rapid Statistical Calculations. A Collection of Distribution-Free and Easy Methods of Estimation and Testing, by M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 136 (1):106, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344431.

Godolphin:1973:BRB

[397] E. J. Godolphin. Book review: Introduction to Statistics, a Fresh Approach, by Gottfried E. Noether. Journal of the Royal Statistical Society. Series A (General), 136(1):106–107, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344432.

Rosenhead:1973:BRB

[398] Jonathan Rosenhead. Book review: Queueing Theory in OR, by E. Page. Journal of the Royal Statistical Society. Series A (General), 136(1):107– 108, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344433.

Champernowne:1973:BRB

[399] D. G. Champernowne. Book review: Decision and Organization. A Volume in Honor of Jacob Marschak, by C. B. McGuire; Roy Radner. Journal of the Royal Statistical Society. Series A (General), 136(1):108–109, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344434.

Fisk:1973:BRBa

[400] P. R. Fisk. Book review: Nonlinear Methods in Econometrics, by Stephen M. Goldfeld; Richard E. Quandt; Dennis E. Smallwood. Journal of the Royal Statistical Society. Series A (General), 136(1):109–110, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344435.

Barker:1973:BRB

[401] Terence S. Barker. Book review: The Analysis and Forecasting of the British Economy, by M. G. C. Surrey. Journal of the Royal Statistical Society. Series A (General), 136(1):110–112, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344436.

Granger:1973:BRB

[402] C. W. J. Granger. Book review: Econometric Models of Cyclical Behavior. Volumes I and II, by Bert G. Hickman. Journal of the Royal Statistical Society. Series A (General), 136(1): 112–113, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344437.

Forsyth:1973:BRB

[403] F. G. Forsyth. Book review: Price Indexes and Quality Change. Studies in New Methods of Measurement, by Zvi Griliches. Journal of the Royal Statistical Society. Series A (General), 136 (1):113–115, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344438.

Thatcher:1973:BRB

[404] A. R. Thatcher. Book review: The Analysis of Family Budgets, by S. J. Prais; H. S. Houthakker. Journal of the Royal Statistical Society. Series A (General), 136(1):115–116, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344439.

Spitz:1973:BRB

[405] Kathleen Spitz. Book review: Surveys of Personal Income in London, by R. U. Redpath; M. G. Powell; S. Kingaby. Journal of the Royal Statistical Society. Series A (General), 136(1):116, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344440.

Williams:1973:BRBb

[406] J. E. Williams. Book review: National Income, Expenditure and Output of the United Kingdom 1855–1965, by C. H. Feinstein. Journal of the Royal Statistical Society. Series A (General), 136 (1):116–117, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344441.

Benjamin:1973:BRB

[407] Bernard Benjamin. Book review: Population: Facts and Methods of Demography, by Nathan Keyfitz; Whilhelm Flieger. Journal of the Royal Statistical Society. Series A (General), 136 (1):117–118, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344442.

Talbot:1973:BRB

[408] Valerie Talbot. Book review: Postal Survey Methods, by Social and Community Planning Research; Sample Design and Selection. by Social and Community Planning Research; Interviewers’ Guidebook. by Social and Community Planning Research. Journal of the Royal Statistical Society. Series A (General), 136(1):118–119, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344443.

Hillmann:1973:BRB

[409] H. C. Hillmann. Book review: The Crisis in Planning, by Mike Faber; Dudley Seers. Journal of the Royal Statistical Society. Series A (General), 136 (1):119–120, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344444.

Goldstein:1973:BRB

[410] H. Goldstein. Book review: Variations in Children’s Services Among British Urban Authorities. A Causal Analysis, by Bleddyn Davies; Andrew Barton; Ian McMillan. Journal of the Royal Statistical Society. Series A (General), 136(1):120, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344445.

Paenson:1973:BRB

[411] I. Paenson. Book review: Russian– English, English–Russian Glossary of Statistical Terms, by M. G. Kendall; W. R. Buckland; Samuel Kotz. Journal of the Royal Statistical Society. Series A (General), 136(1):120–121, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344446.

Anonymous:1973:CNa

[412] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 136(1):122–123, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344447.

Anonymous:1973:RPa

[413] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 136(1):124–147, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344448.

Anonymous:1973:LAa

[414] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 136(1):148–152, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344449.

Anonymous:1973:BMa

[415] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 136(1):i–iv, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344450.

Anonymous:1973:FMb

[416] Anonymous. Front matter. Journal of the Royal Statistical Society.

Series A (General), 136(2):??, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345106.

Anonymous:1973:ENC

[417] Anonymous. Errata: Note concerning volume 135 (1972). Journal of the Royal Statistical Society. Series A (General), 136(2):??, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345107.

Evans:1973:ICT

[418] Dennis A. Evans. The influence of computers on the teaching of statistics. Journal of the Royal Statistical Society. Series A (General), 136(2):153– 190, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345108.

Mead:1973:UCT

[419] R. Mead and R. D. Stern. The use of a computer in the teaching of statistics. Journal of the Royal Statistical Society. Series A (General), 136(2):191– 225, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345109.

Hull:1973:UM

[420] J. Hull, P. G. Moore, and H. Thomas. Utility and its measurement. Journal of the Royal Statistical Society. Series A (General), 136(2):226–247, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345110.

Joseph:1973:HSD

[421] A. W. Joseph. Harmonizing statistical data. Journal of the Royal Statistical Society. Series A (General), 136 (2):248–254, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345111.

Allen:1973:SBL

[422] R. G. D. Allen. The statistician and The British Library of Political and Economic Science. Journal of the Royal Statistical Society. Series A (General), 136(2):255–257, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2345112.

Lindley:1973:BRBb

[423] D. V. Lindley. Book review: On the Mathematical Foundations of Probability, by Tue Tjur. Journal of the Royal Statistical Society. Series A (General), 136(2):258, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345113.

Lindley:1973:BRBc

[424] D. V. Lindley. Book review: Studies in Inductive Logic and Probability. Volume 1, by Rudolf Carnap; Richard C. Jeffrey. Journal of the Royal Statistical Society. Series A (General), 136 (2):259, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345114.

Healy:1973:BRBa

[425] M. J. R. Healy. Book review: Multivariate Analysis, by Anant M. Kshirsagar. Journal of the Royal Statistical Society. Series A (General), 136(2): 259–260, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345115.

Pridmore:1973:BRB

[426] W. A. Pridmore. Book review: Statistical Methods in Research and Production with Special Reference to the Chemical Industry, by Owen L. Davies; Peter L. Goldsmith. Journal of the Royal Statistical Society. Series A (General), 136(2):260–262, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345116.

Stone:1973:BRB

[427] M. Stone. Book review: Introductory Engineering Statistics, by Irwin Guttman; S. S. Wilks; J. Stuart Hunter. Journal of the Royal Statistical Society. Series A (General), 136 (2):262, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345117.

Upton:1973:BRB

[428] G. J. G. Upton. Book review: Statistics of Directional Data, by K. V. Mardia. Journal of the Royal Statistical Society. Series A (General), 136(2):262– 263, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345118.

Ilersic:1973:BRB

[429] A. R. Ilersic. Book review: Basic Statistics, by Dick A. Leabo; Richard D. Irwin. Journal of the Royal Statistical Society. Series A (General), 136 (2):263–264, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345119.

Holgate:1973:BRBa

[430] P. Holgate. Book review: Third Conference Advisory Group of Forest Statisticians, Jouy-en-Josas, 7–11 Septembre 1970, by L’Union Internationale des Organismes de Recherches Forestières. Journal of the Royal Statistical Society. Series A (General), 136 (2):264–265, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345120.

Mayes:1973:BRB

[431] D. G. Mayes. Book review: More Fortran Programs for Economists, by Lucy Joan Slater. Journal of the Royal Statistical Society. Series A (General), 136 (2):265–266, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345121.

Gower:1973:BRB

[432] J. C. Gower. Book review: Fortran Techniques, with Special Reference to Non-Numerical Applications, by A. Colin Day. Journal of the Royal Statistical Society. Series A (General), 136 (2):266–267, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345122.

Healy:1973:BRBb

[433] M. J. R. Healy. Book review: Biometrika Tables for Statisticians, by E. S. Pearson; H. O. Hartley. Journal of the Royal Statistical Society. Series A (General), 136(2):267, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345123.

Oliver:1973:BRBb

[434] F. R. Oliver. Book review: The Trace Econometric Model of the Canadian Economy, by Nanda K. Choudhry; Yehuda Kotowitz; John A. Sawyer; John W. L. Winder. Journal of the Royal Statistical Society. Series A (General), 136(2):268, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345124.

Bibby:1973:BRB

[435] John Bibby. Book review: Production Functions: An Integration of Micro and Macro, Short Run and Long Run Aspects, by Leif Johansen. Journal of the Royal Statistical Society. Series A (General), 136(2):269, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345125.

Coppock:1973:BRB

[436] J. T. Coppock. Book review: Empirical Models of Urban Land Use: Suggestions on Research Objectives and Organization, by James H. Brown; J. Royce Ginn; Franklin J. James; John F. Kain; Mahlon R. Straszheim. Journal of the Royal Statistical Society. Series A (General), 136(2):269–270, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345126.

Pyatt:1973:BRB

[437] F. G. Pyatt. Book review: Repeat-Buying: Theory and Applications, by A. S. C. Ehrenberg. Journal of the Royal Statistical Society. Series A (General), 136(2):270–271, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345127.

Hutchison:1973:BRB

[438] T. W. Hutchison. Book review: Man and the Social Sciences, by William A.

Robson. Journal of the Royal Statistical Society. Series A (General), 136 (2):271–272, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345128.

Sargan:1973:BRB

[439] J. D. Sargan. Book review: Frontiers of Quantitative Economics. Papers Invited for Presentation at the Econometric Society Winter Meetings, by Michael D. Intriligator. Journal of the Royal Statistical Society. Series A (General), 136(2):272–273, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345129.

Topham:1973:BRB

[440] Neville Topham. Book review: Uses of Economics, by G. D. N. Worswick. Journal of the Royal Statistical Society. Series A (General), 136(2):273– 274, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345130.

Shaw:1973:BRB

[441] R. W. Shaw. Book review: The Principles of Modern Management, by Ray Wild; Bryan Lowes. Journal of the Royal Statistical Society. Series A (General), 136(2):274, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345131.

Brothwell:1973:BRB

[442] J. F. Brothwell. Book review: The Theory of Futures Trading, by B. A. Goss. Journal of the Royal Statistical Society. Series A (General), 136(2):275, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345132.

Anonymous:1973:CNb

[443] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 136(2):276–281, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345133.

Anonymous:1973:RPb

[444] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 136(2):282–290, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345134.

Anonymous:1973:LAb

[445] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 136(2):291–292, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345135.

Anonymous:1973:NPM

[446] Anonymous. Notes for the preparation of mathematical papers. Journal of the Royal Statistical Society. Series A (General), 136(2):293–294, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345136.

Anonymous:1973:BMb

[447] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 136(2):i–vi, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345137.

Anonymous:1973:FMc

[448] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 136(3):??, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344993.

Chatfield:1973:BJS

[449] C. Chatfield and D. L. Prothero. Box– Jenkins seasonal forecasting: Problems in a case-study. Journal of the Royal Statistical Society. Series A (General), 136(3):295–336, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2344994. See comments [450].

Box:1973:SCP

[450] G. E. P. Box and G. M. Jenkins. Some comments on a paper by Chatfield and Prothero and on a review by Kendall. Journal of the Royal Statistical Society. Series A (General), 136 (3):337–352, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344995. See

[449, 458].

Lyttkens:1973:FPM

[451] Ejnar Lyttkens. The fix-point method for estimating interdependent systems with the underlying model specification. Journal of the Royal Statistical Society. Series A (General), 136(3): 353–394, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344996.

Lomnicki:1973:SAS

[452] Z. A. Lomnicki. Some aspects of the statistical approach to reliability. Journal of the Royal Statistical Society. Series A (General), 136(3):395–420, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344997.

Gardner:1973:UEE

[453] M. J. Gardner. Using the environment to explain and predict mortality. Journal of the Royal Statistical Society. Series A (General), 136(3):421–440, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344998.

Griffiths:1973:EMV

[454] D. A. Griffiths. The effect of measles vaccination on the incidence of measles in the community. Journal of the Royal Statistical Society. Series A (General), 136(3):441–449, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344999.

Kingman:1973:BRBa

[455] J. F. C. Kingman. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability. Volume II: Probability Theory, by Lucien M. Le Cam; Jerzy Neyman; Elizabeth L. Scott. Journal of the Royal Statistical Society. Series A (General), 136(3):450–451, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345000.

Healy:1973:BRBc

[456] M. J. R. Healy. Book review: Collected Papers of R. A. Fisher. Volume I, 1912–1924, by R. A. Bennett. Journal of the Royal Statistical Society. Series A (General), 136(3):451–452, ????

1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345001.

Barnett:1973:BRB

[457] V. Barnett. Book review: An Introduction to Bayesian Inference and Decision, by Robert L. Winkler. Journal of the Royal Statistical Society. Series A (General), 136(3):452–454, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345002.

Kendall:1973:BRB

[458] David Kendall. Book review: Approximating Countable Markov Chains, by David Freedman. Journal of the Royal Statistical Society. Series A (General), 136(3):454–456, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2345003. See comment [450].

Anderson:1973:BRB

[459] C. W. Anderson. Book review: Numerical Methods in Markov Chains and Bulk Queues, by T. P. Bagchi; J. G. C. Templeton. Journal of the Royal Statistical Society. Series A (General), 136 (3):456–457, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345004.

Mardia:1973:BRB

[460] K. V. Mardia. Book review: Families of Frequency Distributions, by J. K. Ord. Journal of the Royal Statistical Society. Series A (General), 136 (3):457–458, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345005.

Hinkley:1973:BRB

[461] David V. Hinkley. Book review: Robust Estimates of Location: Survey and Advances, by D. F. Andrews; P. J. Bickel; F. R. Hampel; P. J. Huber; W. H. Rogers; J. W. Tukey. Journal of the Royal Statistical Society. Series A (General), 136(3):458–459, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345006.

Holgate:1973:BRBb

[462] P. Holgate. Book review: Success Epochs in Bernoulli Trials (With Applications in Number Theory), by W. Vervaat. Journal of the Royal Statistical Society. Series A (General), 136 (3):459–460, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345007.

Vajda:1973:BRB

[463] S. Vajda. Book review: Stochastic Programming. Methods and Applications, by Jati K. Sengupta. Journal of the Royal Statistical Society. Series A (General), 136(3):460–461, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345008.

Finney:1973:BRB

[464] D. J. Finney. Book review: The Logit Transformation with Special Reference to its Uses in Bioassay, by Winifred D. Ashton. Journal of the Royal Statistical Society. Series A (General), 136 (3):461, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345009.

Holgate:1973:BRBc

[465] P. Holgate. Book review: A Survey of Lie Groups and Lie Algebras with Applications and Computational Methods, by Johan G. F. Belinfante; Bernard Kolman. Journal of the Royal Statistical Society. Series A (General), 136 (3):461–462, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345010.

Bartholomew:1973:BRB

[466] D. J. Bartholomew. Book review: Statistical Decomposition Analysis with Application in the Social and Administrative Sciences, by Henri Theil. Journal of the Royal Statistical Society. Series A (General), 136(3):462–463, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345011.

Chenhall:1973:BRB

[467] R. H. Chenhall. Book review: Risk in Business Decision, by Peter G. Moore. Journal of the Royal Statistical Society. Series A (General), 136(3):463– 464, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345012.

Davies:1973:BRB

[468] Hilda M. Davies. Book review: Practical Exercises in Probability and Statistics for Sixth Forms, Technical Colleges, Colleges of Education, Universities. With Answers and Hints on Solutions, by N. A. Rahman. Journal of the Royal Statistical Society. Series A (General), 136(3):464–465, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345013.

Thomas:1973:BRB

[469] C. J. Thomas. Book review: Introduction to Probability and Statistics, by Henry L. Alder; Edward B. Roessler. Journal of the Royal Statistical Society. Series A (General), 136(3):465– 466, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345014.

Tennant-Smith:1973:BRBa

[470] J. Tennant-Smith. Book review: Elements of Statistics, by E. W. Averill. Journal of the Royal Statistical Society. Series A (General), 136(3):466– 467, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345015.

Tennant-Smith:1973:BRBb

[471] J. Tennant-Smith. Book review: Introductory Statistics: A Service Course, by A. H. Pollard. Journal of the Royal Statistical Society. Series A (General), 136(3):467, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345016.

Hawkes:1973:BRB

[472] A. G. Hawkes. Book review: A Primer of Signal Detection Theory, by D. Mc-Nicol. Journal of the Royal Statistical Society. Series A (General), 136(3): 467–468, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345017.

Deane:1973:BRB

[473] Marjorie Deane. Book review: How to Find out About Statistics, by Gillian A. Burrington. Journal of the Royal Statistical Society. Series A (General), 136 (3):468–469, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345018.

Oldham:1973:BRB

[474] P. D. Oldham. Book review: A Dictionary of Statistical Terms, by Maurice G. Kendall; William R. Buckland. Journal of the Royal Statistical Society. Series A (General), 136(3):469– 470, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345019.

Peretz:1973:BRB

[475] Jane Peretz. Book review: A Quality Check on the 1966 Ten Per Cent Sample Census of England and Wales, by Percy Gray; Frances A. Gee. Journal of the Royal Statistical Society. Series A (General), 136(3):470–471, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345020.

Newlyn:1973:BRB

[476] W. T. Newlyn. Book review: The Frontiers of Development Studies, by Paul Streeten. Journal of the Royal Statistical Society. Series A (General), 136 (3):471–472, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345021.

McCarthy:1973:BRB

[477] M. J. McCarthy. Book review: International Mobility and Movement of Capital, by Fritz Machlup; Walter S. Salant; Lorie Tarshis. Journal of the Royal Statistical Society. Series A (General), 136 (3):473, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345022.

McCormick:1973:BRB

[478] B. McCormick. Book review: Economic Research: Retrospect and Prospect. Human Resources. Fiftieth Anniversary Colloquium VI, by National Bureau of Economic Research. Journal of the Royal Statistical Society. Series A (General), 136(3):473– 474, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345023.

Nicholson:1973:BRB

[479] R. J. Nicholson. Book review: Economic Research: Retrospect and Prospect. Quantitative Economic Research: Trends and Problems. Fiftieth Anniversary Colloquium VII, by Simon Kuznets. Journal of the Royal Statistical Society. Series A (General), 136 (3):474–475, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345024.

Royse:1973:BRB

[480] A. B. Royse. Book review: Key Variables in Social Research. Volume 1: Religion, Housing, Locality, by Elizabeth Gittus. Journal of the Royal Statistical Society. Series A (General), 136 (3):475–476, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345025.

Cairncross:1973:BRB

[481] Alec Cairncross. Book review: Migration and Urban Development. A Reappraisal of British and American Long Cycles, by Brinley Thomas. Journal of the Royal Statistical Society. Series A (General), 136(3):476–477, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345026.

Ashford:1973:BRB

[482] J. R. Ashford. Book review: Essays in the Economics of Health and Medical Care, by Victor R. Fuchs. Journal of the Royal Statistical Society. Series A (General), 136(3):477–478, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345027.

Cliff:1973:BRBa

[483] A. D. Cliff. Book review: Freight Flows and Spatial Aspects of the British Economy, by Michael D. I. Chisholm; Patrick M. O’Sullivan. Journal of the Royal Statistical Society. Series A (General), 136(3):478–479, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345028.

Bowe:1973:BRB

[484] Colette Bowe. Book review: Oligopoly and Conflict: A Dynamic Approach, by Michael Nicholson. Journal of the Royal Statistical Society. Series A (General), 136(3):479–480, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345029.

Anonymous:1973:CNc

[485] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 136(3):481–482, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345030.

Tinbergen:1973:PRF

[486] J. Tinbergen. Professor Ragnar Frisch. Journal of the Royal Statistical Society. Series A (General), 136(3):483, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345031.

Anonymous:1973:RPc

[487] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 136(3):484–500, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345032.

Anonymous:1973:LAc

[488] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 136(3):501–503, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345033.

Anonymous:1973:BMc

[489] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 136(3):i–iv, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345034.

Anonymous:1973:VIb

[490] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 136(4):640– 643, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344747.

Anonymous:1973:FMd

[491] Anonymous. Front matter. Journal of the Royal Statistical Society.

Series A (General), 136(4):??, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344748.

Jones:1973:PCS

[492] H. J. M. Jones, H. B. Lawson, and D. Newman. Population census: Some recent British developments in methodology. Journal of the Royal Statistical Society. Series A (General), 136(4): 505–538, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344749.

Moser:1973:UKO

[493] Claus A. Moser and I. B. Beesley. United Kingdom official statistics and the European Communities. Journal of the Royal Statistical Society. Series A (General), 136(4):539–582, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344750.

Moore:1973:SCP

[494] P. G. Moore. Security of the census of population. Journal of the Royal Statistical Society. Series A (General), 136 (4):583–596, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344751.

Anonymous:1973:PSG

[495] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 136 (4):597, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344752.

Anonymous:1973:POH

[496] Anonymous. Proceedings of the one hundred and thirty-ninth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 136 (4):598, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344753.

Anonymous:1973:ARC

[497] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 136 (4):599–620, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344754.

Reuter:1973:BRB

[498] G. E. H. Reuter. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability, Volume III: Probability Theory, by Lucein M. Le Cam; Jerzy Neyman; Elizabeth L. Scott. Journal of the Royal Statistical Society. Series A (General), 136(4):621–622, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344755.

Fisk:1973:BRBb

[499] P. R. Fisk. Book review: Distributions in Statistics: Continuous Multivariate Distributions, by Norman L. Johnson; Samuel Kotz. Journal of the Royal Statistical Society. Series A (General), 136 (4):622, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344756.

Renshaw:1973:BRB

[500] E. Renshaw. Book review: Stochastic Point Processes: Statistical Analysis, Theory, and Applications, by Peter A. W. Lewis. Journal of the Royal Statistical Society. Series A (General), 136(4):623–624, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344757.

Kingman:1973:BRBb

[501] J. F. C. Kingman. Book review: Martingales a Temps Discret, by Jacques Neveu. Journal of the Royal Statistical Society. Series A (General), 136 (4):624–625, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344758.

Neave:1973:BRB

[502] H. R. Neave. Book review: Digital Time Series Analysis, by Robert K. Otnes; Loren Enochson. Journal of the Royal Statistical Society. Series A (General), 136(4):625–626, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344759.

Williams:1973:BRBc

[503] D. Williams. Book review: Prediction Methods in Multiplicative Models, by R. Teekens. Journal of the Royal Statistical Society. Series A (General), 136 (4):626, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344760.

Brown:1973:BRB

[504] B. M. Brown. Book review: Branching Processes, by K. B. Athreya; P. E. Ney. Journal of the Royal Statistical Society. Series A (General), 136 (4):626–627, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344761.

Clarke:1973:BRB

[505] M. R. B. Clarke. Book review: Modèles et Méthodes de l’Analyse Factorielle, by J. Torrens-Ibern. Journal of the Royal Statistical Society. Series A (General), 136(4):627, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344762.

Pearce:1973:BRB

[506] S. C. Pearce. Book review: Fundamentals of Biometry, by L. N. Balaam. Journal of the Royal Statistical Society. Series A (General), 136(4):628, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344763.

Sutherland:1973:BRBa

[507] Ian Sutherland. Book review: Data Handling in Epidemiology, by W. W. Holland. Journal of the Royal Statistical Society. Series A (General), 136 (4):628–629, ???? 1973. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344764.

Barton:1973:BRB

[508] D. E. Barton. Book review: Introduction to Mathematical Techniques in Pattern Recognition, by Harry C. Andrews. Journal of the Royal Statistical Society. Series A (General), 136(4): 629–630, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344765.

Sutherland:1973:BRBb

[509] Ian Sutherland. Book review: Conflict Analysis, by Michael Nicholson. Journal of the Royal Statistical Society. Series A (General), 136(4):630–631, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344766.

Psacharopoulos:1973:BRB

[510] G. Psacharopoulos. Book review: The Effect of Education on Efficiency in Consumption, by Robert T. Michael. Journal of the Royal Statistical Society. Series A (General), 136(4):631– 632, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344767.

Williams:1973:BRBd

[511] Alan Williams. Book review: Economic Research: Retrospect and Prospect: Public Expenditure and Taxation. Journal of the Royal Statistical Society. Series A (General), 136(4):632–633, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344768.

Cliff:1973:BRBb

[512] A. D. Cliff. Book review: The Detroit Prototype of the NBER Urban Simulation Model, by Gregory K. Ingram; John F. Kain; J. Royce Ginn. Journal of the Royal Statistical Society. Series A (General), 136(4):634, ???? 1973. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344769.

Anonymous:1973:RPd

[513] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 136(4):635–637, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344770.

Anonymous:1973:LAd

[514] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 136(4):638–639, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344771.

Anonymous:1973:BMd

[515] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 136(4):i–ii, ???? 1973. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344772.

Anonymous:1974:VIa

[516] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 137(1):??, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345138.

Anonymous:1974:FMa

[517] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 137(1):i–iv, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345139.

Finney:1974:PDI

[518] D. J. Finney. Problems, data, and inference. Journal of the Royal Statistical Society. Series A (General), 137(1):1– 23, ???? 1974. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345140.

Anonymous:1974:CR

[519] Anonymous. Conference report. Journal of the Royal Statistical Society. Series A (General), 137(1):24, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345141.

Sichel:1974:DRS

[520] H. S. Sichel. On a distribution representing sentence-length in written prose. Journal of the Royal Statistical Society. Series A (General), 137(1):25– 34, ???? 1974. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345142.

Pickford:1974:SAU

[521] Michael Pickford. A statistical analysis of university administration expenditure. Journal of the Royal Statistical Society. Series A (General), 137(1):35– 47, ???? 1974. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345143.

Jones:1974:PSB

[522] C. L. Jones, G. M. Littlejohn, and A. F. McPherson. Predicting sciencebased study at university. Journal of the Royal Statistical Society. Series A (General), 137(1):48–59, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345144.

Gales:1974:TCT

[523] Kathleen E. Gales and P. H. Marks. Twentieth Century trends in the work of women in England and Wales. Journal of the Royal Statistical Society. Series A (General), 137(1):60–74, ????

1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345145.

Fowler:1974:ATI

[524] R. F. Fowler. An ambiguity in the terminology of index number construction. Journal of the Royal Statistical Society. Series A (General), 137(1):75– 88, ???? 1974. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345146.

Kendall:1974:BRB

[525] David Kendall. Book review: Letters on Probability, by Alfréd Rényi; László Vekerdi. Journal of the Royal Statistical Society. Series A (General), 137 (1):89–90, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345147.

Smith:1974:BRBa

[526] Cedric A. B. Smith. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability. Vol. 5: Darwinian, Neo-Darwinian and Non-Darwinian Evolution, by Lucien M. Le Cam; Jerzy Neyman; Elizabeth L. Scott. Journal of the Royal Statistical Society. Series A (General), 137(1):90–91, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345148.

Ashford:1974:BRBa

[527] J. R. Ashford. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability. Volume 6: Effects of Pollution on Health, by Lucien M. Le Cam; Jerzy Neyman; Elizabeth L. Scott. Journal of the Royal Statistical Society. Series A (General), 137(1):91–92, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345149.

Fielding:1974:BRBa

[528] A. Fielding. Book review: Statistical Inference Under Order Restrictions. The Theory and Application of Isotonic Regression, by R. E. Barlow; D. J. Bartholomew; J. M. Bremner; H. D. Brunk. Journal of the Royal Statistical Society. Series A (General), 137(1):92– 93, ???? 1974. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345150.

Sprent:1974:BRBa

[529] P. Sprent. Book review: Linear Statistical Inference and its Applications, by C. Radhakrishna Rao. Journal of the Royal Statistical Society. Series A (General), 137(1):93–94, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345151.

Ord:1974:BRB

[530] Keith Ord. Book review: Discrete Techniques of Parameter Estimation. The Equation Error Formulation, by Jerry M. Mendel. Journal of the Royal Statistical Society. Series A (General), 137(1):94–95, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345152.

Theobald:1974:BRB

[531] C. M. Theobald. Book review: Introduction to The Theory of Statistics, by Harold J. Larson. Journal of the Royal Statistical Society. Series A (General), 137(1):95–96, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345153.

Tong:1974:BRB

[532] H. Tong. Book review: Probability Distributions: An Introduction to Probability Theory with Applications, by Chris P. Tsokos. Journal of the Royal Statistical Society. Series A (General), 137(1):96–97, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345154.

Carrier:1974:BRB

[533] N. H. Carrier. Book review: The Methods and Materials of Demography, Vols. 1 and 2, by Henry S. Shryock; Jacob S. Siegel. Journal of the Royal Statistical Society. Series A (General), 137(1):97–99, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345155.

Herzberg:1974:BRB

[534] Agnes M. Herzberg. Book review: Bibliography on Experiment and Treatment Design Pre-1968, by Walter T. Federer; Leslie N. Balaam. Journal of the Royal Statistical Society. Series A (General), 137(1):99–100, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345156.

Renshaw:1974:BRB

[535] Eric Renshaw. Book review: Stochastic Differential Equations, by I. I. Gihman; A. V. Skorohod; K. Wickwire. Journal of the Royal Statistical Society. Series A (General), 137(1):100–101, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345157.

Rahman:1974:BRB

[536] N. A. Rahman. Book review: Statistics 1. Probability and Probability Distributions, by A. K. Shahani; P. K. Nandi. Journal of the Royal Statistical Society. Series A (General), 137 (1):101–102, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345158.

East:1974:BRB

[537] D. A. East. Book review: Numerical Mathematics. Exercises in Computing with a Desk Calculator, by H. Neill; A. J. Moakes. Journal of the Royal Statistical Society. Series A (General), 137 (1):102–103, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345159.

Prais:1974:BRBa

[538] Zmira Prais. Book review: Sources of Economic and Business Statistics, by Bernard Edwards. Journal of the Royal Statistical Society. Series A (General), 137(1):103, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345160.

Leser:1974:BRB

[539] C. E. V. Leser. Book review: Grundriss der Statistik. Teil 2: Daten. Ihre Gewinnung und Verarbeitung, by Gunter Menges; Heinz J. Skala. Journal of the Royal Statistical Society. Series A (General), 137(1):103–104, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345161.

Wolfe:1974:BRB

[540] J. N. Wolfe. Book review: Cost-Benefit Analysis in Education. A Case Study of Kenya, by Hans H. Thias; Martin Carnoy. Journal of the Royal Statistical Society. Series A (General), 137 (1):104–105, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345162.

Prais:1974:BRBb

[541] Zmira Prais. Book review: Economic Research: Retrospect and Prospect: The Business Cycles Today. Fiftieth Anniversary Colloquium. I, by Victor Zaranowitz. Journal of the Royal Statistical Society. Series A (General), 137 (1):105–106, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345163.

Sumner:1974:BRB

[542] M. T. Sumner. Book review: The Universities — National Bureau Conference on Secular Inflation, by National Bureau of Economic Research. Journal of the Royal Statistical Society. Series A (General), 137(1):107–108, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345164.

Thomas:1974:BRBa

[543] Roy Thomas. Book review: The Pricing System, by Robert Haney Scott. Journal of the Royal Statistical Society. Series A (General), 137(1):108, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345165.

Bowers:1974:BRBa

[544] J. K. Bowers. Book review: Effects of Selective Employment Tax. Final Report, by W. B. Reddaway; D. R. Glynn; J. D. Sugden; P. M. Croxford; C. H. Fletcher; J. S. O’Donnell. Journal of the Royal Statistical Society. Series A (General), 137(1):109–111, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345166.

Anonymous:1974:CNa

[545] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 137(1):112–114, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345167.

Anonymous:1974:RPa

[546] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 137(1):115–124, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345168.

Anonymous:1974:LAa

[547] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 137(1):125–126, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345169.

Anonymous:1974:RSS

[548] Anonymous. Royal Statistical Society regulations of the Research Section. Journal of the Royal Statistical Society. Series A (General), 137(1):127– 129, ???? 1974. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2345170.

Anonymous:1974:BMa

[549] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 137(1):i–ii, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345171.

Anonymous:1974:FMb

[550] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 137(2):??, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344545.

Newbold:1974:EFU

[551] P. Newbold and C. W. J. Granger. Experience with forecasting univariate time series and the combination of forecasts. Journal of the Royal Statistical Society. Series A (General), 137(2): 131–165, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344546.

Sukhatme:1974:PPS

[552] P. V. Sukhatme. The protein problem, its size and nature. Journal of the Royal Statistical Society. Series A (General), 137(2):166–199, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344547.

Cameron:1974:FMR

[553] M. H. Cameron and Judith E. Nash. On forecasting the manpower requirements of an organization with homogeneous workloads. Journal of the Royal Statistical Society. Series A (General), 137(2):200–218, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344548.

Jones:1974:LSE

[554] Susan Jones. The London School of Economics computer-based bibliography of statistical literature. Journal of the Royal Statistical Society. Series A (General), 137(2):219–226, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344549.

Hoerl:1974:RSE

[555] Arthur E. Hoerl and Herbert K. Fallin. Reliability of subjective evaluations in a high incentive situation. Journal of the Royal Statistical Society. Series A (General), 137(2):227–230, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344550.

Teather:1974:STD

[556] D. Teather. Statistical techniques for diagnosis. Journal of the Royal Statistical Society. Series A (General), 137 (2):231–244, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344551.

Turnbull:1974:SDT

[557] P. Turnbull and G. William. Sex differentials in teachers’ pay. Journal of the Royal Statistical Society. Series A (General), 137(2):245–258, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344552.

Kendall:1974:PL

[558] David Kendall. Paul Levy. Journal of the Royal Statistical Society. Series A (General), 137(2):259–260, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344553.

Wilkinson:1974:C

[559] G. N. Wilkinson. E. A. Cornish. Journal of the Royal Statistical Society. Series A (General), 137(2):260, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344554.

Pearce:1974:HB

[560] S. C. Pearce. T. N. Hoblyn, O.B.E. Journal of the Royal Statistical Society. Series A (General), 137(2):261, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344555.

Barnard:1974:BRB

[561] G. A. Barnard. Book review: An Objective Theory of Probability, by D. A. Gillies. Journal of the Royal Statistical Society. Series A (General), 137 (2):262–264, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344556.

Knott:1974:BRB

[562] M. Knott. Book review: Nonparametric Statistical Methods, by Myles Hollander; Douglas A. Wolfe. Journal of the Royal Statistical Society. Series A (General), 137(2):264, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344557.

Campbell:1974:BRB

[563] R. C. Campbell. Book review: Introduction to Biostatistics, by Robert R. Sokal; James F. Rohlf. Journal of the Royal Statistical Society. Series A (General), 137(2):265, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344558.

Sargan:1974:BRB

[564] J. D. Sargan. Book review: An Introduction to Bayesian Inference in Econometrics, by Arnold Zellner. Journal of the Royal Statistical Society. Series A (General), 137(2):265–266, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344559.

Conway:1974:BRB

[565] Freda Conway. Book review: Bayesian Statistics for Social Scientists, by Lawrence D. Phillips. Journal of the Royal Statistical Society. Series A (General), 137(2):266–267, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344560.

Wilson:1974:BRBa

[566] S. R. Wilson. Book review: Stochastic Processes and Applications to Biology and Medicine. (Parts 1 and 2), by M. Iosifescu; P. Tautu. Journal of the Royal Statistical Society. Series A (General), 137(2):267–268, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344561.

Gani:1974:BRB

[567] J. Gani. Book review: Mathematical Models for the Growth of Human Populations, by J. H. Pollard. Journal of the Royal Statistical Society. Series A (General), 137(2):268–269, ????

1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344562.

Vajda:1974:BRBa

[568] S. Vajda. Book review: Projective Planes, by D. R. Hughes; F. C. Piper. Journal of the Royal Statistical Society. Series A (General), 137(2):269, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344563.

Grimshaw:1974:BRB

[569] J. J. Grimshaw. Book review: Table of d′ and β, by P. R. Freeman. Journal of the Royal Statistical Society. Series A (General), 137(2):269–270, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344564.

Fielding:1974:BRBb

[570] A. Fielding. Book review: Statistical Methods for Rates and Proportions, by Joseph L. Fleiss. Journal of the Royal Statistical Society. Series A (General), 137(2):270–271, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344565.

Wilkinson:1974:BRB

[571] R. K. Wilkinson. Book review: Applied Microeconomics. Problems in Estimation, Forecasting and Decision-Making, by Richard Schmalensee. Journal of the Royal Statistical Society. Series A (General), 137(2):271–272, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344566.

Wild:1974:BRB

[572] J. V. Wild. Book review: Statistical Mathematics. A Companion to “A Second Course in Statistics”, by Robert Loveday. Journal of the Royal Statistical Society. Series A (General), 137 (2):272–273, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344567.

Barton:1974:BRBa

[573] D. E. Barton. Book review: An Introduction to Quantitative Methods for Historians, by Roderick Floud. Journal of the Royal Statistical Society. Series A (General), 137(2):273–274, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344568.

Nicholson:1974:BRB

[574] M. Nicholson. Book review: The Wages of War, 1816–1965. A Statistical Handbook, by J. David Singer; Melvin Small. Journal of the Royal Statistical Society. Series A (General), 137 (2):274–275, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344569.

White:1974:BRB

[575] D. J. White. Book review: Dynamic Programming with Management Applications, by N. A. J. Hastings. Journal of the Royal Statistical Society. Series A (General), 137(2):275–276, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344570.

Sprent:1974:BRBb

[576] P. Sprent. Book review: Structural Equation Models in the Social Sciences, by Arthur S. Goldberger; Otis Dudley Duncan. Journal of the Royal Statistical Society. Series A (General), 137 (2):276–277, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344571.

Lewes:1974:BRB

[577] F. M. M. Lewes. Book review: The General Household Survey. Introductory Report. An Inter-Departmental Survey Sponsored, by The Central Statistical Office. by Office of Population Censuses and Surveys. Journal of the Royal Statistical Society. Series A (General), 137(2):277–278, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344572.

Granger:1974:BRB

[578] C. W. J. Granger. Book review: Theory of Demand, Real and Monetary, by M. Morishima. Journal of the Royal Statistical Society. Series A (General), 137(2):278–279, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344573.

Pole:1974:BRB

[579] J. D. Pole. Book review: The Demand for Health: A Theoretical and Empirical Investigation, by Michael Grossman. Journal of the Royal Statistical Society. Series A (General), 137(2): 279–280, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344574.

Noble:1974:BRB

[580] T. Noble. Book review: The Use of Official Statistics in Sociology. A Critique of Positivism and Ethnomethodology, by Barry Hindess. Journal of the Royal Statistical Society. Series A (General), 137(2):280, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344575.

Anonymous:1974:RPb

[581] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 137(2):281–293, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344576.

Anonymous:1974:LAb

[582] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 137(2):294–295, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344577.

Anonymous:1974:BMb

[583] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 137(2):i–iv, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344578.

Anonymous:1974:FMc

[584] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 137(3):??, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344951.

Atkinson:1974:MPF

[585] A. B. Atkinson and T. W. Meade. Methods and preliminary findings in assessing the economic and health services consequences of smoking, with particular reference to lung cancer. Journal of the Royal Statistical Society. Series A (General), 137(3):297– 312, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344952.

Russell:1974:CSF

[586] M. A. H. Russell, J. Peto, and U. A. Patel. The classification of smoking by factorial structure of motives. Journal of the Royal Statistical Society. Series A (General), 137(3):313–346, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344953.

Ashford:1974:DDP

[587] J. R. Ashford and R. G. Hunt. The distribution of doctor–patient contacts in the National Health Service. Journal of the Royal Statistical Society. Series A (General), 137(3):347–383, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344954.

Miller:1974:VRP

[588] W. L. Miller, Gillian Raab, and K. Britto. Voting research and the population census 1918–71: Surrogate data for constituency analyses. Journal of the Royal Statistical Society. Series A (General), 137(3):384–411, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344955.

Anonymous:1974:RJC

[589] Anonymous. Report of the Joint Committee on the Teaching of Statistics in Colleges Providing a Technical or Commercial Education: Wednesday 23 January 1974. Journal of the Royal Statistical Society. Series A (General), 137 (3):412–427, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344956. See

corrections [726, 766].

Grove:1974:BRB

[590] D. M. Grove. Book review: Comparative Statistical Inference, by Vic Barnett. Journal of the Royal Statistical Society. Series A (General), 137(3): 428–429, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344957.

Fisk:1974:BRBa

[591] P. R. Fisk. Book review: Time Series, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 137(3):430, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344958.

Vajda:1974:BRBb

[592] S. Vajda. Book review: Stochastic Models for Social Processes, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series A (General), 137(3):431, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344959.

Barrett:1974:BRB

[593] J. C. Barrett. Book review: A Stochastic Model of Human Reproduction, by Prithwis Das Gupta. Journal of the Royal Statistical Society. Series A (General), 137(3):431–432, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344960.

Ashford:1974:BRBb

[594] J. R. Ashford. Book review: Statistical Principles in Health Care Information, by James S. Kilpatrick. Journal of the Royal Statistical Society. Series A (General), 137(3):432, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344961.

Williams:1974:BRBa

[595] E. J. Williams. Book review: Conditional Inference and Models for Measuring, by Erling B. Anderson. Journal of the Royal Statistical Society. Series A (General), 137(3):433, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344962.

Silvey:1974:BRB

[596] S. D. Silvey. Book review: Distribution Theory for Tests Based on the Sample Distribution Function, by J. Durbin. Journal of the Royal Statistical Society. Series A (General), 137(3):433– 434, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344963.

Haigh:1974:BRB

[597] J. Haigh. Book review: Asymptotic Theory of Rank Tests for Independence, by F. H. Ruymgaart. Journal of the Royal Statistical Society. Series A (General), 137(3):434, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344964.

Brook:1974:BRB

[598] D. Brook. Book review: Concepts of Statistical Inference, by William C.

Guenther. Journal of the Royal Statistical Society. Series A (General), 137 (3):435, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344965.

Loynes:1974:BRB

[599] R. M. Loynes. Book review: Stochastic Analysis. A Tribute to the Memory of Rollo Davidson, by D. G. Kendall; E. F. Harding. Journal of the Royal Statistical Society. Series A (General), 137(3):435–436, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344966.

Barton:1974:BRBb

[600] D. E. Barton. Book review: A Survey of Combinatorial Theory, by J. N. Srivastava; F. Harary; C. R. Rao; G. C. Rota; S. S. Shrikhande. Journal of the Royal Statistical Society. Series A (General), 137(3):436, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344967.

Bartholomew:1974:BRB

[601] D. J. Bartholomew. Book review: Understanding Statistical Reasoning, by Eleanor W. Willemsen. Journal of the Royal Statistical Society. Series A (General), 137(3):436–437, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344968.

Chatfield:1974:BRB

[602] C. Chatfield. Book review: Significance Tests, by Evelyn Caulcott. Journal of the Royal Statistical Society. Series A (General), 137(3):437, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344969.

Maxwell:1974:BRB

[603] A. E. Maxwell. Book review: Fundamental Statistics in Psychology and Education, by J. P. Guilford; Benjamin Fruchter. Journal of the Royal Statistical Society. Series A (General), 137 (3):437–438, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344970.

Holt:1974:BRB

[604] D. Holt. Book review: Statistics and Society. Data Collection and Interpretation, by Walter T. Federer. Journal of the Royal Statistical Society. Series A (General), 137(3):438–439, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344971.

Colquhoun:1974:BRB

[605] D. Colquhoun. Book review: Biostatistics in Pharmacology. Volumes I and II, by A. L. Delaunois. Journal of the Royal Statistical Society. Series A (General), 137(3):439–440, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344972.

Reed:1974:BRB

[606] R. J. Reed. Book review: Approximate Stochastic Behavior of n-Server Systems with Large n, by G. F. Newell. Journal of the Royal Statistical Society. Series A (General), 137(3):440, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344973.

Freeman:1974:BRB

[607] P. R. Freeman. Book review: Uncertainty, Calibration and Probability. The Statistics of Scientific and Industrial Measurement, by C. F. Dietrich. Journal of the Royal Statistical Society. Series A (General), 137(3):441–442, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344974.

Lindley:1974:BRB

[608] D. V. Lindley. Book review: Historical Aspects of the Bayesian Controversy, with Comprehensive Bibliography, by Jean Draper Weber. Journal of the Royal Statistical Society. Series A (General), 137(3):442, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344975.

Land:1974:BRB

[609] F. F. Land. Book review: Introduction to Computing, by Heinz Dinter. Journal of the Royal Statistical Society. Series A (General), 137(3):442, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344976.

Clarke:1974:BRB

[610] M. R. B. Clarke. Book review: Pattern Classification and Scene Analysis, by Richard O. Duda; Peter E. Hart. Journal of the Royal Statistical Society. Series A (General), 137(3):442–443, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344977.

Shearn:1974:BRB

[611] D. C. S. Shearn. Book review: Industrial Systems: Planning, Analysis, Control, by David D. Bedworth. Journal of the Royal Statistical Society. Series A (General), 137(3):443–444, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344978.

Moore:1974:BRBa

[612] P. G. Moore. Book review: Decision, Strategies and New Ventures. Modern Tools for Top Management, by W. G. Byrnes; B. K. Chesterton. Journal of the Royal Statistical Society. Series A (General), 137(3):444–445, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344979.

Easterfield:1974:BRB

[613] T. E. Easterfield. Book review: O.R. in World War 2. Operational Research against the U-Boat, by C. H. Waddington. Journal of the Royal Statistical Society. Series A (General), 137(3):445– 446, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344980.

Thomas:1974:BRBb

[614] Brinley Thomas. Book review: Economic Dynamics, by Roy Harrod. Journal of the Royal Statistical Society. Series A (General), 137(3):446– 447, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344981.

Drettakis:1974:BRB

[615] E. G. Drettakis. Book review: Migration and Economic Growth. A Study of Great Britain and the Atlantic Economy, by Brinley Thomas. Journal of the Royal Statistical Society. Series A (General), 137(3):447–448, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344982.

Burkitt:1974:BRB

[616] B. Burkitt. Book review: Growth and Distribution, by W. A. Eltis. Journal of the Royal Statistical Society. Series A (General), 137(3):448–449, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344983.

Fisk:1974:BRBb

[617] P. R. Fisk. Book review: Theory of Econometrics, by A. Koutsoyiannis. Journal of the Royal Statistical Society. Series A (General), 137(3):449, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344984.

Champernowne:1974:BRB

[618] D. G. Champernowne. Book review: On Economic Inequality, by Armartya Sen. Journal of the Royal Statistical Society. Series A (General), 137(3): 449–450, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344985.

Ford:1974:BRB

[619] J. L. Ford. Book review: Studies in the Pure Theory of International Trade, by Raveendra N. Batra. Journal of the Royal Statistical Society. Series A (General), 137(3):450–452, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344986.

Anonymous:1974:CNb

[620] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 137(3):453–455, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344987.

Cochran:1974:GWS

[621] W. G. Cochran. George W. Snedecor. Journal of the Royal Statistical Society. Series A (General), 137(3):456– 457, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344988.

Robinson:1974:COG

[622] H. W. Robinson. C. Oswald George. Journal of the Royal Statistical Society. Series A (General), 137(3):457– 458, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344989.

Anonymous:1974:RPc

[623] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 137(3):459–479, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344990.

Anonymous:1974:LAc

[624] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 137(3):480–481, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344991.

Anonymous:1974:BMc

[625] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series A (General), 137(3):??, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344992.

Anonymous:1974:VIb

[626] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 137(4):654– 657, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344710.

Anonymous:1974:FMd

[627] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 137(4):??, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344711.

Hutton:1974:PSS

[628] D. S. S. Hutton, Valerie Imber, and H. D. Mitchell. Personal social services statistics. Journal of the Royal Statistical Society. Series A (General), 137 (4):483–531, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344712.

Brass:1974:PPP

[629] W. Brass. Perspectives in population prediction: Illustrated by the statistics of England and Wales. Journal of the Royal Statistical Society. Series A (General), 137(4):532–583, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344713.

Young:1974:NLM

[630] Andrew Young and P-C. G. Vassiliou. A non-linear model on the promotion of staff. Journal of the Royal Statistical Society. Series A (General), 137 (4):584–595, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344714.

Anonymous:1974:POH

[631] Anonymous. Proceedings of the one hundred and fortieth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 137(4):596, ???? 1974. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344715.

Anonymous:1974:ARC

[632] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 137 (4):597–618, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344716.

Barnett:1974:BRB

[633] V. Barnett. Book review: Scientific Inference, by Harold Jeffreys. Journal of the Royal Statistical Society. Series A (General), 137(4):619–620, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344717.

Downton:1974:BRB

[634] F. Downton. Book review: Statistics: A Guide to the Unknown, by Judith M. Tanur. Journal of the Royal Statistical Society. Series A (General), 137 (4):620–621, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344718.

Coleman:1974:BRB

[635] Rodney Coleman. Book review: Stochastic Geometry: A Tribute to the Memory of Rollo Davidson, by E. F. Harding; D. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 137(4):621–622, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344719.

Pearce:1974:BRB

[636] S. C. Pearce. Book review: Statistics for Biologists, by R. C. Campbell. Journal of the Royal Statistical Society. Series A (General), 137(4):622– 623, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344720.

Pridmore:1974:BRB

[637] W. A. Pridmore. Book review: Statistics in Small Doses, by Winifred M. Castle. Journal of the Royal Statistical Society. Series A (General), 137 (4):623–624, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344721.

Oldham:1974:BRB

[638] P. D. Oldham. Book review: An Introduction to Medical Statistics, by H. O. Lancaster. Journal of the Royal Statistical Society. Series A (General), 137 (4):624–625, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344722.

Moore:1974:BRBb

[639] P. G. Moore. Book review: Basic Statistics with Business Applications, by R. C. Clelland; J. S. DeCani; F.

E. Brown. Journal of the Royal Statistical Society. Series A (General), 137 (4):625–626, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344723.

Kmietowicz:1974:BRB

[640] Z. W. Kmietowicz. Book review: A Manual of Sampling Techniques, by R. K. Som. Journal of the Royal Statistical Society. Series A (General), 137 (4):626–627, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344724.

Smith:1974:BRBb

[641] T. M. F. Smith. Book review: The Design of Sample Surveys, by Des Raj. Journal of the Royal Statistical Society. Series A (General), 137(4):627, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344725.

Harris:1974:BRB

[642] Amelia I. Harris. Book review: Questionnaire Design Manual, by Social & Community Planning Research. Journal of the Royal Statistical Society. Series A (General), 137(4):627–628, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344726.

Rowntree:1974:BRB

[643] J. A. Rowntree. Book review: Patterns in Sickness Absence: Experience of London Transport Staff Over Two Decades, by J. E. Ager; P. A. B. Raffle. Journal of the Royal Statistical Society. Series A (General), 137(4):628– 629, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344727.

Bowers:1974:BRBb

[644] J. K. Bowers. Book review: Statistical Methods for the Analysis of Earnings Data with Special Application to Salaries in Swedish Industry, by A. Klevmarken. Journal of the Royal Statistical Society. Series A (General), 137 (4):629, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344728.

Shahani:1974:BRB

[645] A. K. Shahani. Book review: Statistical Quality Control, by Eugene L. Grant; Richard S. Leavenworth. Journal of the Royal Statistical Society. Series A (General), 137(4):629–630, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344729.

Barton:1974:BRBc

[646] D. E. Barton. Book review: The Dimensions of Quantitative Research in History, by W. O. Aydelotte; A. G. Bogue; R. W. Fogel. Journal of the Royal Statistical Society. Series A (General), 137(4):630–631, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344730.

Thomas:1974:BRBc

[647] C. G. Thomas. Book review: The Bureau of the Census, by Ross Ecklar. Journal of the Royal Statistical Society. Series A (General), 137(4):632, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344731.

Wilson:1974:BRBb

[648] T. D. Wilson. Book review: The National Archives and Statistical Research, by M. H. Fishbein. Journal of the Royal Statistical Society. Series A (General), 137(4):632–633, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344732.

Lloyd:1974:BRB

[649] Emlyn Lloyd. Book review: Inventory Control and Water Storage, by A. Prekopa. Journal of the Royal Statistical Society. Series A (General), 137 (4):633–634, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344733.

Bessell:1974:BRB

[650] J. E. Bessell. Book review: Population and Land in World Agriculture: Recent Trends and Relationships, by Joginder Kumar. Journal of the Royal Statistical Society. Series A (General), 137 (4):634–635, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344734.

Wilson:1974:BRBc

[651] A. G. Wilson. Book review: World Urbanization 1950–1970. Volume II: Analysis of Trends, Relationships and Developments, by Kingsley Davis. Journal of the Royal Statistical Society. Series A (General), 137(4):636, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344735.

Ehrenberg:1974:BRB

[652] A. S. C. Ehrenberg. Book review: New Commodities and Consumer Behaviour, by D. S. Ironmonger. Journal of the Royal Statistical Society. Series A (General), 137(4):636–637, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344736.

Binmore:1974:BRB

[653] K. G. Binmore. Book review: The Computation of Economic Equilibria, by Herbert Scarf; Terje Hansen. Journal of the Royal Statistical Society. Series A (General), 137(4):637–638, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344737.

Godfrey:1974:BRB

[654] Michael D. Godfrey. Book review: The Computer from Pascal to von Neumann, by H. H. Goldstine. Journal of the Royal Statistical Society. Series A (General), 137(4):638–639, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344738.

Hansford-Miller:1974:BRB

[655] Frank Hansford-Miller. Book review: The School Mathematics Project: The First Ten Years, Being Primarily the Collection of the Director’s Reports for the Years 1961–71, by Bryan Thwaites. Journal of the Royal Statistical Society. Series A (General), 137(4):639– 640, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344739.

Williams:1974:BRBb

[656] H. P. Williams. Book review: FOR-TRAN Programs for Mathematical Programming, by A. Land; S. Powell.

Journal of the Royal Statistical Society. Series A (General), 137(4):640– 641, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344740.

McLone:1974:BRB

[657] R. R. McLone. Book review: Applicable Mathematics — A Course for Scientists and Engineers, by R. G. Goult; R. F. Hoskins; J. A. Milner; M. J. Pratt. Journal of the Royal Statistical Society. Series A (General), 137(4): 641–642, ???? 1974. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344741.

Woodroofe:1974:LHO

[658] Ernest Woodroofe. Lord Heyworth of Oxton. Journal of the Royal Statistical Society. Series A (General), 137 (4):643–644, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344742.

Jones:1974:DCJ

[659] R. Huws Jones. David Caradog Jones, 1883–1974. Journal of the Royal Statistical Society. Series A (General), 137 (4):644–645, ???? 1974. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344743.

Anonymous:1974:RPd

[660] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 137(4):646–651, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344744.

Anonymous:1974:LAd

[661] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 137(4):652–653, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344745.

Anonymous:1974:BMd

[662] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 137(4):??, ???? 1974. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344746.

Anonymous:1975:VIa

[663] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 138(1):??, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345244.

Anonymous:1975:FMa

[664] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 138(1):i–ii, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345245.

Daniels:1975:SUP

[665] H. E. Daniels. Statistics in universities — a personal view. Journal of the Royal Statistical Society. Series A (General), 138(1):1–17, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345246.

Irwin:1975:GWDa

[666] J. O. Irwin. The generalized Waring distribution. Part I. Journal of the Royal Statistical Society. Series A (General), 138(1):18–31, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345247.

Hart:1975:EEB

[667] R. A. Hart and D. I. MacKay. Engineering earnings in Britain, 1914–68. Journal of the Royal Statistical Society. Series A (General), 138(1):32–50, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345248.

Greenfield:1975:EMC

[668] C. C. Greenfield. On the estimation of a missing cell in a 2 × 2 contingency table. Journal of the Royal Statistical Society. Series A (General), 138(1):51– 61, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345249.

Marshall:1975:EAD

[669] Malcomn L. Marshall. Equilibrium age distributions for graded systems. Journal of the Royal Statistical Society. Series A (General), 138(1):62–69, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345250.

Grinyer:1975:PEM

[670] P. H. Grinyer and D. Norburn. Planning for existing markets: Perceptions of executives and financial performance. Journal of the Royal Statistical Society. Series A (General), 138(1):70– 97, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345251.

Kingman:1975:BRBa

[671] J. F. C. Kingman. Book review: Theory of Probability, A Critical Introductory Treatment, by Bruno de Finetti; Antonio Machi; Adrian Smith. Journal of the Royal Statistical Society. Series A (General), 138(1):98–99, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345252.

Nelder:1975:BRB

[672] J. A. Nelder. Book review: The Analysis of Categorical Data, by R. L. Plackett. Journal of the Royal Statistical Society. Series A (General), 138(1):99– 100, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345253.

Holder:1975:BRB

[673] R. L. Holder. Book review: Multivariate Statistical Inference, by D. G. Kabe; R. P. Gupta. Journal of the Royal Statistical Society. Series A (General), 138(1):100–101, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345254.

Gray:1975:BRB

[674] J. R. Gray. Book review: Elementary Principles of Probability and Information, by R. F. Wrighton. Journal of the Royal Statistical Society. Series A (General), 138(1):101, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345255.

Lutz:1975:BRB

[675] W. Lutz. Book review: Statistische Methoden, by L. Sachs. Journal of the Royal Statistical Society. Series A (General), 138(1):102, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345256.

Kemp:1975:BRBa

[676] C. D. Kemp. Book review: Biometrical Interpretation, by Neil Gilbert. Journal of the Royal Statistical Society. Series A (General), 138(1):102–103, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345257.

Cormack:1975:BRBa

[677] R. M. Cormack. Book review: Spatial Autocorrelation, by A. D. Cliff; J. K. Ord. Journal of the Royal Statistical Society. Series A (General), 138 (1):103–104, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345258.

Burrows:1975:BRB

[678] Paul Burrows. Book review: Statistical and Mathematical Aspects of Pollution Problems, by John W. Pratt. Journal of the Royal Statistical Society. Series A (General), 138(1):104, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345259.

Vajda:1975:BRB

[679] S. Vajda. Book review: Optimisation Methods, by R. W. Cottle; J. Krarup. Journal of the Royal Statistical Society. Series A (General), 138(1):105, ????

1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345260.

Bartholomew:1975:BRBa

[680] D. J. Bartholomew. Book review: Statistics for Social Scientists, a Co-Ordinated Learning System, by F. J. Kohout. Journal of the Royal Statistical Society. Series A (General), 138 (1):105–106, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345261.

Ray:1975:BRBa

[681] W. D. Ray. Book review: Statistical Techniques in Business and Economics, by Robert D. Mason. Journal of the Royal Statistical Society. Series A (General), 138(1):106–107, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345262.

Morgan:1975:BRBa

[682] R. W. Morgan. Book review: Mathematics for Management Science, by Judith Atherton; Michael J. C. Martin; Geraint Roberts. Journal of the Royal Statistical Society. Series A (General), 138(1):107, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345263.

Cox:1975:BRBa

[683] Joan G. Cox. Book review: Manpower and Management Science, by D. J. Bartholomew; A. R. Smith. Journal of the Royal Statistical Society. Series A (General), 138(1):108–109, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345264.

Cox:1975:BRBb

[684] P. R. Cox. Book review: A Workbook in Demography, by Roland Pressat. Journal of the Royal Statistical Society. Series A (General), 138(1):109, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345265.

Murphy:1975:BRB

[685] M. Murphy. Book review: Sources of Social Statistics, by Bernard Edwards. Journal of the Royal Statistical Society. Series A (General), 138(1):109– 111, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345266.

Atkinson:1975:BRB

[686] A. B. Atkinson. Book review: The Distribution of Income between Persons, by D. G. Champernowne. Journal of the Royal Statistical Society. Series A (General), 138(1):111–112, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345267.

Hutton:1975:BRB

[687] J. P. Hutton. Book review: Outline of International Trade: Commodity Flows and Division of Production between Countries, by L. Sirc. Journal of the Royal Statistical Society. Series A (General), 138(1):112–113, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345268.

Clayton:1975:BRB

[688] G. Clayton. Book review: The Banks and the Monetary System in the U.K.

1959–1971, by J. E. Wadsworth. Journal of the Royal Statistical Society. Series A (General), 138(1):113–114, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345269.

Rahman:1975:BRB

[689] N. A. Rahman. Book review: How to Solve Problems: Elements of a Theory of Problems and Problem Solving, by Wayne Wickelgren. Journal of the Royal Statistical Society. Series A (General), 138(1):114–115, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345270.

Whittle:1975:BRB

[690] P. Whittle. Book review: Statistical Mechanics and Field Theory, by R. N. Sen; C. Weil. Journal of the Royal Statistical Society. Series A (General), 138 (1):115–116, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345271.

Anonymous:1975:RPa

[691] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 138(1):117–125, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345272.

Anonymous:1975:LAa

[692] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 138(1):126–127, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345273.

Anonymous:1975:BMa

[693] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 138(1):??, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345274.

Anonymous:1975:FMb

[694] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 138(2):??, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984644.

Moore:1975:TES

[695] Peter G. Moore. A tribute to Egon Sharpe Pearson. Journal of the Royal Statistical Society. Series A (General), 138(2):129–130, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2984645.

Curnow:1975:MMF

[696] R. N. Curnow and Charles Smith. Multifactorial models for familial diseases in man. Journal of the Royal Statistical Society. Series A (General), 138 (2):131–169, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984646.

Ward:1975:WME

[697] D. H. Ward. Weights and measures and the EEC. Journal of the Royal Statistical Society. Series A (General), 138 (2):170–203, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984647.

Irwin:1975:GWDb

[698] J. O. Irwin. The generalized Waring distribution. Part II. Journal of the Royal Statistical Society. Series A (General), 138(2):204–227, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984648.

Downton:1975:CSB

[699] F. Downton and Carmen Lockwood. Computer studies of baccarat, I: Chemin-De-Fer. Journal of the Royal Statistical Society. Series A (General), 138(2):228–238, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2984649.

Ehrenberg:1975:CHW

[700] A. S. C. Ehrenberg. Children’s heights and weights in 1905. Journal of the Royal Statistical Society. Series A (General), 138(2):239–241, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984650.

McDonald:1975:ASR

[701] John McDonald. An analysis of the significance of revisions to some quarterly U.K. national income time series. Journal of the Royal Statistical Society. Series A (General), 138(2):242–256, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984651.

Plackett:1975:BRBa

[702] R. L. Plackett. Book review: Introduction to Probability and Statistics. Part I: Probability, by Narayan C. Giri. Journal of the Royal Statistical Society. Series A (General), 138(2):257, ????

1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984652.

Barnett:1975:BRB

[703] V. Barnett. Book review: Applied Linear Statistical Models, by John Neter; William Wasserman. Journal of the Royal Statistical Society. Series A (General), 138(2):258, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984653.

Galbraith:1975:BRB

[704] Jane Galbraith. Book review: Applied Statistics: Analysis of Variance and Regression, by Olive Jean Dunn; Virginia A. Clark. Journal of the Royal Statistical Society. Series A (General), 138(2):259, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984654.

Seheult:1975:BRB

[705] A. H. Seheult. Book review: Introduction to Probability Theory and Statistical Inference, by Harold J. Larson. Journal of the Royal Statistical Society. Series A (General), 138(2):259– 260, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2984655.

Ray:1975:BRBb

[706] W. D. Ray. Book review: Multivariate Statistical Methods for Business and Economics, by Ben W. Bolch; Cliff J. Huang. Journal of the Royal Statistical Society. Series A (General), 138 (2):260–261, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984656.

Moore:1975:BRBa

[707] P. G. Moore. Book review: Statistical Analysis and Modelling for Management Decision-Making, by R. E. Willis; N. L. Chervany. Journal of the Royal Statistical Society. Series A (General), 138(2):261, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984657.

Rosenbaum:1975:BRB

[708] S. Rosenbaum. Book review: Statistics — An Interdisciplinary Approach, by Open University Course Team. Journal of the Royal Statistical Society. Series A (General), 138(2):261–263, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2984658.

Benjamin:1975:BRBa

[709] Bernard Benjamin. Book review: Significance Testing; Applied Non-Parametric Statistics; Quality Control, by Open University Course Team. Journal of the Royal Statistical Society. Series A (General), 138(2):263– 264, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2984659.

Benjamin:1975:BRBb

[710] Bernard Benjamin. Book review: Vital Statistics, by John Boreham; Cliff Holmes. Journal of the Royal Statistical Society. Series A (General), 138 (2):264–265, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984660.

Benjamin:1975:BRBc

[711] Bernard Benjamin. Book review: Basic Statistical Techniques, by K. D. C.

Stoodley. Journal of the Royal Statistical Society. Series A (General), 138 (2):265, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2984661.

Chatfield:1975:BRB

[712] C. Chatfield. Book review: Essentials of Statistics in Marketing, by C. S. Greensted; A. K. S. Jardine; J. D. MacFarlane. Journal of the Royal Statistical Society. Series A (General), 138 (2):265–266, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984662.

Cormack:1975:BRBb

[713] R. M. Cormack. Book review: Cluster Analysis, by Brian Everitt. Journal of the Royal Statistical Society. Series A (General), 138(2):266–267, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984663.

Healy:1975:BRB

[714] M. J. R. Healy. Book review: The Interpretation of Multiple Observations, by F. H. C. Marriott. Journal of the Royal Statistical Society. Series A (General), 138(2):267, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984664.

Kingman:1975:BRBb

[715] J. F. C. Kingman. Book review: Probability Theory: A Historical Sketch, by L. E. Maistrov; Samuel Kotz. Journal of the Royal Statistical Society. Series A (General), 138(2):267–268, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984665.

Rabinowicz:1975:BRB

[716] D. Rabinowicz and M. Stone. Book review: Inference and Decision, by D. A. S. Fraser. Journal of the Royal Statistical Society. Series A (General), 138 (2):268, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2984666.

Hill:1975:BRB

[717] Austin Bradford Hill. Book review: The Challenge of Facts. Selected Public Health Papers of Edgar Sydenstricker, by Edgar Sydenstricker; Richard V. Kasius. Journal of the Royal Statistical Society. Series A (General), 138 (2):268–269, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984667.

Holgate:1975:BRB

[718] P. Holgate. Book review: The Genetic Structure of Populations, by A. Jacquard. Journal of the Royal Statistical Society. Series A (General), 138 (2):269–271, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984668.

Hammersley:1975:BRBa

[719] J. M. Hammersley. Book review: Graphs, Groups and Surfaces, by Arthur T. White. Journal of the Royal Statistical Society. Series A (General), 138(2):271, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984669.

Gower:1975:BRB

[720] J. C. Gower. Book review: Non-Negative Matrices, by E. Seneta. Journal of the Royal Statistical Society. Series A (General), 138(2):271–272, ????

1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984670.

Whalley:1975:BRB

[721] J. Whalley. Book review: Allocation Under Uncertainty: Equilibrium and Optimality, by Jacques H. Dreze. Journal of the Royal Statistical Society. Series A (General), 138(2):272, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984671.

Hill:1975:DPS

[722] Austin Bradford Hill. Dr Percy Stocks, 1889–1974. Journal of the Royal Statistical Society. Series A (General), 138 (2):273–274, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984672.

Anonymous:1975:RPb

[723] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 138(2):275–291, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984673.

Anonymous:1975:LAb

[724] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 138(2):292–295, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984674.

Lee:1975:CNL

[725] Ronald Lee. Comment on “A New Look at Some Old Data: The Beveridge Wheat Price Series” by Granger and Hughes. Journal of the Royal Statistical Society. Series A (General), 138(2):296, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984675. See

[179].

Anonymous:1975:CRJ

[726] Anonymous. Correction: Report of the Joint Committee on the Teaching of Statistics in Colleges Providing a Technical or Commercial Education: Wednesday 23 January 1974. Journal of the Royal Statistical Society. Series A (General), 138(2):296, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984676. See [589].

Anonymous:1975:BMb

[727] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 138(2):i–iv, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984677.

Anonymous:1975:FMc

[728] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 138(3):??, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344579.

Cole:1975:LPR

[729] T. J. Cole. Linear and proportional regression models in the prediction of ventilatory function. Journal of the Royal Statistical Society. Series A (General), 138(3):297–338, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344580.

Bray:1975:OCN

[730] Jeremy Bray. Optimal control of a noisy economy with the UK as an example. Journal of the Royal Statistical Society. Series A (General), 138 (3):339–373, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344581.

Irwin:1975:GWDc

[731] J. O. Irwin. The generalized Waring distribution. Part III. Journal of the Royal Statistical Society. Series A (General), 138(3):374–384, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344582.

Durbin:1975:SAB

[732] J. Durbin and M. J. Murphy. Seasonal adjustment based on a mixed additive–multiplicative model. Journal of the Royal Statistical Society. Series A (General), 138(3):385–410, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344583.

Johnson:1975:ARB

[733] D. G. Johnson. The accuracy of regression based cost indices. Journal of the Royal Statistical Society. Series A (General), 138(3):411–422, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344584.

Hart:1975:MDE

[734] P. E. Hart. Moment distributions in economics: An exposition. Journal of the Royal Statistical Society. Series A (General), 138(3):423–434, ????

1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344585.

Kingman:1975:BRBc

[735] J. F. C. Kingman. Book review: Random Processes, by Murray Rosenblatt. Journal of the Royal Statistical Society. Series A (General), 138(3):435, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344586.

Plackett:1975:BRBb

[736] R. L. Plackett. Book review: The Analysis of Frequency Data, by Shelby J. Haberman. Journal of the Royal Statistical Society. Series A (General), 138 (3):435–436, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344587.

Coleman:1975:BRB

[737] Rodney Coleman. Book review: Fundamentals of Queueing Theory, by Donald Gross; Carl M. Harris. Journal of the Royal Statistical Society. Series A (General), 138(3):436–437, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344588.

Hammersley:1975:BRBb

[738] J. M. Hammersley. Book review: Quelques Problèmes Mathématiques en Physique Statistique, by M. Kac. Journal of the Royal Statistical Society. Series A (General), 138(3):437–438, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344589.

Collings:1975:BRB

[739] Stanley Collings. Book review: Basic Statistics: A Modern Approach, by Morris Hamburg. Journal of the Royal Statistical Society. Series A (General), 138(3):438–439, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344590.

Irving:1975:BRB

[740] Doreen Irving. Book review: Biostatistics: A Foundation for Analysis in the Health Sciences, by Wayne W. Daniel. Journal of the Royal Statistical Society. Series A (General), 138(3):439, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344591.

Kemp:1975:BRBb

[741] D. J. Kemp. Book review: Statistics for Management and Economics, by W. Mendenhall; J. E. Reinmuth. Journal of the Royal Statistical Society. Series A (General), 138(3):440, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344592.

Holt:1975:BRB

[742] D. Holt. Book review: Statistics: A Tool for the Social Sciences, by William Mendenhall; Lyman Ott; R. F. Larson. Journal of the Royal Statistical Society. Series A (General), 138(3):440– 441, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344593.

Bartholomew:1975:BRBb

[743] D. J. Bartholomew. Book review: Commonsense Statistics for Economists and Others, by F. R. Jolliffe. Journal of the Royal Statistical Society. Series A (General), 138(3): 441–442, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344594.

Burns:1975:BRB

[744] M. E. Burns. Book review: Statistics in Economics, by John D. Hey. Journal of the Royal Statistical Society. Series A (General), 138(3):442, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344595.

Lury:1975:BRB

[745] D. A. Lury. Book review: Living Statistics: An Introductory Text for West African Students, by J. Bibby. Journal of the Royal Statistical Society. Series A (General), 138(3):443, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344596.

Ord:1975:BRBa

[746] Keith Ord. Book review: Statistical Analysis of Spatial Dispersion, by A. Rogers. Journal of the Royal Statistical Society. Series A (General), 138 (3):443–444, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344597.

Cliff:1975:BRB

[747] A. D. Cliff. Book review: Organisation, Location and Behaviour: Decision-Making in Economic Geography, by P. Toyne. Journal of the Royal Statistical Society. Series A (General), 138(3): 444, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344598.

Hopes:1975:BRB

[748] R. Hopes. Book review: Manpower Planning Models, by D. J. Clough; C. G. Lewis; A. L. Oliver. Journal of the Royal Statistical Society. Series A (General), 138(3):444–446, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344599.

Garside:1975:BRB

[749] M. J. Garside. Book review: Statistical Techniques in Simulation, by Jack P. C. Kleijnen. Journal of the Royal Statistical Society. Series A (General), 138(3):446, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344600.

Huitson:1975:BRB

[750] A. Huitson. Book review: Statistical Distributions: A Handbook for Students and Practitioners, by N. A. J. Hastings; J. B. Peacock. Journal of the Royal Statistical Society. Series A (General), 138(3):447, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344601.

Keen:1975:BRB

[751] Joan Keen. Book review: Tables for Statisticians, by John White; Alan Yeats; Gordon Skipworth. Journal of the Royal Statistical Society. Series A (General), 138(3):447–449, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344602.

Kyprianou:1975:BRB

[752] E. Kyprianou. Book review: English– Greek Mathematical Dictionary, by C.

P. Tzelekis. Journal of the Royal Statistical Society. Series A (General), 138 (3):449, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344603.

Coxon:1975:BRBa

[753] A. P. M. Coxon. Book review: The Role of the Computer in Economic and Social Research in Latin America, by Nancy D. Ruggles. Journal of the Royal Statistical Society. Series A (General), 138(3):449–450, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344604.

Mann:1975:BRB

[754] Peter H. Mann. Book review: Thesaurus of Sociological Research Terminology, by C. Van de Merwe. Journal of the Royal Statistical Society. Series A (General), 138(3):450–451, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344605.

Benjamin:1975:JID

[755] B. Benjamin. Jessie Iris Douglas. Journal of the Royal Statistical Society. Series A (General), 138(3):452, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344606.

Anonymous:1975:RPc

[756] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 138(3):453–457, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344607.

Anonymous:1975:LAc

[757] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 138(3):458–461, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344608.

Anonymous:1975:BMc

[758] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 138(3):i–vi, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344609.

Anonymous:1975:VIb

[759] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 138(4):601– 604, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345209.

Anonymous:1975:FMd

[760] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 138(4):??, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345210.

Brown:1975:ENF

[761] Philip Brown and Clive Payne. Election night forecasting. Journal of the Royal Statistical Society. Series A (General), 138(4):463–498, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345211.

Anonymous:1975:SJB

[762] Anonymous. Are statistical journals becoming too theoretical? Journal of the Royal Statistical Society. Series A (General), 138(4):499–503, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/

stable/2345212. See corrigenda [866].

Pickering:1975:DEC

[763] J. F. Pickering and B. C. Isherwood. Determinants of expenditure on consumer durables. Journal of the Royal Statistical Society. Series A (General), 138(4):504–530, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345213.

McCreath:1975:PFI

[764] Margaret D. McCreath. Project on factors influencing choice of higher education. Journal of the Royal Statistical Society. Series A (General), 138 (4):531–539, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345214.

Nicholson:1975:WCL

[765] J. L. Nicholson. Whose cost of living? Journal of the Royal Statistical Society. Series A (General), 138(4): 540–542, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345215.

Ehrenberg:1975:TSC

[766] A. S. C. Ehrenberg. The teaching of statistics: Corrections and comments. Journal of the Royal Statistical Society. Series A (General), 138 (4):543–545, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345216. See

[589].

Anonymous:1975:POH

[767] Anonymous. Proceedings of the one hundred and forty-first Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 138(4):546, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2345217.

Anonymous:1975:ARC

[768] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 138 (4):547–567, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345218.

Silvey:1975:BRB

[769] S. D. Silvey. Book review: Theoretical Statistics, by D. R. Cox; D. V. Hinkley. Journal of the Royal Statistical Society. Series A (General), 138 (4):568–569, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345219.

Morgan:1975:BRBb

[770] Byron J. T. Morgan. Book review: Elementary Probability Theory with Stochastic Processes, by Kai Lai Chung. Journal of the Royal Statistical Society. Series A (General), 138 (4):570, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345220.

Haigh:1975:BRB

[771] John Haigh. Book review: Stochastic Processes, by R. Coleman. Journal of the Royal Statistical Society. Series A (General), 138(4):570–571, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345221.

Pearce:1975:BRBa

[772] S. C. Pearce. Book review: Design of Experiments: A Realistic Approach, by V. L. Anderson; R. A. McLean. Journal of the Royal Statistical Society. Series A (General), 138(4):571–572, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345222.

Pearce:1975:BRBb

[773] S. C. Pearce. Book review: Risk, Choice and Prediction, by W. J. Youden. Journal of the Royal Statistical Society. Series A (General), 138 (4):572, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345223.

Theobald:1975:BRB

[774] C. M. Theobald. Book review: Regression Estimation from Grouped Observations, by Y. Haitovsky. Journal of the Royal Statistical Society. Series A (General), 138(4):573, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345224.

Plackett:1975:BRBc

[775] R. L. Plackett. Book review: Progress in Statistics: European Meeting of Statisticians, Budapest, 1972, Vols I and II, by J. Gani; K. Sarkadi; I. Vincze. Journal of the Royal Statistical Society. Series A (General), 138 (4):573–574, ???? 1975. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345225.

Anderson:1975:BRB

[776] C. W. Anderson. Book review: Mathematical Methods in Queueing Theory, by A. B. Clarke. Journal of the Royal Statistical Society. Series A (General), 138(4):574–575, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345226.

Hiorns:1975:BRB

[777] R. W. Hiorns. Book review: Mathematical Models in Biology and Medicine, by N. T. J. Bailey; B. Sendov; R. Tsanov. Journal of the Royal Statistical Society. Series A (General), 138(4):575– 576, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345227.

Ord:1975:BRBb

[778] J. K. Ord. Book review: Characterization Problems in Mathematical Statistics, by A. M. Kagan; Yu. V. Linnik; C. R. Rao; B. Ramachandran. Journal of the Royal Statistical Society. Series A (General), 138(4):576–577, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345228.

Kingman:1975:BRBd

[779] J. F. C. Kingman. Book review: Asymptotic Expansions and the Deficiency Concept in Statistics, by W. Albers. Journal of the Royal Statistical Society. Series A (General), 138(4): 577–578, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345229.

Kingman:1975:BRBe

[780] J. F. C. Kingman. Book review: Conditional Probability Distributions, by Tue Tjur. Journal of the Royal Statistical Society. Series A (General), 138(4): 578, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345230.

Radcliffe:1975:BRB

[781] J. Radcliffe. Book review: Deterministic Threshold Models in the Theory of Epidemics, by Paul Waltman. Journal of the Royal Statistical Society. Series A (General), 138(4):578–579, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345231.

Benjamin:1975:BRBd

[782] B. Benjamin. Book review: Population Growth Estimation, by Eli S. Marks; William Seltzer; Karol J. Krotki. Journal of the Royal Statistical Society. Series A (General), 138(4):579–580, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345232.

Sumner:1975:BRB

[783] M. T. Sumner. Book review: Modelling the Economy, by G. A. Renton. Journal of the Royal Statistical Society. Series A (General), 138(4):580–581, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345233.

Barna:1975:BRB

[784] T. Barna. Book review: Economic Development and Planning. Essays in Honour of Jan Tinbergen, by W. Sellekaerts. Journal of the Royal Statistical Society. Series A (General), 138 (4):582, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345234.

Moore:1975:BRBb

[785] P. G. Moore. Book review: Business Statistics, by Example. by Alexander E. Innes. Journal of the Royal Statistical Society. Series A (General), 138 (4):583, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345235.

Goldstein:1975:BRB

[786] H. Goldstein. Book review: Social Information Processing and Statistical Systems — Change and Reform, by Edgar S. Dunn. Journal of the Royal Statistical Society. Series A (General), 138(4):583–584, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345236.

Coxon:1975:BRBb

[787] A. P. M. Coxon. Book review: The Social Grading of Occupations. A New Approach and Scale, by J. H. Goldthorpe; K. Hope. Journal of the Royal Statistical Society. Series A (General), 138(4):584–585, ???? 1975. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345237.

Lewis:1975:BRB

[788] J. Parry Lewis. Book review: The Structure, Size and Costs of Urban Settlements, by P. A. Stone. Journal of the Royal Statistical Society. Series A (General), 138(4):585–586, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345238.

Nelder:1975:RWM

[789] J. A. Nelder. Robert William MacLagan Wedderburn, 1947–1975. Journal of the Royal Statistical Society. Series A (General), 138(4):587, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345239.

Anonymous:1975:RPd

[790] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 138(4):588–594, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345240.

Anonymous:1975:LAd

[791] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 138(4):595–599, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345241.

Anonymous:1975:PSG

[792] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 138 (4):600, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345242.

Anonymous:1975:BMd

[793] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 138(4):??, ???? 1975. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345243.

Anonymous:1976:VIa

[794] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 139(1):??, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344379.

Anonymous:1976:FMa

[795] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 139(1):??, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344380.

Cunliffe:1976:I

[796] Stella V. Cunliffe. Interaction. Journal of the Royal Statistical Society. Series A (General), 139(1):1–19, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344381.

Freeman:1976:BAM

[797] P. R. Freeman. A Bayesian analysis of the megalithic yard. Journal of the Royal Statistical Society. Series A (General), 139(1):20–55, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344382.

Lindley:1976:IIM

[798] Robert M. Lindley. Inter-industry mobility of male employees in Great Britain, 1959–68. Journal of the Royal Statistical Society. Series A (General), 139(1):56–79, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344383.

Kish:1976:OPL

[799] Leslie Kish. Optima and proxima in linear sample designs. Journal of the Royal Statistical Society. Series A (General), 139(1):80–95, ???? 1976. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344384. See corrigenda

[1301].

Greenfield:1976:SAU

[800] C. C. Greenfield and S. M. Tam. A simple approximation for the upper limit to the value of a missing cell in 2 × 2 contingency table. Journal of the Royal Statistical Society. Series A (General), 139(1):96–103, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344385.

Brown:1976:RSS

[801] P. J. Brown. Remarks on some statistical methods for medical diagnosis. Journal of the Royal Statistical Society. Series A (General), 139(1):104– 107, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344386.

Hart:1976:CSD

[802] P. E. Hart. The comparative statics and dynamics of income distributions. Journal of the Royal Statistical Society. Series A (General), 139(1):108– 125, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344387.

Judge:1976:BRB

[803] Ken Judge. Book review: Reviews of United Kingdom Statistical Sources.: Volume I: Personal Social Services, by W. F. Maunder; B. P. Davies; Reviews of United Kingdom Statistical Sources.: Voluntary Organizations in the Personal Social Service Field, by W. F. Maunder; G. J. Murray. Journal of the Royal Statistical Society. Series A (General), 139(1):126–127, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344388.

Lavers:1976:BRB

[804] R. J. Lavers. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (General), 139(1):127, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344389.

Wilkinson:1976:BRBa

[805] R. K. Wilkinson. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (General), 139(1): 127–128, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344390.

Plackett:1976:BRBa

[806] R. L. Plackett. Book review: Introduction to the Theory of Statistics, by Harold J. Larson. Journal of the Royal Statistical Society. Series A (General), 139(1):128–129, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344391.

Pluck:1976:BRB

[807] R. A. Pluck. Book review: Introduction to Statistics, 2nd ed, by R. E. Walpole. Journal of the Royal Statistical Society. Series A (General), 139 (1):129–130, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344392.

Barton:1976:BRB

[808] D. E. Barton. Book review: Symmetric Functions in Statistics, by D. S. Tracey; Paul S. Dwyer. Journal of the Royal Statistical Society. Series A (General), 139(1):130, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344393.

Barrett:1976:BRB

[809] C. R. Barrett. Book review: Linear Algebra and Linear Economics, by G. Heal; G. Hughes; R. Tarling. Journal of the Royal Statistical Society. Series A (General), 139(1):130–131, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344394.

Ray:1976:BRB

[810] W. D. Ray. Book review: Generalised Inverses: Theory and Applications, by Adi Ben-Israel; Thomas N. E. Greville. Journal of the Royal Statistical Society. Series A (General), 139(1):131– 132, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344395.

Cormack:1976:BRB

[811] R. M. Cormack. Book review: Cluster Analysis: A Survey, by Benjamin S. Duran; Patrick L. Odell. Journal of the Royal Statistical Society. Series A (General), 139(1):132, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344396.

Freeman:1976:BRB

[812] G. H. Freeman. Book review: Analysis of Variance in Complex Experimental Designs, by Harold R. Lindman. Journal of the Royal Statistical Society. Series A (General), 139(1):132–133, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344397.

Pearce:1976:BRBa

[813] S. C. Pearce. Book review: Statistical Theory of the Analysis of Experimental Designs, by J. Ogawa. Journal of the Royal Statistical Society. Series A (General), 139(1):133, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344398.

Kalton:1976:BRB

[814] G. Kalton. Book review: Statistical Theory of Sample Survey Design and Analysis, by H. S. Konijn. Journal of the Royal Statistical Society. Series A (General), 139(1):133–134, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344399.

Smeed:1976:BRB

[815] R. J. Smeed. Book review: Transportation and Traffic Theory, by D. J. Buckley. Journal of the Royal Statistical Society. Series A (General), 139 (1):134–135, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344400.

MacDonald:1976:BRB

[816] K. I. MacDonald. Book review: The Use of Computers in University Social Science Departments, by John Utting; John Hall. Journal of the Royal Statistical Society. Series A (General), 139 (1):135–136, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344401.

Abrams:1976:BRB

[817] Mark Abrams. Book review: “Who’s Listening?”, by Robert Silvey. Journal of the Royal Statistical Society. Series A (General), 139(1):136–137, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344402.

Fisk:1976:BRBa

[818] P. R. Fisk. Book review: An Approach to Econometrics, by R. J. Allard. Journal of the Royal Statistical Society. Series A (General), 139(1): 137, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344403.

Mizon:1976:BRB

[819] Grayham E. Mizon. Book review: Exercises in Mathematical Economics and Econometrics, by J. E. Spencer; R. C. Geary. Journal of the Royal Statistical Society. Series A (General), 139 (1):137–138, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344404.

Matatka:1976:BRB

[820] J. Matatka. Book review: Selected Tables in Mathematical Statistics, Vol. 2, by H. L. Harter; D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 139(1):138, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344405.

Bulpitt:1976:BRB

[821] C. J. Bulpitt. Book review: Morbidity and Mortality in the United States, by C. L. Erhardt; J. E. Berlin. Journal of the Royal Statistical Society. Series A (General), 139(1):139, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344406.

Berman:1976:BRB

[822] L. S. Berman. Book review: Input-Output and Throughput. Proceedings of the 1971 Norwich Conference, by W. F. Gossling; Estimating and Projecting Input-Output Coefficients, by R. I. G. Allen; W. F. Gossling; Capital Coefficients and Dynamic Input-Output Models, by W. F. Gossling. Journal of the Royal Statistical Society. Series A (General), 139(1):139–140, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344407.

Hillmann:1976:BRB

[823] H. C. Hillmann. Book review: War and Economic Development. Essays in Memory of David Joslin, by J. M. Winter. Journal of the Royal Statistical Society. Series A (General), 139(1):140– 141, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344408.

Ford:1976:BRB

[824] J. L. Ford. Book review: The Pure Theory of International Trade under Uncertainty, by Raveendra N. Batra. Journal of the Royal Statistical Society. Series A (General), 139(1):141– 143, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344409.

Stoodley:1976:CR

[825] Keith D. C. Stoodley. Corrections: Reviews. Journal of the Royal Statistical Society. Series A (General), 139(1): 143, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344410.

Jolliffe:1976:CR

[826] Flavia R. Jolliffe. Corrections: Reviews. Journal of the Royal Statistical Society. Series A (General), 139(1): 143, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344411.

Bibby:1976:CR

[827] John Bibby. Corrections: Reviews. Journal of the Royal Statistical Society. Series A (General), 139(1):143, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344412.

Anonymous:1976:RPa

[828] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 139(1):144–155, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344413.

Anonymous:1976:LAa

[829] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 139(1):156–159, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344414.

Anonymous:1976:BMa

[830] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series A (General), 139(1):i–iv, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344415.

Anonymous:1976:FMb

[831] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 139(2):??, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345172.

Anderson:1976:MMC

[832] J. A. Anderson. Multivariate methods in clinical assessment. Journal of the Royal Statistical Society. Series A (General), 139(2):161–182, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345173.

Smith:1976:FSS

[833] T. M. F. Smith. The foundations of survey sampling: A review. Journal of the Royal Statistical Society. Series A (General), 139(2):183–204, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345174.

McClean:1976:TSM

[834] Sally McClean. The two-stage model of personnel behaviour. Journal of the Royal Statistical Society. Series A (General), 139(2):205–217, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345175.

Swee-Hock:1976:OMV

[835] Saw Swee-Hock. Occupational mortality variations in Singapore, 1970. Journal of the Royal Statistical Society. Series A (General), 139(2):218–226, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345176.

Walter:1976:UAS

[836] S. D. Walter, C. T. Miller, and J. A. H. Lee. The use of age-specific mean cohort slopes in the analysis of epidemiological incidence and mortality data. Journal of the Royal Statistical Society. Series A (General), 139(2):227– 245, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345177.

Granger:1976:TSM

[837] C. W. J. Granger and M. J. Morris. Time series modelling and interpretation. Journal of the Royal Statistical Society. Series A (General), 139(2): 246–257, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345178.

Iro:1976:MFW

[838] Madugba I. Iro. The main features of a working life table of the female labour force in Nigeria, 1965. Journal of the Royal Statistical Society. Series A (General), 139(2):258–264, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345179.

Anonymous:1976:PSG

[839] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 139 (2):265, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345180.

Silvey:1976:BRB

[840] S. D. Silvey. Book review: Introduction to Mathematical Statistics, by L. Schmetterer; Kenneth Wickwire. Journal of the Royal Statistical Society. Series A (General), 139(2):266–267, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345181.

Plackett:1976:BRBb

[841] R. L. Plackett. Book review: Introduction to Probability and Statistics, Part II: Statistics, by Narayan C. Giri. Journal of the Royal Statistical Society. Series A (General), 139(2):267, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345182.

Loynes:1976:BRB

[842] R. M. Loynes. Book review: Probability with Applications, by Michael Woodroofe. Journal of the Royal Statistical Society. Series A (General), 139 (2):267–268, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345183.

Pridmore:1976:BRB

[843] W. A. P. Pridmore. Book review: Data Reduction: Analysing and Interpreting Statistical Data, by A. S. C. Ehrenberg. Journal of the Royal Statistical Society. Series A (General), 139(2): 268–269, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345184.

Wilkinson:1976:BRBb

[844] R. K. Wilkinson. Book review: Statistics at School Level. Proceedings of the Third International Round Table Conference on the Teaching of Statistics, by Lennard Rade. Journal of the Royal Statistical Society. Series A (General), 139(2):269–270, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345185.

Sargan:1976:BRB

[845] J. D. Sargan. Book review: Studies in Bayesian Econometrics and Statistics. In Honor of Leonard J. Savage, by S. E. Fienberg; A. Zellner. Journal of the Royal Statistical Society. Series A (General), 139(2):270, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345186.

Hills:1976:BRB

[846] M. Hills. Book review: Discriminant Analysis, by P. A. Lachenbruch. Journal of the Royal Statistical Society. Series A (General), 139(2):270–271, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345187.

Lomnicki:1976:BRB

[847] Z. A. Lomnicki. Book review: Methods for Statistical Analysis of Reliability and Life Data, by Nancy R. Mann; Ray E. Schafer; Nozer D. Singpurwalla. Journal of the Royal Statistical Society. Series A (General), 139(2):271– 272, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345188.

Trustrum:1976:BRB

[848] Kathleen Trustrum. Book review: Problems in Linear and Nonlinear Programming, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 139(2):273, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345189.

Chambers:1976:BRBa

[849] M. L. Chambers. Book review: Introduction to Operations Research, by Frederick S. Hillier; Gerald J. Lieberman. Journal of the Royal Statistical Society. Series A (General), 139(2): 273, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345190.

Ord:1976:BRB

[850] Keith Ord. Book review: Display and Analysis of Spatial Data, by John C. Davis; Michael J. McCullagh. Journal of the Royal Statistical Society. Series A (General), 139(2):274, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345191.

Thatcher:1976:BRB

[851] A. R. Thatcher. Book review: Cost of Living Index Numbers, by Kali S. Banerjee. Journal of the Royal Statistical Society. Series A (General), 139 (2):274–275, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345192.

Bartlett:1976:BRB

[852] M. S. Bartlett. Book review: Stochastic Population Theories, by D. Ludwig. Journal of the Royal Statistical Society. Series A (General), 139(2):275– 276, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345193.

Vajda:1976:BRBa

[853] S. Vajda. Book review: Introduction to Minimax, by V. F. Dem’yanov; V. N. Malozemov; D. Louvish. Journal of the Royal Statistical Society. Series A (General), 139(2):276, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345194.

Lancaster:1976:BRB

[854] T. Lancaster. Book review: Introduction to Structural Equation Models, by O. D. Duncan. Journal of the Royal Statistical Society. Series A (General), 139(2):276–277, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2345195.

Ripley:1976:BRB

[855] B. D. Ripley. Book review: Random Sets and Integral Geometry, by G. Matheron. Journal of the Royal Statistical Society. Series A (General), 139 (2):277–278, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345196.

Harris:1976:BRB

[856] R. R. Harris. Book review: Estimation Theory, by Demetrios G. Lainiotis. Journal of the Royal Statistical Society. Series A (General), 139(2):278– 279, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345197.

Nove:1976:BRB

[857] Alec Nove. Book review: The Mathematical Revolution in Soviet Economics, by A. Zauberman. Journal of the Royal Statistical Society. Series A (General), 139(2):279–280, ????

1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345198.

Bailey:1976:BRB

[858] B. J. R. Bailey. Book review: Mathematical Problems in Biology: Victoria Conference, 1973, by Pauline Van den Driessche. Journal of the Royal Statistical Society. Series A (General), 139(2):280, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345199.

Anonymous:1976:BRB

[859] Anonymous. Book review: Statistics of Politics, by Philip McGhee. Journal of the Royal Statistical Society. Series A (General), 139(2):280–281, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345200.

Chedzoy:1976:BRB

[860] O. B. Chedzoy. Book review: Vital Statistics, by Arthur Gore. Journal of the Royal Statistical Society. Series A (General), 139(2):281, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345201.

Grant:1976:SRH

[861] A. T. K. Grant. Sir Ralph Hawtrey, 1879–1975. Journal of the Royal Statistical Society. Series A (General), 139 (2):282–283, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345202.

Deeming:1976:DCR

[862] W. Edwards Deeming. Donald Crockett Riley, 1905–1973. Journal of the Royal Statistical Society. Series A (General), 139(2):283, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345203.

Anonymous:1976:RPb

[863] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 139(2):284–294, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345204.

Anonymous:1976:LAb

[864] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 139(2):295–297, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345205.

Anonymous:1976:CI

[865] Anonymous. Corrigenda: Interaction. Journal of the Royal Statistical Society. Series A (General), 139(2):298, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345206.

Anonymous:1976:CSJ

[866] Anonymous. Corrigenda: Are statistical journals becoming too theoretical. Journal of the Royal Statistical Society. Series A (General), 139(2):298, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345207. See [762].

Anonymous:1976:BMb

[867] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 139(2):i–iv, ????

1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345208.

Anonymous:1976:FMc

[868] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 139(3):??, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344837.

Gascoigne:1976:DBS

[869] J. C. Gascoigne and I. D. Hill. The draft British Standard 6002: “Sampling inspection by variables”. Journal of the Royal Statistical Society. Series A (General), 139(3):299–317, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344838.

Barnett:1976:OMD

[870] V. Barnett. The ordering of multivariate data. Journal of the Royal Statistical Society. Series A (General), 139 (3):318–355, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344839.

Downton:1976:CSB

[871] F. Downton and Carmen Lockwood. Computer studies of baccarat, II: Baccarat-Banque. Journal of the Royal Statistical Society. Series A (General), 139(3):356–364, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344840.

Drettakis:1976:DLM

[872] E. G. Drettakis. Distributed lag models for the quarterly migration flows of West Germany, 1962–72. Journal of the Royal Statistical Society. Series A (General), 139(3):365–373, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344841.

Cook:1976:HUG

[873] W. R. Cook. How the University Grants Committee determines allocations of recurrent grants — a curious correlation. Journal of the Royal Statistical Society. Series A (General), 139 (3):374–384, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344842. See

[969] and reply [1052].

Conway:1976:WSW

[874] Freda Conway. What is statistics and who should teach it? Journal of the Royal Statistical Society. Series A (General), 139(3):385–388, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344843.

Greenfield:1976:RPD

[875] C. C. Greenfield. A revised procedure for dual record systems in estimating vital events. Journal of the Royal Statistical Society. Series A (General), 139 (3):389–401, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344844.

Plackett:1976:BRBc

[876] R. L. Plackett. Book review: Discrete Multivariate Analysis: Theory and Practice, by Yvonne M. M. Bishop; Stephen E. Fienberg; Paul W. Holland. Journal of the Royal Statistical Society. Series A (General), 139(3):402– 403, ???? 1976. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2344845.

Kingman:1976:BRBa

[877] J. F. C. Kingman. Book review: Theory of Probability, A Critical Introductory Treatment, by Bruno de Finetti; Antonio Machi; Adrian F. M. Smith. Journal of the Royal Statistical Society. Series A (General), 139(3):403, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344846.

Collings:1976:BRB

[878] S. N. Collings. Book review: Introduction to Statistics, by H. Robbins; J. Van Ryzin. Journal of the Royal Statistical Society. Series A (General), 139 (3):403–404, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344847.

Huitson:1976:BRBa

[879] A. Huitson. Book review: Stat Lab: An Empirical Introduction to Statistics, by J. L. Hodges; David Krech; Richard S. Crutchfield. Journal of the Royal Statistical Society. Series A (General), 139 (3):404–405, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344848.

Hammerton:1976:BRB

[880] M. Hammerton. Book review: Introduction to Statistics for Psychology and Education, by J. C. Nunally. Journal of the Royal Statistical Society. Series A (General), 139(3):405, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344849.

Marriott:1976:BRB

[881] F. H. C. Marriott. Book review: Analyse des Données Multi-Dimensionelles, by P. Bertier; J.-M. Bonroche. Journal of the Royal Statistical Society. Series A (General), 139(3):405–406, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344850.

Pearce:1976:BRBb

[882] S. C. Pearce. Book review: A Survey of Statistical Design and Linear Models, by J. N. Srivastava. Journal of the Royal Statistical Society. Series A (General), 139(3):406, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344851.

Grimshaw:1976:BRB

[883] J. J. Grimshaw. Book review: Normal Values in Clinical Chemistry: A Guide to Statistical Analysis of Laboratory Data, by H. F. Martin; B. J. Gudzinowicz; H. Fanger. Journal of the Royal Statistical Society. Series A (General), 139(3):406–407, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344852.

Stone:1976:BRB

[884] M. Stone. Book review: Concepts in Geostatistics, by Richard B. McCammon. Journal of the Royal Statistical Society. Series A (General), 139(3): 407, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344853.

Fisk:1976:BRBb

[885] P. R. Fisk. Book review: Studies in Correlation, by Sydney Afriat; M. V.

Rama Sastry; Gerhard Tintner. Journal of the Royal Statistical Society. Series A (General), 139(3):407–408, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344854.

Hawkes:1976:BRB

[886] John Hawkes. Book review: Advances in Probability and Related Topics, by P. Ney; S. Port. Journal of the Royal Statistical Society. Series A (General), 139(3):408–409, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344855.

Whittle:1976:BRB

[887] P. Whittle. Book review: Stochastic Differential Equations: Theory and Applications, by L. Arnold. Journal of the Royal Statistical Society. Series A (General), 139(3):409, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344856.

Chambers:1976:BRBb

[888] M. L. Chambers. Book review: Examples and Exercises in Operations Research, by A. Ghosal; S. G. Loo; N. Singh. Journal of the Royal Statistical Society. Series A (General), 139(3): 410, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344857.

Ashford:1976:BRB

[889] J. R. Ashford. Book review: Systems Aspects of Health Planning, by Norman T. J. Bailey; M. Thompson. Journal of the Royal Statistical Society. Series A (General), 139(3):410–411, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344858.

Yannopoulos:1976:BRB

[890] G. N. Yannopoulos. Book review: Input-Output Analysis and its Applications, by R. O’Connor; E. W. Henry. Journal of the Royal Statistical Society. Series A (General), 139(3):411, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344859.

Hart:1976:BRB

[891] P. E. Hart. Book review: Royal Commission on the Distribution of Income and Wealth. Report No. 1. Initial Report on the Standing Reference. Journal of the Royal Statistical Society. Series A (General), 139(3):412, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344860.

Hill:1976:BRB

[892] Austin Bradford Hill. Book review: The Registrar-General’s Statistical Review of England and Wales for the Three Years 1968–70. Supplement on Cancer, by O. P. C. S. Journal of the Royal Statistical Society. Series A (General), 139(3):412–413, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344861.

Winterbottom:1976:BRB

[893] Alan Winterbottom. Book review: Statistics and Reliability, by Roger Harrison. Journal of the Royal Statistical Society. Series A (General), 139 (3):413–414, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344862.

Wilkinson:1976:BRBc

[894] R. K. Wilkinson. Book review: Housing, by R. Thomas; R. Peacock. Journal of the Royal Statistical Society. Series A (General), 139(3):414, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344863.

Fuller:1976:BRB

[895] Mike Fuller. Book review: Statistical Sources: Unit 7, Production; Unit 8, Consumption; Unit 9, Investment; Unit 10, Labour; Unit 11, Inflation; Unit 12, Income Distribution, by Thelma Liesner; Colin Leicester; Frank Blackaby; Andrew Dean; Fred Johnson; Milton Keynes. Journal of the Royal Statistical Society. Series A (General), 139(3):414–415, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344864.

Pridmore:1976:AS

[896] W. A. Pridmore. Albert Swan 1892– 1975. Journal of the Royal Statistical Society. Series A (General), 139(3): 416, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344865.

Williams:1976:MHB

[897] E. J. Williams. Maurice Henry Belz 1897–1975. Journal of the Royal Statistical Society. Series A (General), 139 (3):417, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344866.

Anonymous:1976:RPc

[898] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 139(3):418–427, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344867.

Anonymous:1976:LAc

[899] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 139(3):428–430, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344868.

Anonymous:1976:BMc

[900] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 139(3):i–iv, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344869.

Anonymous:1976:VIb

[901] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 139(4):572– 574, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344347.

Anonymous:1976:FMd

[902] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 139(4):??, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344348.

Atkinson:1976:CGB

[903] A. C. Atkinson and M. C. Pearce. The computer generation of beta, gamma and normal random variables. Journal of the Royal Statistical Society. Series A (General), 139(4):431–461, ????

1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344349.

Atkinson:1976:SAG

[904] A. C. Atkinson and J. Whittaker. A switching algorithm for the generation of beta random variables with at least one parameter less than 1. Journal of the Royal Statistical Society. Series A (General), 139(4):462–467, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344350.

Prothero:1976:MMT

[905] David L. Prothero and Kenneth F. Wallis. Modelling macroeconomic time series. Journal of the Royal Statistical Society. Series A (General), 139(4): 468–500, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344351.

McClean:1976:SMC

[906] Sally McClean. Some models for company growth. Journal of the Royal Statistical Society. Series A (General), 139 (4):501–507, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344352.

Fiegehen:1976:MPN

[907] G. C. Fiegehen and P. S. Lansley. The measurement of poverty: A note on household size and income units. Journal of the Royal Statistical Society. Series A (General), 139(4):508–518, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344353.

Anonymous:1976:POH

[908] Anonymous. Proceedings of the one hundred and forty-second Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 139 (4):519, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344354.

Anonymous:1976:ARC

[909] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 139 (4):520–540, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344355.

Haigh:1976:BRB

[910] J. Haigh. Book review: Random Variables, by L. E. Clarke. Journal of the Royal Statistical Society. Series A (General), 139(4):541–542, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344356.

Oldham:1976:BRB

[911] P. D. Oldham. Book review: Sequential Medical Trials, by P. Armitage. Journal of the Royal Statistical Society. Series A (General), 139(4):542–543, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344357.

Moore:1976:BRB

[912] P. G. Moore. Book review: Management Decision and Decision Analysis, by Colin Eden; John Harris. Journal of the Royal Statistical Society. Series A (General), 139(4):543, ????

1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344358.

Haley:1976:BRB

[913] K. B. Haley. Book review: Fundamentals of Operations Research for Management: An Introduction to Quantitative Methods, by Shiv K. Gupta; John M. Cozzolino. Journal of the Royal Statistical Society. Series A (General), 139 (4):543–544, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344359.

Losty:1976:BRB

[914] P. A. Losty. Book review: An Introduction to General Systems Thinking, by Gerald M. Weinberg. Journal of the Royal Statistical Society. Series A (General), 139(4):544–545, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344360.

Barbour:1976:BRB

[915] A. D. Barbour. Book review: Sums of Independent Random Variables, by V. V. Petrov; A. A. Brown. Journal of the Royal Statistical Society. Series A (General), 139(4):545, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344361.

Vajda:1976:BRBb

[916] S. Vajda. Book review: Mathematical Description of Linear Systems, by Wilson J. Rugh. Journal of the Royal Statistical Society. Series A (General), 139 (4):545–546, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344362.

Bather:1976:BRB

[917] J. A. Bather. Book review: Deterministic and Stochastic Optimal Control, by W. H. Fleming; R. W. Rishel. Journal of the Royal Statistical Society. Series A (General), 139(4):546, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344363.

Lawrance:1976:BRB

[918] A. J. Lawrance. Book review: Random Point Processes, by Donald L. Snyder. Journal of the Royal Statistical Society. Series A (General), 139(4):547, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344364.

Kingman:1976:BRBb

[919] J. F. C. Kingman. Book review: Discrete-Parameter Martingales, by Jacques Neveu. Journal of the Royal Statistical Society. Series A (General), 139(4):547–548, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344365.

Conway:1976:BRB

[920] Freda Conway. Book review: A Programmed Text in Statistics.: Book One: Summarising Data, by J. Hine; G. B. Wetherill; A Programmed Text in Statistics.: Book Two: Basic Theory. by J. Hine; G. B. Wetherill; A Programmed Text in Statistics.: Book Three: The t-Test and χ^{2} Goodness of Fit. by J. Hine; G. B. Wetherill; A Programmed Text in Statistics.: Book Four: Tests on Variance and Regression. by J. Hine; G. B. Wetherill. Journal of the Royal Statistical Society. Series A (General), 139(4):548, ???? 1976. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344366.

Goldstein:1976:BRB

[921] Harvey Goldstein. Book review: Applied Multiple Regression/Correlation Analysis for the Behavioural Sciences, by Jacob Cohen; Patricia Cohen. Journal of the Royal Statistical Society. Series A (General), 139(4):549, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344367.

Fielding:1976:BRB

[922] A. Fielding. Book review: Inferences from Sociological Survey Data: A Unified Approach, by J. K. Lindsey. Journal of the Royal Statistical Society. Series A (General), 139(4):549–551, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344368.

Huitson:1976:BRBb

[923] A. Huitson. Book review: Statistical Models in Applied Science, by Karl V. Bury. Journal of the Royal Statistical Society. Series A (General), 139(4): 551, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344369.

Sprent:1976:BRB

[924] P. Sprent. Book review: Introduction to Statistics and Computer Programming, by Carl F. Kossach; Claudia I. Henschke. Journal of the Royal Statistical Society. Series A (General), 139 (4):551–552, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344370.

Gilje:1976:BRB

[925] E. K. Gilje. Book review: Towards a System of Social and Demographic Statistics, by Richard Stone. Journal of the Royal Statistical Society. Series A (General), 139(4):552–553, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344371.

Fisk:1976:BRBc

[926] P. R. Fisk. Book review: Index Numbers in Theory and Practice, by R. G. D. Allen. Journal of the Royal Statistical Society. Series A (General), 139 (4):553–554, ???? 1976. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344372.

Chatfield:1976:BRB

[927] C. Chatfield. Book review: Seasonality in Tourism, by R. R. V. Baron. Journal of the Royal Statistical Society. Series A (General), 139(4):554, ???? 1976. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344373.

Coppock:1976:BRB

[928] J. T. Coppock. Book review: Reviews of United Kingdom Statistical Sources, by W. F. Maunder; Reviews of United Kingdom Statistical Sources. Volume IV: Leisure, by F. M. M. Lewes; S. R. Parker; Tourism, by L. J. Lickorish. Journal of the Royal Statistical Society. Series A (General), 139(4):554– 555, ???? 1976. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344374.

Schagen:1976:BRB

[929] I. P. Schagen. Book review: The Search for Oil: Some Statistical Methods and Techniques, by D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 139(4):555, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344375.

Anonymous:1976:RPd

[930] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 139(4):556–568, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344376.

Anonymous:1976:LAd

[931] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 139(4):569–571, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344377.

Anonymous:1976:BMd

[932] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 139(4):i–iv, ???? 1976. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344378.

Anonymous:1977:VIa

[933] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 140(1):??, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344514.

Anonymous:1977:FMa

[934] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 140(1):i–vi, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344515.

Lawrance:1977:SMR

[935] A. J. Lawrance and N. T. Kottegoda. Stochastic modelling of riverflow time series. Journal of the Royal Statistical Society. Series A (General), 140(1):1– 47, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344516.

Nelder:1977:RLM

[936] J. A. Nelder. A reformulation of linear models. Journal of the Royal Statistical Society. Series A (General), 140(1):48– 77, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344517.

McCarthy:1977:EVT

[937] Colm McCarthy and Terence M. Ryan. Estimates of voter transition probabilities from the British General Elections of 1974. Journal of the Royal Statistical Society. Series A (General), 140(1):78– 85, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344518.

Upton:1977:MMV

[938] G. J. G. Upton. A memory model for voting transitions in British elections. Journal of the Royal Statistical Society. Series A (General), 140(1):86–94, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344519.

Barnett:1977:BRBa

[939] V. Barnett. Book review: The Logical Foundations of Statistical Inference, by Henry E. Kyburg. Journal of the Royal Statistical Society. Series A (General), 140(1):95–96, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344520.

Oliver:1977:BRBa

[940] F. R. Oliver. Book review: Bayesian Analysis, Probability and Decision, by Hans Werner Gottinger. Journal of the Royal Statistical Society. Series A (General), 140(1):96, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344521.

Loynes:1977:BRB

[941] R. M. Loynes. Book review: Time Series Analysis of Forecasting: The Box– Jenkins Approach, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 140(1):97, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344522.

Cressie:1977:BRB

[942] Noel Cressie. Book review: Lecture Notes on Queueing Systems, by Brian Conolly. Journal of the Royal Statistical Society. Series A (General), 140 (1):97–98, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344523.

Lury:1977:BRB

[943] D. A. Lury. Book review: Population Density and Concentration in Great Britain 1931, 1951 and 1961, by J.

Craig. Journal of the Royal Statistical Society. Series A (General), 140(1):98– 99, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344524.

Gray:1977:BRBa

[944] J. R. Gray. Book review: Credibility– Theory and Applications, by P. M. Kahn. Journal of the Royal Statistical Society. Series A (General), 140(1): 99–100, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344525.

Chambers:1977:BRB

[945] M. L. Chambers. Book review: Practical Aspects of Forecasting, by M. A. Gordon. Journal of the Royal Statistical Society. Series A (General), 140 (1):100–101, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344526.

Holgate:1977:BRB

[946] P. Holgate. Book review: Branching Processes with Biological Applications, by P. Jagers. Journal of the Royal Statistical Society. Series A (General), 140 (1):101, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344527.

Stern:1977:BRB

[947] G. J. A. Stern. Book review: Statistische Tafeln zur Multivariaten Analysis: Ein Handbuch mit Hinweisen zur Anwendung, by Heinz Kres. Journal of the Royal Statistical Society. Series A (General), 140(1):102, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344528.

Bulmer:1977:BRB

[948] M. G. Bulmer. Book review: Contributions to Applied Statistics, by W. J. Ziegler. Journal of the Royal Statistical Society. Series A (General), 140 (1):102, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344529.

Flowerdew:1977:BRB

[949] A. D. J. Flowerdew. Book review: Noise and Prices, by A. A. Walters. Journal of the Royal Statistical Society. Series A (General), 140(1):103, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344530.

Downton:1977:BRB

[950] F. Downton. Book review: Statistics in the Real World, by Richard J. Larson; Donna Fox Stroup. Journal of the Royal Statistical Society. Series A (General), 140(1):103–104, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344531.

Coleman:1977:BRB

[951] Rodney Coleman. Book review: Probability, Statistics and Time. A Collection of Essays, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 140(1):105, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344532.

Turner:1977:BRB

[952] Robert Turner. Book review: International Recommendations on Labour Statistics, by International Labour Office. Journal of the Royal Statistical Society. Series A (General), 140(1):105, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344533.

Marshall:1977:BRB

[953] G. P. Marshall. Book review: Economics of the Family: Marriage, Children and Human Capital, by T. W. Schultz. Journal of the Royal Statistical Society. Series A (General), 140 (1):106–108, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344534.

Haigh:1977:BRBa

[954] John Haigh. Book review: Limit Theorems of Probability Theory, by P. Revesz. Journal of the Royal Statistical Society. Series A (General), 140 (1):108, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344535.

Haigh:1977:BRBb

[955] John Haigh. Book review: Nonparametrics. Statistical Methods Based on Ranks, by E. L. Lehmann. Journal of the Royal Statistical Society. Series A (General), 140(1):108–109, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344536.

Fielding:1977:BRB

[956] A. Fielding. Book review: Statistics: The Essentials for Research, by Henry E. Klugh. Journal of the Royal Statistical Society. Series A (General), 140 (1):109, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344537.

Paine:1977:BRB

[957] G. Paine. Book review: Vital Statistics: A Memorial Volume of Selections from the Reports and Writings of William Farr, by Mervyn Susser; Abraham Adelstein. Journal of the Royal Statistical Society. Series A (General), 140(1):109–110, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344538.

Ross:1977:BRBa

[958] G. J. S. Ross. Book review: Methods of Optimization, by G. R. Walsh. Journal of the Royal Statistical Society. Series A (General), 140(1):110, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344539.

Wilson:1977:HL

[959] Harold Wilson. Hector Leak, 1887– 1976. Journal of the Royal Statistical Society. Series A (General), 140(1): 111, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344540.

Linder:1977:HK

[960] Arthur Linder. Hans Kellerer, 1902– 1976. Journal of the Royal Statistical Society. Series A (General), 140(1): 111–112, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344541.

Anonymous:1977:RPa

[961] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 140(1):113–124, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344542.

Anonymous:1977:LAa

[962] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 140(1):125–126, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344543.

Anonymous:1977:BMa

[963] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 140(1):i–vi, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344544.

Anonymous:1977:FMb

[964] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 140(2):??, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344870.

Anonymous:1977:TSA

[965] Anonymous. A tribute to Sir Austin Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 140 (2):127–128, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344871.

Moore:1977:MSU

[966] P. G. Moore. The manager’s struggles with uncertainty. Journal of the Royal Statistical Society. Series A (General), 140(2):129–165, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344872.

Cooper:1977:ASS

[967] B. E. Cooper. Advances in statistical system design. Journal of the Royal Statistical Society. Series A (General), 140(2):166–198, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344873.

Dainton:1977:C

[968] Frederick Dainton. [comments]. Journal of the Royal Statistical Society. Series A (General), 140(2):199, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344874.

Barnard:1977:HUG

[969] G. A. Barnard. How the University Grants Committee determines allocations of university grants — a case study in statistical fallacies. Journal of the Royal Statistical Society. Series A (General), 140(2):200–202, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344875. See [873].

Green:1977:AUG

[970] P. J. Green and C. Chatfield. The allocation of university grants. Journal of the Royal Statistical Society. Series A (General), 140(2):202–209, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344876.

Anonymous:1977:ERS

[971] Anonymous. Evidence from the Royal Statistical Society to the data protection committee. Journal of the Royal Statistical Society. Series A (General), 140(2):210–216, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344877.

Fisk:1977:SAI

[972] P. R. Fisk. Some approximations to an “ideal” index number. Journal of the Royal Statistical Society. Series A (General), 140(2):217–231, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344878.

Atkinson:1977:EPA

[973] A. C. Atkinson. An easily programmed algorithm for generating gamma random variables. Journal of the Royal Statistical Society. Series A (General), 140(2):232–234, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344879.

Plackett:1977:BRBa

[974] R. L. Plackett. Book review: An Introduction to Probability Theory and Mathematical Statistics, by V. K. Rohatgi. Journal of the Royal Statistical Society. Series A (General), 140(2): 235–236, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344880.

Haigh:1977:BRBc

[975] John Haigh. Book review: Normal Approximations and Asymptotic Expansions, by R. N. Battacharya; R. Ranga Rao. Journal of the Royal Statistical Society. Series A (General), 140(2): 236, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344881.

Chatfield:1977:BRBa

[976] C. Chatfield. Book review: Fourier Analysis of Time Series: An Introduction, by P. Bloomfield. Journal of the Royal Statistical Society.

Series A (General), 140(2):237, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344882.

Barnett:1977:BRBb

[977] V. Barnett. Book review: The Advanced Theory of Statistics. Vol. 3. Design and Analysis, and Time Series, by Maurice Kendall; A. Stuart. Journal of the Royal Statistical Society. Series A (General), 140(2):237–238, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344883.

Bartlett:1977:BRB

[978] M. S. Bartlett. Book review: Studies in Probability and Statistics. Papers in Honour of Edwin J. G. Pitman, by E. J. Williams. Journal of the Royal Statistical Society. Series A (General), 140 (2):238–239, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344884.

Gray:1977:BRBb

[979] J. R. Gray. Book review: A First Course in Probability, by Sheldon Ross. Journal of the Royal Statistical Society. Series A (General), 140(2):239– 240, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344885.

Hornstein:1977:BRBa

[980] Z. Hornstein. Book review: Elementary Statistical Concepts, by R. E. Walpole. Journal of the Royal Statistical Society. Series A (General), 140(2):240, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344886.

Bibby:1977:BRBa

[981] J. Bibby. Book review: Introduction to Statistics and Probability, by E. J. Dudewicz. Journal of the Royal Statistical Society. Series A (General), 140 (2):240, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344887.

Prentice:1977:BRB

[982] M. J. Prentice. Book review: Statistical Theory, by B. W. Lindgren. Journal of the Royal Statistical Society. Series A (General), 140(2):241, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344888.

Pridmore:1977:BRB

[983] W. A. Pridmore. Book review: Statistics in Small Doses, by Winifred M. Castle. Journal of the Royal Statistical Society. Series A (General), 140 (2):241, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344889.

Marriott:1977:BRBa

[984] F. H. C. Marriott. Book review: Elementary Statistics, by P. G. Hoel. Journal of the Royal Statistical Society. Series A (General), 140(2):241–242, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344890.

Freeman:1977:BRBa

[985] P. R. Freeman. Book review: Introduction to Statistics, by Robert Fried. Journal of the Royal Statistical Society. Series A (General), 140(2):242, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344891.

Moore:1977:BRBa

[986] P. G. Moore. Book review: Statistics for Business, Finance and Accounting, by J. P. Dickinson. Journal of the Royal Statistical Society. Series A (General), 140(2):242–243, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344892.

Kanji:1977:BRB

[987] G. K. Kanji. Book review: Statistics Made Relevant: A Case Book of Real-Life Examples, by P. Baum; E. M. Scheuer. Journal of the Royal Statistical Society. Series A (General), 140 (2):243–244, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344893.

Woolf:1977:BRB

[988] Myra Woolf. Book review: Statistical Calculations: How to Solve Statistical Problems, by E. J. Middlebrooks. Journal of the Royal Statistical Society. Series A (General), 140(2):244, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344894.

Bartholomew:1977:BRB

[989] D. J. Bartholomew. Book review: Markov Chains: Theory and Applications, by D. L. Isaacson; R. W. Madsen. Journal of the Royal Statistical Society. Series A (General), 140(2):244– 245, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344895.

Foster:1977:BRB

[990] F. G. Foster. Book review: Queueing Models Involving Buffers, by F. Göbel. Journal of the Royal Statistical Society. Series A (General), 140(2):245, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344896.

Titterington:1977:BRB

[991] D. M. Titterington. Book review: Invariance and Minimax Statistical Tests, by N. Giri. Journal of the Royal Statistical Society. Series A (General), 140 (2):245–246, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344897.

Ross:1977:BRBb

[992] G. J. S. Ross. Book review: Analyse Statistique à Plusieurs Variables, by Pierre Dagnelie. Journal of the Royal Statistical Society. Series A (General), 140(2):246, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344898.

Hills:1977:BRB

[993] M. Hills. Book review: An Introduction to Multivariate Techniques for Social and Behavioural Sciences, by Spencer Bennett; David Bowers. Journal of the Royal Statistical Society. Series A (General), 140(2):246–247, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344899.

Marriott:1977:BRBb

[994] F. H. C. Marriott. Book review: Computational Methods of Multivariate Analysis in Physical Geography, by P.

M. Mather. Journal of the Royal Statistical Society. Series A (General), 140 (2):247–248, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344900.

Ord:1977:BRB

[995] Keith Ord. Book review: The Statistical Analysis of Spatial Pattern, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 140 (2):248, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344901.

Freeman:1977:BRBb

[996] P. R. Freeman. Book review: Spatial Analysis in Archaeology, by Ian Hodder; Clive Orton. Journal of the Royal Statistical Society. Series A (General), 140(2):248–249, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344902.

Bissell:1977:BRB

[997] Derek Bissell. Book review: Statistical Quality Control Methods, by Irving W. Burr. Journal of the Royal Statistical Society. Series A (General), 140 (2):249–250, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344903.

Robinson:1977:BRB

[998] Colin Robinson. Book review: Forecasting for Business: Methods and Applications, by D. Wood; R. Fildes. Journal of the Royal Statistical Society. Series A (General), 140(2):250– 251, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344904.

Moore:1977:BRBb

[999] P. G. Moore. Book review: Concepts and Applications of Decision Models, by V. M. Rao Tummula; Richard G. Henshaw. Journal of the Royal Statistical Society. Series A (General), 140 (2):251–252, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344905.

Hornstein:1977:BRBb

[1000] Z. Hornstein. Book review: General Sources of Statistics, by G. F. Lock. Journal of the Royal Statistical Society. Series A (General), 140(2):252– 253, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344906.

Plackett:1977:BRBb

[1001] R. L. Plackett. Book review: On the History of Statistics and Probability, by D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 140 (2):253, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344907.

Tippett:1977:BRB

[1002] L. H. C. Tippett. Book review: The Statistical Movement in Early Victorian Britain, by Michael J. Cullen. Journal of the Royal Statistical Society. Series A (General), 140(2):253– 254, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344908.

Land:1977:BRB

[1003] Ailsa Land. Book review: Operations Research ’75, by K. B. Haley. Journal of the Royal Statistical Society. Series A (General), 140(2):254–255, ????

1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344909.

Bunday:1977:BRB

[1004] B. D. Bunday. Book review: Managing Mathematically, by Terry Green; John Webster. Journal of the Royal Statistical Society. Series A (General), 140 (2):255–256, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344910.

Nicholson:1977:BRBa

[1005] R. J. Nicholson. Book review: Understanding Econometrics, by Jon Stewart. Journal of the Royal Statistical Society. Series A (General), 140(2):256, ???? 1977. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344911.

Sargan:1977:BRBa

[1006] J. D. Sargan. Book review: Econometrics, by Peter Schmidt. Journal of the Royal Statistical Society. Series A (General), 140(2):257, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344912.

Moore:1977:BRBc

[1007] P. G. Moore. Book review: The Delphi Method: Techniques and Applications, by H. A. Lingstone; M. Turoff. Journal of the Royal Statistical Society. Series A (General), 140(2):257–258, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344913.

Fessey:1977:BRB

[1008] M. C. Fessey. Book review: Trade Statistics and the Balance of Payments, by Geoff Fogg. Journal of the Royal Statistical Society. Series A (General), 140(2):258–259, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344914.

Nevin:1977:BRB

[1009] Edward Nevin. Book review: International Comparisons of Levels of Development, by Geraldina del Drago. Journal of the Royal Statistical Society. Series A (General), 140(2):259, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344915.

Spitz:1977:BRB

[1010] Kathleen Spitz. Book review: Introduction to Designing Research and Evaluation, by Clifford J. Drew. Journal of the Royal Statistical Society. Series A (General), 140(2):260, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344916.

Taylor:1977:LHF

[1011] Wallis Taylor. Lancelot Hogben, F.R.S. (1895–1975). Journal of the Royal Statistical Society. Series A (General), 140 (2):261–262, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344917.

Anonymous:1977:RPb

[1012] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 140(2):263–272, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344918.

Anonymous:1977:LAb

[1013] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 140(2):273–275, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344919.

Anonymous:1977:BMb

[1014] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 140(2):i–ii, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344920.

Anonymous:1977:FMc

[1015] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 140(3):??, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344921.

Ehrenberg:1977:RN

[1016] A. S. C. Ehrenberg. Rudiments of numeracy. Journal of the Royal Statistical Society. Series A (General), 140 (3):277–297, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344922.

Mahon:1977:SDI

[1017] B. H. Mahon. Statistics and decisions: The importance of communication and the power of graphical presentation. Journal of the Royal Statistical Society. Series A (General), 140 (3):298–323, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344923.

Eaves:1977:ICH

[1018] Lindon J. Eaves. Inferring the causes of human variation. Journal of the Royal Statistical Society. Series A (General), 140(3):324–355, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344924.

Anonymous:1977:ACT

[1019] Anonymous. Advisory Committee on Trunk Road Assessment. Journal of the Royal Statistical Society. Series A (General), 140(3):356–358, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344925.

Krishnan:1977:LWL

[1020] P. Krishnan. The length of working life: India, 1971. Journal of the Royal Statistical Society. Series A (General), 140 (3):359–365, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344926.

Benjamin:1977:PMS

[1021] B. Benjamin. Progress in medical statistics. Journal of the Royal Statistical Society. Series A (General), 140 (3):366–376, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344927.

Beesley:1977:BRB

[1022] Ian Beesley. Book review: Setting Statistical Priorities — A Report, by U.S. National Research Council. Journal of the Royal Statistical Society. Series A (General), 140(3):377–378, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344928.

Marriott:1977:BRBc

[1023] F. H. C. Marriott. Book review: Multivariate Analysis in Behavioural Research, by A. E. Maxwell. Journal of the Royal Statistical Society. Series A (General), 140(3):378, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344929.

Clarke:1977:BRB

[1024] G. M. Clarke. Book review: Multivariate Statistical Methods, by D. F. Morrison. Journal of the Royal Statistical Society. Series A (General), 140 (3):378–379, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344930.

Chatfield:1977:BRBb

[1025] C. Chatfield. Book review: Introduction to Statistical Time Series, by W. A. Fuller. Journal of the Royal Statistical Society. Series A (General), 140 (3):379–380, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344931.

Ray:1977:BRBa

[1026] W. D. Ray. Book review: Statistical Forecasting, by Warren Gilchrist. Journal of the Royal Statistical Society. Series A (General), 140(3):380, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344932.

Oliver:1977:BRBb

[1027] F. R. Oliver. Book review: Statistics for Beginners, by S. O. Adamu; T. L. Johnson. Journal of the Royal Statistical Society. Series A (General), 140 (3):380–381, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344933.

Keen:1977:BRB

[1028] Joan Keen. Book review: Statistics: A Beginning, by R. R. Kuebler; H. Smith. Journal of the Royal Statistical Society. Series A (General), 140 (3):381–382, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344934.

Wilkinson:1977:BRB

[1029] R. K. Wilkinson. Book review: Introductory Economic Statistics, by Ann C. Mayes; David G. Mayes. Journal of the Royal Statistical Society. Series A (General), 140(3):382–383, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344935.

Smith:1977:BRBa

[1030] T. M. F. Smith. Book review: Basic Ideas of Scientific Sampling, by Alan Stuart. Journal of the Royal Statistical Society. Series A (General), 140 (3):383, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344936.

Ray:1977:BRBb

[1031] W. D. Ray. Book review: Handbook of Statistical Distributions, by Jagdish K. Patel; C. H. Kapadia; D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 140(3):383, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344937.

Ray:1977:BRBc

[1032] W. D. Ray. Book review: Studies in Non-linear Estimation, by Stephen M. Goldfield; Richard E. Quandt. Journal of the Royal Statistical Society. Series A (General), 140(3):383–384, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344938.

Nelder:1977:BRB

[1033] J. A. Nelder. Book review: Theory and Application of the Linear Model, by Franklin A. Graybill. Journal of the Royal Statistical Society. Series A (General), 140(3):384–385, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344939.

Bibby:1977:BRBb

[1034] John Bibby. Book review: Multiple Regression and Analysis of Variance: An Introduction for Computer Users in Management and Economics, by George O. Wesolowsky. Journal of the Royal Statistical Society. Series A (General), 140(3):385–386, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344940.

Dyson:1977:BRB

[1035] Tim Dyson. Book review: Demography, by P. R. Cox. Journal of the Royal Statistical Society. Series A (General), 140(3):386–387, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344941.

Armitage:1977:BRB

[1036] P. Armitage. Book review: Mathematical Models in Medicine, by J. Berger; W. Buhler; R. Repges; P. Tautu. Journal of the Royal Statistical Society. Series A (General), 140(3):387, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344942.

Leser:1977:BRB

[1037] C. E. V. Leser. Book review: Entscheidungs- und Spieltheorie, by H. Buhlmann; H. Loeffel; E. Nievergelt. Journal of the Royal Statistical Society. Series A (General), 140(3):387, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344943.

Lesser:1977:BRB

[1038] C. E. V. Lesser. Book review: Theory and Measurement of Consumer Demand, by Henri Theil. Journal of the Royal Statistical Society. Series A (General), 140(3):388, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344944.

Deaton:1977:BRB

[1039] Angus Deaton. Book review: An Annual Growth Model of the U.S. Economy, by B. G. Hickman; R. M. Coen. Journal of the Royal Statistical Society. Series A (General), 140(3):388– 389, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344945.

Sargan:1977:BRBb

[1040] J. D. Sargan. Book review: A Model of Output, Employment, Wages and Prices in the U.K, by I. F. Pearce; P. K. Trivedi; C. T. Stromback; G. J. Anderson. Journal of the Royal Statistical Society. Series A (General), 140 (3):389–390, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344946.

Johnson:1977:BRB

[1041] F. J. Johnson. Book review: Input-Output Techniques: Proceedings of the Second Hungarian Conference on Input-Output Techniques, by A. Csepinszky. Journal of the Royal Statistical Society. Series A (General), 140 (3):390–391, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344947.

Anonymous:1977:RPc

[1042] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 140(3):392–406, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344948.

Anonymous:1977:LAc

[1043] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 140(3):407–409, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344949.

Anonymous:1977:BMc

[1044] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 140(3):i–iv, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344950.

Anonymous:1977:VIb

[1045] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 140(4):572– 574, ???? 1977. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345275.

Anonymous:1977:FMd

[1046] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 140(4):??, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345276.

Campbell:1977:SSW

[1047] M. J. Campbell and A. M. Walker. A survey of statistical work on the Mackenzie River series of annual Canadian lynx trappings for the years 1821– 1934 and a new analysis. Journal of the Royal Statistical Society. Series A (General), 140(4):411–431, ???? 1977. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345277. See comments

[1048].

Tong:1977:SCC

[1048] H. Tong. Some comments on the Canadian lynx data. Journal of the Royal Statistical Society. Series A (General), 140(4):432–436, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345278. See [1047].

Morris:1977:FSC

[1049] M. J. Morris. Forecasting the sunspot cycle. Journal of the Royal Statistical Society. Series A (General), 140(4): 437–468, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345279.

Liddell:1977:MCA

[1050] F. D. K. Liddell, J. C. McDonald, D. C.

Thomas, and Stella V. Cunliffe. Methods of cohort analysis: Appraisal by application to asbestos mining. Journal of the Royal Statistical Society. Series A (General), 140(4):469–491, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345280.

Chatfield:1977:SRD

[1051] C. Chatfield. Some recent developments in time-series analysis. Journal of the Royal Statistical Society. Series A (General), 140(4):492–510, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345281.

Cook:1977:CCR

[1052] W. R. Cook. Curious correlations — a reply. Journal of the Royal Statistical Society. Series A (General), 140 (4):511–513, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345282. See

[873].

Anonymous:1977:PSG

[1053] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 140 (4):514–516, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345283.

Anonymous:1977:POH

[1054] Anonymous. Proceedings of the one hundred and forty-third Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 140 (4):517, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345284.

Anonymous:1977:PR

[1055] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 140(4):518–521, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345285.

Anonymous:1977:HTR

[1056] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 140(4): 522, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345286.

Anonymous:1977:ARC

[1057] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 140 (4):523–544, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345287.

Kingman:1977:BRB

[1058] J. F. C. Kingman. Book review: Problems of Analytical Statistics, by Y. V. Linnik. Journal of the Royal Statistical Society. Series A (General), 140 (4):545, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2345288.

Taylor:1977:BRB

[1059] S. J. Taylor. Book review: Probability Theory with the Essential Analysis, by J. S. Milton; C. P. Tsokos.

Journal of the Royal Statistical Society. Series A (General), 140(4):546, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345289.

Wetherill:1977:BRB

[1060] G. B. Wetherill. Book review: Linear Regression Analysis, by G. A. F. Seber. Journal of the Royal Statistical Society. Series A (General), 140(4): 546–547, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345290.

Champernowne:1977:BRB

[1061] D. G. Champernowne. Book review:

Skew Distributions and the Size of Business Firms, by Y. Ijiri; H. A. Simon. Journal of the Royal Statistical Society. Series A (General), 140(4): 547–548, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345291.

Matatko:1977:BRB

[1062] John Matatko. Book review: Selected Tables in Mathematical Statistics. Volume 3, by Institute of Mathematical Statistics. Journal of the Royal Statistical Society. Series A (General), 140 (4):548–549, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345292.

Tocher:1977:BRB

[1063] K. D. Tocher. Book review: Theory of Modelling and Simulation, by B. P. Zeigler. Journal of the Royal Statistical Society. Series A (General), 140 (4):549–551, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345293.

Haigh:1977:BRBd

[1064] John Haigh. Book review: Stochastic Models in Population Genetics, by Wen-Hsiung Li. Journal of the Royal Statistical Society. Series A (General), 140(4):551–552, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2345294.

Hajnal:1977:BRB

[1065] J. Hajnal. Book review: Demographic Genetics, by Kenneth M. Weiss; Paul A. Ballonoff. Journal of the Royal Statistical Society. Series A (General), 140 (4):552–553, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345295.

Vajda:1977:BRB

[1066] S. Vajda. Book review: Manpower Planning in the Civil Service, by A. R. Smith. Journal of the Royal Statistical Society. Series A (General), 140 (4):553–554, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345296.

Jeffers:1977:BRB

[1067] J. N. R. Jeffers. Book review: Field Experimentation with Fruit Trees and Other Perennial Plants, by S. C. Pearce. Journal of the Royal Statistical Society. Series A (General), 140 (4):554–555, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345297.

Black:1977:BRB

[1068] I. G. Black. Book review: Urban Transport Appraisal, by Ian S. Jones. Journal of the Royal Statistical Society. Series A (General), 140(4):555, ????

1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345298.

Moore:1977:BRBd

[1069] P. G. Moore. Book review: Modern Decision Analysis: Selected Readings, by G. M. Kaufman; H. Thomas. Journal of the Royal Statistical Society. Series A (General), 140(4):555–556, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345299.

Smith:1977:BRBb

[1070] A. F. M. Smith. Book review: Decisions with Multiple Objectives: Preferences and Value Tradeoffs, by R. L. Keeney; H. Raiffa. Journal of the Royal Statistical Society. Series A (General), 140(4):556–557, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345300.

Lindley:1977:BRB

[1071] D. V. Lindley. Book review: Statistical Decision Theory and Related Topics II, by Shanti S. Gupta; David S. Moore. Journal of the Royal Statistical Society. Series A (General), 140 (4):557, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345301.

Jenkins:1977:BRB

[1072] G. M. Jenkins. Book review: Analysis and Control of Dynamic Ecnomic Systems, by Gregory C. Chow. Journal of the Royal Statistical Society. Series A (General), 140(4):557–560, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345302.

Fisk:1977:BRB

[1073] P. R. Fisk. Book review: Foundations of Econometrics, by A. Madansky. Journal of the Royal Statistical Society. Series A (General), 140(4):560, ???? 1977. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2345303.

Topham:1977:BRB

[1074] Neville Topham. Book review: Econometrics of Investment, by J. C. R. Rowley; P. K. Trivedi. Journal of the Royal Statistical Society. Series A (General), 140(4):560–561, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2345304.

Hendry:1977:BRB

[1075] David F. Hendry. Book review: The Models of Project LINK, by J. L. Waelbroeck. Journal of the Royal Statistical Society. Series A (General), 140 (4):561–562, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2345305.

Nicholson:1977:BRBb

[1076] R. J. Nicholson. Book review: Statistical Tables of National Income, Expenditure and Output of the U.K. 1855– 1965, by C. H. Feinstein. Journal of the Royal Statistical Society. Series A (General), 140(4):562, ???? 1977. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345306.

Grimes:1977:BRB

[1077] D. H. Grimes. Book review: Stock Market Logic: A Sophisticated Approach to Profits on Wall Street, by N. G. Fosback. Journal of the Royal Statistical Society. Series A (General), 140(4): 563, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345307.

Norton:1977:ADB

[1078] Bernard Norton. Allan D. Birnbaum 1923–1976. Journal of the Royal Statistical Society. Series A (General), 140 (4):564–565, ???? 1977. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345308.

Witcomb:1977:MJF

[1079] Roger Witcomb. M. J. Farrell, 1926– 1975. Journal of the Royal Statistical Society. Series A (General), 140(4): 565, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345309.

Kendall:1977:MF

[1080] David Kendall. Maurice Frechet, 1878– 1973. Journal of the Royal Statistical Society. Series A (General), 140(4): 566, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345310.

Rudoe:1977:P

[1081] W. Rudoe. T. S. Pilling, 1922–1977.

Journal of the Royal Statistical Society. Series A (General), 140(4):567– 568, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345311.

Barnard:1977:MHQ

[1082] G. A. Barnard. Maurice Henry Quenouille, 1924–1973. Journal of the Royal Statistical Society. Series A (General), 140(4):568–569, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345312.

Hill:1977:DDR

[1083] Austin Bradford Hill. Donald Darnley Reid, 1914–1977. Journal of the Royal Statistical Society. Series A (General), 140(4):569–570, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345313.

Wardrop:1977:RS

[1084] J. G. Wardrop. Reuben Smeed, 1909– 1976. Journal of the Royal Statistical Society. Series A (General), 140(4): 570–571, ???? 1977. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345314.

Anonymous:1977:BMd

[1085] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 140(4):??, ???? 1977. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345315.

Anonymous:1978:FMa

[1086] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 141(1):??, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344773.

Wynn:1978:FSI

[1087] H. P. Wynn. Freedom of statistical information. Journal of the Royal Statistical Society. Series A (General), 141 (1):1–13, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344774.

Tanner:1978:LTF

[1088] J. C. Tanner. Long-term forecasting of vehicle ownership and road traffic. Journal of the Royal Statistical Society. Series A (General), 141(1):14–63, ???? 1978. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344775.

Brooks:1978:NFC

[1089] R. J. Brooks, A. P. Dawid, J. I. Galbriath, R. F. Galbraith, M. Stone, and A. F. M. Smith. A note on forecasting car ownership. Journal of the Royal Statistical Society. Series A (General), 141(1):64–68, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344776.

Watson:1978:VDA

[1090] S. R. Watson and R. V. Brown. The valuation of decision analysis. Journal of the Royal Statistical Society. Series A (General), 141(1):69–78, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344777.

Hull:1978:AMS

[1091] J. C. Hull. The accuracy of the means and standard deviations of subjective probability distributions. Journal of the Royal Statistical Society. Series A (General), 141(1):79–85, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344778.

Vassiliou:1978:HON

[1092] P-C. G. Vassiliou. A high order nonlinear Markovian model for promotion in manpower systems. Journal of the Royal Statistical Society. Series A (General), 141(1):86–94, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344779.

Townsend:1978:SLC

[1093] Joy L. Townsend. Smoking and lung cancer: A cohort data study of men and women in England and Wales 1935–70. Journal of the Royal Statistical Society. Series A (General), 141(1):95–107, ???? 1978. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344780.

Barnett:1978:BRB

[1094] V. Barnett. Book review: Probability and Statistics: Theory and Applications, by D. A. S. Fraser. Journal of the Royal Statistical Society. Series A (General), 141(1):108–109, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344781.

Clarke:1978:BRBa

[1095] G. M. Clarke. Book review: The Advanced Theory of Statistics, by Maurice Kendall; A. Stuart. Journal of the Royal Statistical Society. Series A (General), 141(1):109–110, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344782.

Dawid:1978:BRB

[1096] Philip Dawid. Book review: Statistical Prediction Analysis, by J. Aitchison; I. R. Dunsmore. Journal of the Royal Statistical Society. Series A (General), 141 (1):110–111, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344783.

Marriott:1978:BRBa

[1097] F. H. C. Marriott. Book review: Experiments: Design and Analysis, by J. A. John; M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 141(1):111–112, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344784.

Huitson:1978:BRB

[1098] A. Huitson. Book review: Sampling Inspection in Statistical Quality Control, by W. C. Guenther. Journal of the Royal Statistical Society. Series A (General), 141(1):112, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344785.

Spitz:1978:BRB

[1099] K. E. Spitz. Book review: Case Studies in Sample Design, by A. C. Rosander. Journal of the Royal Statistical Society. Series A (General), 141(1):112– 113, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344786.

Benjamin:1978:BRBa

[1100] B. Benjamin. Book review: Statistics and Public Policy, by William B. Fairley; Frederick Mosteller. Journal of the Royal Statistical Society. Series A (General), 141(1):113, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344787.

Chatfield:1978:BRBa

[1101] C. Chatfield. Book review: Forecasting Methods for Management, by S.

C. Wheelwright; S. Makridakis. Journal of the Royal Statistical Society. Series A (General), 141(1):113–114, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344788.

Oliver:1978:BRBa

[1102] F. R. Oliver. Book review: Introduction to Probability and Statistics, by Henry L. Alder; Edward B. Roessler. Journal of the Royal Statistical Society. Series A (General), 141(1):114–115, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344789.

Dalton:1978:BRB

[1103] E. Rosemary Dalton. Book review: Introductory Medical Statistics, by R. F. Mould. Journal of the Royal Statistical Society. Series A (General), 141 (1):115, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344790.

Keen:1978:BRB

[1104] J. Keen. Book review: Descriptive Statistics: A Contemporary Approach, by R. P. Runyon. Journal of the Royal Statistical Society. Series A (General), 141(1):115–116, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344791.

Hills:1978:BRB

[1105] M. Hills. Book review: Winning with Statistics, by R. P. Runyon. Journal of the Royal Statistical Society. Series A (General), 141(1):117, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344792.

Wilkinson:1978:BRB

[1106] R. K. Wilkinson. Book review: Projects in Economic and Social Statistics, by David G. Mayes. Journal of the Royal Statistical Society. Series A (General), 141(1):117–118, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344793.

Whittle:1978:BRB

[1107] P. Whittle. Book review: Latent Variables in Socio-Economic Models, by D. J. Aigner; A. S. Goldberger. Journal of the Royal Statistical Society. Series A (General), 141(1):118, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344794.

Benjamin:1978:BRBb

[1108] B. Benjamin. Book review: The Natural History of Chronic Bronchitis and Emphysema, by Charles Fletcher; Richard Peto; Cecily Tinker; Frank E. Spiezer. Journal of the Royal Statistical Society. Series A (General), 141 (1):119, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344795.

Benjamin:1978:BRBc

[1109] B. Benjamin. Book review: Seminars in Community Medicine, Vol. 2 — Health Information, Planning and Monitoring, by Roy M. Acheson; David J. Hall; Lesley Aird. Journal of the Royal Statistical Society. Series A (General), 141(1):119–120, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344796.

Stone:1978:BRBa

[1110] M. Stone. Book review: Statistics of Smoking in the United Kingdom, by P. N. Lee. Journal of the Royal Statistical Society. Series A (General), 141 (1):120–121, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344797.

Barton:1978:BRB

[1111] D. E. Barton. Book review: Studies in the History of Statistics and Probability. Vol. II, by M. G. Kendall; R. L. Plackett. Journal of the Royal Statistical Society. Series A (General), 141 (1):121, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344798.

Stone:1978:BRBb

[1112] M. Stone. Book review: Prediction Analysis of Cross-Classifications, by David K. Hildebrand; James D. Laing; Howard Rosenthal. Journal of the Royal Statistical Society. Series A (General), 141(1):122, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344799.

Binmore:1978:BRB

[1113] K. G. Binmore. Book review: Mathematical Economics and Game Theory — Essays in Honor of Oskar Morgenstern, by R. Henn; O. Moeschlin. Journal of the Royal Statistical Society. Series A (General), 141(1):122–123, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344800.

Anonymous:1978:RPa

[1114] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 141(1):124–152, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344801.

Anonymous:1978:BMa

[1115] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 141(1):??, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344802.

Anonymous:1978:FMb

[1116] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 141(2):??, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344451.

Duncan:1978:ORH

[1117] I. B. Duncan and R. N. Curnow. Operational research in the health and social services. Journal of the Royal Statistical Society. Series A (General), 141 (2):153–194, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344452.

Aitkin:1978:AUC

[1118] Murray Aitkin. The analysis of unbalanced cross-classifications. Journal of the Royal Statistical Society. Series A (General), 141(2):195–223, ???? 1978. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344453. See corrigenda

Gail:1978:AHI

[1119] Mitchell Gail. The analysis of heterogeneity for indirect standardized mortality ratios. Journal of the Royal Statistical Society. Series A (General), 141 (2):224–234, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344454.

Edmonston:1978:INE

[1120] Barry Edmonston and Omar Davies. Interpreting the negative exponential density gradient. Journal of the Royal Statistical Society. Series A (General), 141(2):235–241, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344455.

Woolson:1978:AID

[1121] R. F. Woolson, J. D. Leeper, and W. R. Clarke. Analysis of incomplete data from longitudinal and mixed longitudinal studies. Journal of the Royal Statistical Society. Series A (General), 141 (2):242–252, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344456.

Thom:1978:DBS

[1122] A. Thom. The distances between stones in stone rows. Journal of the Royal Statistical Society. Series A (General), 141(2):253–257, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344457.

Hajnal:1978:BRB

[1123] J. Hajnal. Book review: Perspectives in Probability and Statistics. Papers in Honour of M. S. Bartlett, by J. Gani.

Journal of the Royal Statistical Society. Series A (General), 141(2):258– 259, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344458.

Kanji:1978:BRB

[1124] G. K. Kanji. Book review: Probability and Statistics for Engineers and Scientists, by R. Walpole; R. Myers. Journal of the Royal Statistical Society. Series A (General), 141(2):259–260, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344459.

Folks:1978:BRB

[1125] J. Leroy Folks. Book review: Parameter Estimation in Engineering and Science, by James V. Beck; Kenneth J. Arnold. Journal of the Royal Statistical Society. Series A (General), 141 (2):260, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344460.

Evans:1978:BRB

[1126] I. G. Evans. Book review: Statistical Analysis for Decision Making, by Morris Hamburg. Journal of the Royal Statistical Society. Series A (General), 141 (2):260–261, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344461.

Jackson:1978:BRB

[1127] P. H. Jackson. Book review: Elements of Statistical Inference, by D. Huntsberger; P. Billingsley. Journal of the Royal Statistical Society. Series A (General), 141(2):261–262, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344462.

Dunsmore:1978:BRB

[1128] I. R. Dunsmore. Book review: Statistical Concepts and Methods, by G. K. Bhattacharyya; R. A. Johnson. Journal of the Royal Statistical Society. Series A (General), 141(2):262, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344463.

Ray:1978:BRBa

[1129] W. D. Ray. Book review: Probability and Statistical Inference, by Robert V. Hogg; Elliot A. Tanis. Journal of the Royal Statistical Society. Series A (General), 141(2):262–263, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344464.

Silvey:1978:BRB

[1130] S. D. Silvey. Book review: Probability and Statistics, by A. M. Mathai; P. N. Rathie. Journal of the Royal Statistical Society. Series A (General), 141 (2):263, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344465.

Holmes:1978:BRB

[1131] P. Holmes. Book review: Elementary Probability and Statistical Reasoning, by H. E. Reinhardt; D. O. Loftsgaarden. Journal of the Royal Statistical Society. Series A (General), 141 (2):263–264, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344466.

Haigh:1978:BRBa

[1132] John Haigh. Book review: Nonparametric Statistics: A Contemporary Approach, by Richard P. Runyon. Journal of the Royal Statistical Society. Series A (General), 141(2):264, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344467.

Marriott:1978:BRBb

[1133] F. H. C. Marriott. Book review: Pocket Book of Statistical Tables, by R. E. Odeh; D. B. Owen; Z. W. Birnbaum; L. Fisher. Journal of the Royal Statistical Society. Series A (General), 141 (2):264–265, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344468.

Munford:1978:BRB

[1134] A. G. Munford. Book review: Urn Models and their Application: An Approach to Modern Discrete Probability Theory, by Norman L. Johnson; Samuel Kotz. Journal of the Royal Statistical Society. Series A (General), 141 (2):265, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344469.

Gower:1978:BRBa

[1135] J. C. Gower. Book review: Methods for Statistical Data Analysis of Multivariate Observations, by R. Gnanadesikan. Journal of the Royal Statistical Society. Series A (General), 141(2): 265–267, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344470.

Ross:1978:BRB

[1136] G. J. S. Ross. Book review: Multivariate Statistical Methods: Within Groups Covariation, by E. H. Bryant; W. R. Atcherley; Multivariate Statistical Methods: Among Groups Covariation, by W. R. Atcherley; E. H. Bryant. Journal of the Royal Statistical Society. Series A (General), 141(2):267– 268, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344471.

Ray:1978:BRBb

[1137] W. D. Ray. Book review: Multivariate Analysis IV, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 141(2):268, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344472.

Ray:1978:BRBc

[1138] W. D. Ray. Book review: Maximum Likelihood Estimation in Small Samples, by L. R. Shenton; K. O. Bowman. Journal of the Royal Statistical Society. Series A (General), 141(2):268, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344473.

Chatfield:1978:BRBb

[1139] C. Chatfield. Book review: Time Series, by Maurice Kendall. Journal of the Royal Statistical Society. Series A (General), 141(2):269, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344474.

Kemp:1978:BRB

[1140] A. W. Kemp. Book review: Decomposition of Superpositions of Density Functions and Discrete Distributions, by Pal Medgyessy. Journal of the Royal Statistical Society. Series A (General), 141 (2):269–270, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344475.

Holgate:1978:BRBa

[1141] P. Holgate. Book review: Selection in One- and Two-Locus Systems, by T. Nagylaki. Journal of the Royal Statistical Society. Series A (General), 141 (2):270–271, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344476.

Eversley:1978:BRB

[1142] D. E. C. Eversley. Book review:

Zero Growth?, by A. Sauvy. Journal of the Royal Statistical Society. Series A (General), 141(2):271–272, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344477.

Anonymous:1978:RPb

[1143] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 141(2):273–285, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344478.

Herzberg:1978:CRW

[1144] A. M. Herzberg and D. R. Cox. Correction: Recent work on the design of experiments: A bibliography and a review. Journal of the Royal Statistical Society. Series A (General), 141(2):

286, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344479. See [1].

Anonymous:1978:BMb

[1145] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 141(2):i–ii, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344480.

Anonymous:1978:FMc

[1146] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 141(3):??, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344803.

Stigler:1978:FYE

[1147] Stephen M. Stigler. Francis Ysidro Edgeworth, statistician. Journal of the Royal Statistical Society. Series A (General), 141(3):287–322, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344804.

Gani:1978:SPE

[1148] J. Gani. Some problems of epidemic theory. Journal of the Royal Statistical Society. Series A (General), 141 (3):323–347, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344805.

Forsyth:1978:PCC

[1149] F. G. Forsyth. The practical construction of a chain price index number. Journal of the Royal Statistical Society. Series A (General), 141(3):348– 358, ???? 1978. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2344806.

Byron:1978:ELS

[1150] R. P. Byron. The estimation of large social account matrices. Journal of the Royal Statistical Society. Series A (General), 141(3):359–367, ???? 1978. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344807. See corrigenda

[1304].

Prais:1978:SPL

[1151] S. J. Prais. The strike-proneness of large plants in Britain. Journal of the Royal Statistical Society. Series A (General), 141(3):368–384, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344808.

Healy:1978:SS

[1152] M. J. R. Healy. Is statistics a science? Journal of the Royal Statistical Society. Series A (General), 141 (3):385–393, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344809.

Uche:1978:SAR

[1153] P. I. Uche. Statistical analysis of recurrent expenditure of universities in a developing country: A Nigerian example. Journal of the Royal Statistical Society. Series A (General), 141 (3):394–400, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344810.

Flaherty:1978:RBC

[1154] David H. Flaherty. Report of the Bellagio Conference. Journal of the Royal Statistical Society. Series A (General), 141(3):401–405, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344811.

Brillinger:1978:HS

[1155] David R. Brillinger. Henry Scheffe, 1907–1977. Journal of the Royal Statistical Society. Series A (General), 141 (3):406–407, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344812.

Campion:1978:HM

[1156] H. Campion. Herbert Marshall, 1887– 1977. Journal of the Royal Statistical Society. Series A (General), 141(3): 407–408, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344813.

Hill:1978:JCG

[1157] I. D. Hill. John Cleland Gascoigne, 1913–1978. Journal of the Royal Statistical Society. Series A (General), 141 (3):409, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344814.

Gower:1978:BRBb

[1158] J. C. Gower. Book review: Multivariate Statistical Inference, by Narayan C. Giri. Journal of the Royal Statistical Society. Series A (General), 141 (3):410–411, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344815.

Marriott:1978:BRBc

[1159] F. H. C. Marriott. Book review: BMDP-77. Biomedical Computer Programs, P -Series, by W. J. Dixon; M. B. Brown; L. Engleman; J. W. Frame; R. I. Jennrich. Journal of the Royal Statistical Society. Series A (General), 141(3):411, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344816.

Hill:1978:BRB

[1160] I. D. Hill. Book review: Programming and Algorithms, by A. J. Guttmann. Journal of the Royal Statistical Society. Series A (General), 141(3):411– 412, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344817.

Clarke:1978:BRBb

[1161] G. M. Clarke. Book review: Statistical Treatment of Experimental Data, by J. R. Green; D. Margerison. Journal of the Royal Statistical Society. Series A (General), 141(3):412–413, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344818.

Marriott:1978:BRBd

[1162] F. H. C. Marriott. Book review: Statistics: A Biomedical Introduction, by B. W. Brown; M. Hollander. Journal of the Royal Statistical Society. Series A (General), 141(3):413, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344819.

Mardia:1978:BRB

[1163] K. V. Mardia. Book review: Empirical Distributions and Rank Statistics, by M. C. A. Van Zuijlen. Journal of the Royal Statistical Society. Series A (General), 141(3):413–414, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344820.

Daley:1978:BRB

[1164] D. J. Daley. Book review: The General Point Process: Applications to Structural Fatigue, Bioscience, and Medical Research, by V. K. Murthy. Journal of the Royal Statistical Society. Series A (General), 141(3):414, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344821.

Freeman:1978:BRB

[1165] P. R. Freeman. Book review: Contracting Markov Decision Processes, by J. A. E. E. van Nunen. Journal of the Royal Statistical Society. Series A (General), 141(3):414, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344822.

Holgate:1978:BRBb

[1166] P. Holgate. Book review: Measuring Selection in Natural Populations, by F. B. Christiansen; T. M. Fenchel. Journal of the Royal Statistical Society. Series A (General), 141(3):414–415, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344823.

Ray:1978:BRBd

[1167] W. D. Ray. Book review: Stochastic Models for Spike Trains of Single Neurons, by G. Sampath; S. K. Srinivasan. Journal of the Royal Statistical Society. Series A (General), 141(3):415, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344824.

Hammersley:1978:BRB

[1168] J. M. Hammersley. Book review: Diffusion Processes and Related Topics in Biology, by L. M. Ricciardi. Journal of the Royal Statistical Society. Series A (General), 141(3):415–416, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344825.

Chatfield:1978:BRBc

[1169] C. Chatfield. Book review: Population Projections 1975–2015, by Government Actuary. Journal of the Royal Statistical Society. Series A (General), 141 (3):416, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344826.

Stone:1978:BRBc

[1170] M. Stone. Book review: Applications of Statistics: Proceedings of the Symposium Held at Wright State University, June 1976, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 141(3):417, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344827.

Farebrother:1978:BRBa

[1171] R. W. Farebrother. Book review: Basic Econometrics, by Michael S. Common. Journal of the Royal Statistical Society. Series A (General), 141(3): 417–418, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344828.

Allen:1978:BRB

[1172] R. G. D. Allen. Book review: The Price Index, by S. N. Afriat. Journal of the Royal Statistical Society.

Series A (General), 141(3):418, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344829.

Carruthers:1978:BRB

[1173] A. G. Carruthers. Book review: British Employment Statistics, by N. K. Buxton; D. I. MacKay. Journal of the Royal Statistical Society. Series A (General), 141(3):418–420, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344830.

Oliver:1978:BRBb

[1174] F. R. Oliver. Book review: Structure, System and Economic Policy, by Wassily Leontief. Journal of the Royal Statistical Society. Series A (General), 141 (3):420–421, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344831.

Armstrong:1978:BRB

[1175] A. G. Armstrong. Book review: Medium-Term Dynamic Forecasting: Proceedings of the 1975 London Input-Output Conference, by W. F. Gossling. Journal of the Royal Statistical Society. Series A (General), 141(3):421, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344832.

Nicholson:1978:BRBa

[1176] R. J. Nicholson. Book review: Measuring Inequality, by F. A. Cowell. Journal of the Royal Statistical Society. Series A (General), 141(3):421–422, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344833.

Hall:1978:BRB

[1177] David J. Hall. Book review: National Health Survey Systems in the European Economic Community, by P. Armitage. Journal of the Royal Statistical Society. Series A (General), 141(3):422– 423, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344834.

Anonymous:1978:RPc

[1178] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 141(3):424–436, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344835.

Anonymous:1978:BMc

[1179] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 141(3):i–vi, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344836.

Anonymous:1978:VI

[1180] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 141(4):586– 589, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344481.

Anonymous:1978:FMd

[1181] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 141(4):??, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344482.

Burch:1978:SLC

[1182] P. R. J. Burch. Smoking and lung cancer: The problem of inferring cause. Journal of the Royal Statistical Society. Series A (General), 141(4):437– 477, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344483.

Fessey:1978:BSD

[1183] M. C. Fessey. Business statistics: A decade of central collection. Journal of the Royal Statistical Society. Series A (General), 141(4):478–506, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344484.

Upton:1978:NEV

[1184] Graham J. G. Upton. A note on the estimation of voter transition probabilities. Journal of the Royal Statistical Society. Series A (General), 141 (4):507–512, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344485.

Kemp:1978:CGB

[1185] C. D. Kemp and S. Loukas. The computer generation of bivariate discrete random variables. Journal of the Royal Statistical Society. Series A (General), 141(4):513–519, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344486.

Anonymous:1978:PSG

[1186] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 141 (4):520, ???? 1978. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2344487.

Anonymous:1978:POH

[1187] Anonymous. Proceedings of the one hundred and forty-fourth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 141 (4):520, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344488.

Anonymous:1978:PR

[1188] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 141(4):521–522, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344489.

Anonymous:1978:HTR

[1189] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 141(4): 523, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344490.

Anonymous:1978:ARC

[1190] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 141 (4):524–547, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344491.

Osborn:1978:BRBa

[1191] J. F. Osborn. Book review: Exploratory Data Analysis, by John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 141(4):548– 549, ???? 1978. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2344492.

Osborn:1978:BRBb

[1192] J. F. Osborn. Book review: Data Analysis and Regression, by F. Mosteller; John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 141(4):549–550, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344493.

Nelder:1978:BRB

[1193] J. A. Nelder. Book review: Computational Methods for Data Analysis, by John M. Chambers. Journal of the Royal Statistical Society. Series A (General), 141(4):550–551, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344494.

Holt:1978:BRB

[1194] D. Holt. Book review: The Analysis of Cross-Classified Categorical Data, by Stephen E. Fienberg. Journal of the Royal Statistical Society. Series A (General), 141(4):551, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344495.

Ray:1978:BRBe

[1195] W. D. Ray. Book review: Stochastic Processes, by John Lamperti. Journal of the Royal Statistical Society. Series A (General), 141(4):551–552, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344496.

Marriott:1978:BRBe

[1196] F. H. C. Marriott. Book review: Models of Spatial Processes, by A. Getis; B. Boots. Journal of the Royal Statistical Society. Series A (General), 141 (4):552, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344497.

Abakuks:1978:BRB

[1197] A. Abakuks. Book review: Integrodifferential Equations and Delay Models in Population Dynamics, by J. M. Cushing. Journal of the Royal Statistical Society. Series A (General), 141 (4):552–553, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344498.

Haigh:1978:BRBb

[1198] John Haigh. Book review: Stochastic Problems in Population Genetics, by Takeo Maruyama. Journal of the Royal Statistical Society. Series A (General), 141(4):553–554, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344499.

Scott:1978:BRB

[1199] Chris Scott. Book review: Survey Research Practice, by Gerald Hoinville; Roger Jowell. Journal of the Royal Statistical Society. Series A (General), 141 (4):554, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344500.

Wilson:1978:BRB

[1200] A. G. Wilson. Book review: Locational Analysis in Human Geography: Volume. 1: Locational Models, by P. Haggett; A. D. Cliff; A. Frey; Locational Analysis in Human Geography: Volume. 2: Locational Methods, by P. Haggett; A. D. Cliff; A. Frey. Journal of the Royal Statistical Society. Series A (General), 141(4):554–555, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344501.

Farebrother:1978:BRBb

[1201] R. W. Farebrother. Book review: The Statistical Implications of Pre-Test and Stein-Rule Estimators in Econometrics, by George G. Judge; M. E. Bock. Journal of the Royal Statistical Society. Series A (General), 141(4):555– 556, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344502.

Kmietowicz:1978:BRB

[1202] Z. W. Kmietowicz. Book review: Statistics for Modern Business Decisions, by Lawrence Lapin. Journal of the Royal Statistical Society. Series A (General), 141(4):556–557, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344503.

Bailey:1978:BRB

[1203] B. J. R. Bailey. Book review: Selected Tables in Mathematical Statistics. Volume 5, by Institute of Mathematical Statistics. Journal of the Royal Statistical Society. Series A (General), 141 (4):557–558, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344504.

Smith:1978:BRBa

[1204] A. F. M. Smith. Book review: Conflicting Objectives in Decisions, by David E. Bell; Ralph L. Keeney; Howard Raiffa. Journal of the Royal Statistical Society. Series A (General), 141 (4):558–559, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344505.

Smith:1978:BRBb

[1205] A. F. M. Smith. Book review: New Developments in the Applications of Bayesian Methods, by A. Aykac; C. Brumat. Journal of the Royal Statistical Society. Series A (General), 141 (4):559, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344506.

Liddell:1978:BRB

[1206] F. D. K. Liddell. Book review: Statistics and the Environment, a Forum for Interdisciplinary Interaction, by Richard H. Foote. Journal of the Royal Statistical Society. Series A (General), 141(4):559–561, ???? 1978. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344507.

Archer:1978:BRB

[1207] Catherine A. Archer. Book review:

Theory and Applications of Economic Indices, by W. Eichhorn. Journal of the Royal Statistical Society. Series A (General), 141(4):561, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344508.

Nicholson:1978:BRBb

[1208] R. J. Nicholson. Book review: Regional Accounts, by Central Statistical Office. Journal of the Royal Statistical Society. Series A (General), 141(4):562, ???? 1978. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344509.

Nicholson:1978:BRBc

[1209] R. J. Nicholson. Book review: Distribution of Personal Wealth in Britain, by A. B. Atkinson; A. J. Harrison. Journal of the Royal Statistical Society. Series A (General), 141(4):562–563, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344510.

Conolly:1978:BRB

[1210] B. W. Conolly. Book review: Handbook of Hypergeometric Integrals, by Harold Exton. Journal of the Royal Statistical Society. Series A (General), 141 (4):563–564, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344511.

Anonymous:1978:RPd

[1211] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 141(4):565–585, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344512.

Anonymous:1978:BMd

[1212] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 141(4):i–ii, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344513.

Anonymous:1979:FMa

[1213] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 142(1):??, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344650.

Wetherill:1979:SIS

[1214] G. Barrie Wetherill and Julian Kollerstrom. Sampling inspection simplified. Journal of the Royal Statistical Society. Series A (General), 142(1):1–32, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344651. See corrigenda

[1303].

Holt:1979:PS

[1215] D. Holt and T. M. F. Smith. Post stratification. Journal of the Royal Statistical Society. Series A (General), 142 (1):33–46, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344652.

Clark:1979:CCV

[1216] R. M. Clark. Calibration, crossvalidation and carbon-14. I. Journal of the Royal Statistical Society. Series A (General), 142(1):47–62, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344653.

Hill:1979:BRB

[1217] I. D. Hill. Book review: The Mathematical Work of Charles Babbage, by J. M. Dubbey. Journal of the Royal Statistical Society. Series A (General), 142(1):63–64, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344654.

Healy:1979:BRBa

[1218] M. J. R. Healy. Book review: Outliers in Statistical Data, by V. Barnett; T. Lewis. Journal of the Royal Statistical Society. Series A (General), 142(1):64– 65, ???? 1979. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344655.

Law:1979:BRB

[1219] I. A. Law. Book review: A Basic Course in Statistics, by G. M. Clarke; D. Cooke. Journal of the Royal Statistical Society. Series A (General), 142 (1):65–66, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344656.

Neal:1979:BRB

[1220] F. Neal. Book review: Introductory Statistics, by T. H. Wonnacott; R. J. Wonnacott. Journal of the Royal Statistical Society. Series A (General), 142 (1):66–67, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344657.

Kingman:1979:BRBa

[1221] J. F. C. Kingman. Book review: Information and Exponential Families in Statistical Theory, by O. Barndorff-Nielsen. Journal of the Royal Statistical Society. Series A (General), 142(1):67, ???? 1979. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344658.

Bather:1979:BRB

[1222] J. A. Bather. Book review: Markov Decision Theory, by H. C. Tijms; J. Wessels. Journal of the Royal Statistical Society. Series A (General), 142(1):67, ???? 1979. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344659.

Smith:1979:BRB

[1223] A. F. M. Smith. Book review: Prediction and Improved Estimation in Linear Models, by J. Bibby; H. Toutenburg. Journal of the Royal Statistical Society. Series A (General), 142(1):68– 69, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344660.

Kemp:1979:BRB

[1224] A. W. Kemp. Book review: Characterizations of the Normal Probability Law, by A. M. Mathai; G. Pederzoli. Journal of the Royal Statistical Society. Series A (General), 142(1):69–70, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344661.

Ray:1979:BRBa

[1225] W. D. Ray. Book review: Forecasting and Planning, by R. Fildes; D. Wood. Journal of the Royal Statistical Society. Series A (General), 142(1):70, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344662.

Burman:1979:BRB

[1226] J. P. Burman. Book review: Interactive Forecasting: Univariate and Multivariate Methods, by S. Makridakis; S. C. Wheelwright. Journal of the Royal Statistical Society. Series A (General), 142(1):71, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344663.

Ray:1979:BRBb

[1227] W. D. Ray. Book review: Managerial Planning: An Optimum and Stochastic Control Approach. (Volumes 1 and 2), by Charles S. Tapiero. Journal of the Royal Statistical Society. Series A (General), 142(1):71–72, ????

1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344664.

Moore:1979:BRB

[1228] P. G. Moore. Book review: Statistical Analysis for Decision Making, by Morris Hanbury. Journal of the Royal Statistical Society. Series A (General), 142 (1):72, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344665.

Rowe:1979:BRB

[1229] B. C. Rowe. Book review: Computer Models in the Social Sciences, by R. B. Coats; A. Parkin. Journal of the Royal Statistical Society. Series A (General), 142(1):73, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344666.

Vajda:1979:BRBa

[1230] S. Vajda. Book review: Model Building in Mathematical Programming, by H. P. Williams. Journal of the Royal Statistical Society. Series A (General), 142(1):73–74, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344667.

Barbour:1979:BRB

[1231] A. D. Barbour. Book review: Network Models in Population Biology, by E. R. Lewis. Journal of the Royal Statistical Society. Series A (General), 142(1):74– 75, ???? 1979. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344668.

Benjamin:1979:BRB

[1232] Bernard Benjamin. Book review: Occupational Mortality–Decennial Supplement to the Registrar General’s Statistical Review for England and Wales 1970–72. Journal of the Royal Statistical Society. Series A (General), 142 (1):75–76, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344669.

Cox:1979:BRB

[1233] P. R. Cox. Book review: Mathematical Demography, by D. Smith; N. Keyfitz. Journal of the Royal Statistical Society. Series A (General), 142(1):77, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344670.

Buckland:1979:BRB

[1234] W. R. Buckland. Book review: Material Accountability: Theory, Verification, Applications, by R. Avenhaus. Journal of the Royal Statistical Society. Series A (General), 142(1):77–78, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344671.

Webb:1979:BRB

[1235] Michael G. Webb. Book review: Industrial Pricing in the U.K, by K. Coutts; W. Godley; W. Nordhaus. Journal of the Royal Statistical Society. Series A (General), 142(1):78–79, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344672.

Archer:1979:BRB

[1236] Catherine A. Archer. Book review: On the Factorial Approach Providing the True Index of Cost of Living, by Kali S. Banerjee. Journal of the Royal Statistical Society. Series A (General), 142(1):

79, ???? 1979. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344673.

Stone:1979:JM

[1237] Richard Stone. Jacob Marschak, 1898– 1977. Journal of the Royal Statistical Society. Series A (General), 142(1):80– 81, ???? 1979. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344674.

Anonymous:1979:RPa

[1238] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 142(1):82–96, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344675.

Anonymous:1979:BMa

[1239] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 142(1):i–ii, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344676.

Anonymous:1979:FMb

[1240] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 142(2):??, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2345076.

Makridakis:1979:AFE

[1241] Spyros Makridakis, Michele Hibon, and Claus Moser. Accuracy of forecasting: An empirical investigation. Journal of the Royal Statistical Society. Series A (General), 142(2):97–145, ????

1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345077.

Lindley:1979:RPA

[1242] D. V. Lindley, A. Tversky, and R. V.

Brown. On the reconciliation of probability assessments. Journal of the Royal Statistical Society. Series A (General), 142(2):146–180, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345078.

Fryer:1979:CEM

[1243] J. G. Fryer, R. A. Harding, M. D.

Macdonald, K. L. Q. Read, and G. R. Crocker. Comparing the early mortality rates of the local authorities in England and Wales. Journal of the Royal Statistical Society. Series A (General), 142(2):181–198, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345079.

Bhansali:1979:MSA

[1244] R. J. Bhansali. A mixed spectrum analysis of the Lynx data. Journal of the Royal Statistical Society. Series A (General), 142(2):199–209, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345080.

Kalton:1979:UCS

[1245] Graham Kalton. Ultimate cluster sampling. Journal of the Royal Statistical Society. Series A (General), 142(2): 210–222, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345081.

Rees:1979:RPP

[1246] Philip H. Rees. Regional population project models and accounting methods. Journal of the Royal Statistical Society. Series A (General), 142(2):223– 255, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345082.

Burrows:1979:SPP

[1247] Peter M. Burrows. Selected percentage points of Greenwood’s statistics. Journal of the Royal Statistical Society. Series A (General), 142(2):256–258, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345083.

Kendall:1979:BRBa

[1248] David Kendall. Book review: I. J. Bienayme: Statistical Theory Anticipated, by C. C. Heyde; E. Seneta. Journal of the Royal Statistical Society. Series A (General), 142(2):259–260, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345084.

Silvey:1979:BRBa

[1249] S. D. Silvey. Book review: Problems and Solutions in Theoretical Statistics, by D. R. Cox; D. V. Hinkley. Journal of the Royal Statistical Society. Series A (General), 142(2):260, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345085.

Richardson:1979:BRB

[1250] S. C. Richardson. Book review: Graphical Techniques for Multivariate Data, by Brian Everitt. Journal of the Royal Statistical Society. Series A (General), 142(2):261, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345086.

Oliver:1979:BRB

[1251] F. R. Oliver. Book review: Essentials of Statistics, by Stephen A. Book. Journal of the Royal Statistical Society. Series A (General), 142(2):261– 262, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345087.

Kmietowicz:1979:BRB

[1252] Z. W. Kmietowicz. Book review:

Statistics in Theory and Practice (7th Edition), by L. R. Connor; A. J. H. Morrell. Journal of the Royal Statistical Society. Series A (General), 142 (2):262–263, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345088.

Kingman:1979:BRBb

[1253] J. F. C. Kingman. Book review: Infinitely Divisible Point Processes, by J. Keersten; K. Matthes; J. Mecke. Journal of the Royal Statistical Society. Series A (General), 142(2):263, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345089.

Holgate:1979:BRBa

[1254] P. Holgate. Book review: Random Allocation, by V. F. Kolchin; B. A. Sevast’yanov; V. P. Chistyakov; A. V. Balakrishnan. Journal of the Royal Statistical Society. Series A (General), 142 (2):263–264, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345090.

Wood:1979:BRB

[1255] Jean Wood. Book review: Contributions to Survey Sampling and Applied Statistics. Papers in Honor of H.

O. Hartley, by H. A. David. Journal of the Royal Statistical Society. Series A (General), 142(2):264–265, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345091.

Kalton:1979:BRB

[1256] Graham Kalton. Book review: Statistical Survey Techniques, by R. J. Jessen. Journal of the Royal Statistical Society. Series A (General), 142(2):265, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345092.

Wilkinson:1979:BRBa

[1257] R. K. Wilkinson. Book review: Introduction to Quantitative Economics, by Brian Haines. Journal of the Royal Statistical Society. Series A (General), 142 (2):265–266, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345093.

Hornstein:1979:BRB

[1258] Zmira Hornstein. Book review: Correlational Procedures for Research, by Robert M. Thorndike. Journal of the Royal Statistical Society. Series A (General), 142(2):266, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345094.

Silvey:1979:BRBb

[1259] S. D. Silvey. Book review: Asymptotic Optimality of Likelihood Ratio Tests in Exponential Families, by W. C. M. Kallenberg. Journal of the Royal Statistical Society. Series A (General), 142 (2):266–267, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345095.

Thomas:1979:BRBa

[1260] J. J. Thomas. Book review: Disturbances in the Linear Model, Estimation and Hypothesis Testing, by C. Dubbleman. Journal of the Royal Statistical Society. Series A (General), 142(2): 267, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345096.

Olkin:1979:BRB

[1261] Ingram Olkin. Book review: Selected Tables in Mathematical Statistics, Volume 4. Dirichlet Distribution Type 1, by M. Sobel; V. R. R. Uppuluri; K. Frankowski. Journal of the Royal Statistical Society. Series A (General), 142 (2):268, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345097.

Vajda:1979:BRBb

[1262] S. Vajda. Book review: Numerical Methods for Unconstrained Optimization: An Introduction, by M. A. Wolfe. Journal of the Royal Statistical Society. Series A (General), 142(2):268– 269, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2345098.

Kendall:1979:BRBb

[1263] David Kendall. Book review: The Measurement of Biological Shape and Change, by F. L. Bookstein. Journal of the Royal Statistical Society. Series A (General), 142(2):269, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345099.

Bartholomew:1979:BRBa

[1264] D. J. Bartholomew. Book review: Management Science Approaches to Manpower Planning and Organization Design, by A. Charnes; W. W. Cooper; R. J. Niehaus. Journal of the Royal Statistical Society. Series A (General), 142(2):269–270, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2345100.

Simpson:1979:BRB

[1265] I. G. Simpson. Book review: Econometric Simulation Model of the U.S. Farm Sector and its Policies and Food Exports, by Earl O. Heady; Thomas M. Reynolds; Donald O. Mitchell. Journal of the Royal Statistical Society. Series A (General), 142(2):270–271, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345101.

Leser:1979:BRB

[1266] C. E. V. Leser. Book review: The Distribution of Economic Well-Being, by F. Thomas Juster. Journal of the Royal Statistical Society. Series A (General), 142(2):271, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345102.

Harris:1979:BRB

[1267] A. H. Harris. Book review: Resources, Environment and Economics: Applications of the Materials/ Energy Balance Principle, by Robert U. Ayres. Journal of the Royal Statistical Society. Series A (General), 142(2):271–272, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345103.

Anonymous:1979:RPb

[1268] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 142(2):273–284, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345104.

Anonymous:1979:BMb

[1269] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 142(2):i–viii, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2345105.

Anonymous:1979:FMc

[1270] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 142(3):??, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982481.

Dalenius:1979:DPL

[1271] Tore Dalenius. Data protection legislation in Sweden: A Statistician’s perspective. Journal of the Royal Statistical Society. Series A (General), 142 (3):285–298, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982482. See

discussion [1274].

Durbin:1979:SRD

[1272] J. Durbin. Statistics and the report of the Data Protection Committee. Journal of the Royal Statistical Society. Series A (General), 142(3):299– 306, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982483. See dis-

cussion [1274].

Anonymous:1979:SRA

[1273] Anonymous. Statistics, research and archives. Journal of the Royal Statistical Society. Series A (General), 142 (3):307–316, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982484.

Anonymous:1979:DPP

[1274] Anonymous. Discussion of the papers by Professor Dalenius and by Professor Durbin. Journal of the Royal Statistical Society. Series A (General), 142 (3):316–329, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982485. See

[1271, 1272].

Gleed:1979:DRC

[1275] R. H. Gleed and R. D. Rees. The derivation of regional capital stock estimates for UK manufacturing industries. 1951–1973. Journal of the Royal Statistical Society. Series A (General), 142(3):330–346, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982486.

Fine:1979:JBM

[1276] Paul E. M. Fine. John Brownlee and the measurement of infectiousness: An historical study in epidemic theory. Journal of the Royal Statistical Society. Series A (General), 142(3):347– 362, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982487.

Chatfield:1979:IA

[1277] C. Chatfield. Inverse autocorrelations. Journal of the Royal Statistical Society. Series A (General), 142(3):363– 377, ???? 1979. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2982488.

Hill:1979:ADG

[1278] I. D. Hill. Approximating the distribution of Greenwood’s statistic with Johnson distributions. Journal of the Royal Statistical Society. Series A (General), 142(3):378–380, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982489. See corrigenda

[1332].

Ray:1979:BRBc

[1279] W. D. Ray. Book review: Statistics for Experiments. An Introduction to Design, Data Analysis and Model Building, by George E. P. Box; William G. Hunter; J. Stuart Hunter. Journal of the Royal Statistical Society. Series A (General), 142(3):381–382, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982490.

Ray:1979:BRBd

[1280] W. D. Ray. Book review: Developments in Statistics, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 142(3):382, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982491.

Mackie:1979:BRB

[1281] Peter Mackie. Book review: Reviews of U. K. Statistical Sources: Volume VII: Road Passenger Transport, by W. F. Maunder; D. L. Munby; Reviews of U. K. Statistical Sources: Road Goods Transport. by W. F. Maunder; A. H.

Watson. Journal of the Royal Statistical Society. Series A (General), 142 (3):382–383, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982492.

Wilkinson:1979:BRBb

[1282] Margaret Wilkinson. Book review: British Economic Policy, 1960–74, by F. T. Blackaby. Journal of the Royal Statistical Society. Series A (General), 142(3):383–384, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982493.

Bartholomew:1979:BRBb

[1283] D. J. Bartholomew. Book review: Mathematics of Manpower Planning, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 142 (3):384–385, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982494.

Nelder:1979:BRB

[1284] J. A. Nelder. Book review: The General Linear Model, Data Analysis in the Social and Behavioural Sciences, by R. L. Horton. Journal of the Royal Statistical Society. Series A (General), 142 (3):385, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982495.

Kingman:1979:BRBc

[1285] J. F. C. Kingman. Book review: Introduction to Probability and Measure, by K. R. Parthasarathy. Journal of the Royal Statistical Society. Series A (General), 142(3):385–386, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982496.

Holgate:1979:BRBb

[1286] P. Holgate. Book review: Geometric Probability, by H. Solomon. Journal of the Royal Statistical Society. Series A (General), 142(3):386, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982497.

Marriott:1979:BRBa

[1287] F. H. C. Marriott. Book review: Discrete Discriminant Analysis, by M. Goldstein; W. R. Dillon. Journal of the Royal Statistical Society. Series A (General), 142(3):386–387, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982498.

Freeman:1979:BRB

[1288] P. R. Freeman. Book review: Finite Dynamic Programming, by D. J. White. Journal of the Royal Statistical Society. Series A (General), 142 (3):387, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982499.

Haley:1979:BRBa

[1289] K. B. Haley. Book review: Mathematical Programming and Control Theory, by B. D. Craven. Journal of the Royal Statistical Society. Series A (General), 142(3):387–388, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982500.

Glazebrook:1979:BRB

[1290] K. D. Glazebrook. Book review: Principles of Dynamic Programming,Part 1. Basic Analytic and Computational Methods, by R. E. Larson; J. L. Casti.

Journal of the Royal Statistical Society. Series A (General), 142(3):388, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982501.

Haley:1979:BRBb

[1291] K. B. Haley. Book review: Mathematical Programming in the Design of Large Physical Systems, by S. Petriccione; L. Piccioni. Journal of the Royal Statistical Society. Series A (General), 142 (3):388–389, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982502.

Thomas:1979:BRBb

[1292] D. A. H. Thomas. Book review: An Introduction to Statistics, by A. H. Hall. Journal of the Royal Statistical Society. Series A (General), 142(3):389, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982503.

Shanbhag:1979:BRB

[1293] D. N. Shanbhag. Book review: Queueing Theory: Working Examples and Problems, by J. Murdoch. Journal of the Royal Statistical Society. Series A (General), 142(3):389–390, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982504.

Hawkes:1979:BRB

[1294] Alan G. Hawkes. Book review: Biomathematics and Cell Kinetics, by A. J. Valleron; P. D. M. Macdonald. Journal of the Royal Statistical Society. Series A (General), 142(3):390, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982505.

Wilkinson:1979:BRBc

[1295] R. K. Wilkinson. Book review: Contemporary Economic Analysis, by M. J. Artis; A. R. Nobay. Journal of the Royal Statistical Society. Series A (General), 142(3):390–391, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982506.

Wilkinson:1979:BRBd

[1296] R. K. Wilkinson. Book review: Using Statistics in Managing the Economy, by R. Thomas; Bernard Eccleston; Milton Keynes. Journal of the Royal Statistical Society. Series A (General), 142 (3):391, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982507.

Greenfield:1979:BRB

[1297] Tony Greenfield. Book review: Multiple Sclerosis: An Analysis of 812 Cases by Means of Electronic Data Processing, by S. Poser. Journal of the Royal Statistical Society. Series A (General), 142(3):392, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982508.

Davidson:1979:BRB

[1298] James Davidson. Book review: A Micro-to-macro Model of the Swedish Economy, by G. Eliasson. Journal of the Royal Statistical Society. Series A (General), 142(3):392–393, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982509.

Tukey:1979:HFD

[1299] John W. Tukey. Harold F. Dodge, 1893–1976. Journal of the Royal Statistical Society. Series A (General), 142 (3):394, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982510.

Anonymous:1979:RPc

[1300] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 142(3):395–403, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982511.

Kish:1979:COP

[1301] Leslie Kish. Corrigenda: Optima and proxima in linear sample designs. Journal of the Royal Statistical Society. Series A (General), 142(3):404, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982512. See [799].

Aitkin:1979:CAU

[1302] Murray Aitkin. Corrigenda: The analysis of unbalanced cross-classifications. Journal of the Royal Statistical Society. Series A (General), 142(3):404, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982513. See [1118].

Wetherill:1979:CSI

[1303] G. Barrie Wetherill and Julian Köllerström. Corrigenda: Sampling inspection simplified. Journal of the Royal Statistical Society. Series A (General), 142(3):404, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982514. See [1214].

Byron:1979:CEL

[1304] R. P. Byron. Corrigenda: The estimation of large social account matrices. Journal of the Royal Statistical Society. Series A (General), 142(3):405, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982515. See [1150].

Anonymous:1979:BMc

[1305] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 142(3):i–ii, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982516.

Anonymous:1979:VI

[1306] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 142(4):537– 540, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982549.

Anonymous:1979:FMd

[1307] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 142(4):??, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982550.

Goldstein:1979:SMA

[1308] Harvey Goldstein. Some models for analysing longitudinal data on educational attainment. Journal of the Royal Statistical Society. Series A (General), 142(4):407–442, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982551.

Stuart:1979:POP

[1309] A. Stuart, Norman L. Webb, and D. Butler. Public opinion polls. Journal of the Royal Statistical Society. Series A (General), 142(4):443–467, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982552.

Whitmore:1979:IGM

[1310] G. A. Whitmore. An inverse Gaussian model for labour turnover. Journal of the Royal Statistical Society. Series A (General), 142(4):468–478, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982553.

Anonymous:1979:POH

[1311] Anonymous. Proceedings of the one hundred and forty-fifth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 142(4):479, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982554.

Anonymous:1979:HTR

[1312] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 142(4): 480, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982555.

Anonymous:1979:ARC

[1313] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 142 (4):481–503, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982556.

Yates:1979:BRB

[1314] F. Yates. Book review: R. A. Fisher.

The Life of a Scientist, by Joan Fisher Box. Journal of the Royal Statistical Society. Series A (General), 142(4): 504–506, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982557.

Morgan:1979:BRB

[1315] R. W. Morgan. Book review: Statistical Techniques for Manpower Planning, by David J. Bartholomew; Andrew F. Forbes. Journal of the Royal Statistical Society. Series A (General), 142 (4):506–507, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982558.

Healy:1979:BRBb

[1316] M. J. R. Healy. Book review: Statistical Method in Biological Assay, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 142 (4):507, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982559.

Lindley:1979:BRB

[1317] D. V. Lindley. Book review: Evidence, Proof and Probability, by Richard Eggleston. Journal of the Royal Statistical Society. Series A (General), 142 (4):507–508, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982560.

Cooper:1979:BRB

[1318] P. J. Cooper. Book review: The Theory of Competing Risks, by H. A. David; M. L. Moeschlberger. Journal of the Royal Statistical Society. Series A (General), 142(4):508–509, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982561.

Kingman:1979:BRBd

[1319] J. F. C. Kingman. Book review: Probability Theory: Independence, Interchangeability, Martingales, by Y. S. Chow; H. Teicher; Diffusions, Markov Processes and Martingales, Vol. 1: Foundations, by D. Williams. Journal of the Royal Statistical Society. Series A (General), 142(4):509, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982562.

Puzey:1979:BRB

[1320] A. S. Puzey. Book review: Manual of Mortality Analysis, by H. Campbell; H. Hansluwka; S. Koller; S. Mandel; Life Table and Mortality Analysis, by Chin Long Chiang. Journal of the Royal Statistical Society. Series A (General), 142 (4):510–511, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982563.

Marriott:1979:BRBb

[1321] F. H. C. Marriott. Book review: Geometrical Probability and Biological Structures: Buffon’s 200th Anniversary, by R. E. Miles; J. Serra. Journal of the Royal Statistical Society. Series A (General), 142(4):511, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982564.

Nankivell:1979:BRB

[1322] Owen Nankivell. Book review: Long Range Forecasting — From Crystal Ball to Computer, by J. Scott-Armstrong. Journal of the Royal Statistical Society. Series A (General), 142 (4):512, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982565.

Goldstein:1979:BRB

[1323] Michael Goldstein. Book review:

Regression Analysis by Example, by S. Chatterjee; B. Price. Journal of the Royal Statistical Society. Series A (General), 142(4):512–513, ???? 1979. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982566.

Ray:1979:BRBe

[1324] W. D. Ray. Book review: Gaussian Random Processes, by I. A. Ibragimov; Y. A. Rozanov. Journal of the Royal Statistical Society. Series A (General), 142(4):513, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982567.

Diggle:1979:BRB

[1325] Peter J. Diggle. Book review: Competition for Space and the Structure of Ecological Communities, by Peter Yodzis. Journal of the Royal Statistical Society. Series A (General), 142(4): 513–514, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982568.

Skene:1979:BRB

[1326] A. M. Skene. Book review: The Analysis of Cross-Tabulated Data, by G. J. G. Upton. Journal of the Royal Statistical Society. Series A (General), 142 (4):514–515, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982569.

Shearn:1979:BRB

[1327] D. C. S. Shearn. Book review: Interfaces between Computer Sciences and Operations Research, by J. K. Lenstra; A. H. G. Rinnooy Kan; P. van Emde Boas. Journal of the Royal Statistical Society. Series A (General), 142(4): 515, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982570.

Yates:1979:GMC

[1328] Frank Yates. Gertrude Mary Cox, 1900–1978. Journal of the Royal Statistical Society. Series A (General), 142 (4):516–517, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982571.

Berman:1979:JS

[1329] L. S. Berman. Julius Shiskin, 1912– 1978. Journal of the Royal Statistical Society. Series A (General), 142(4): 517–518, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982572.

Anonymous:1979:RPd

[1330] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 142(4):519–536, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982573.

Anonymous:1979:BMd

[1331] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 142(4):i–vi, ???? 1979. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982574.

Hill:1981:CAD

[1332] I. D. Hill. Corrigenda: Approximating the distribution of Greenwood’s statistic with Johnson distributions. Journal of the Royal Statistical Society. Series A (General), 144(3):388, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981821.